### v1

# Accuracy of Discretely Sampled Stochastic Policies in Continuous-time Reinforcement Learning

Abstract.
Stochastic policies are widely used in continuous-time reinforcement learning algorithms. However, executing a stochastic policy and evaluating its performance in a continuous-time environment remain open challenges. This work introduces and rigorously analyzes a policy execution framework that samples actions from a stochastic policy at discrete time points and implements them as piecewise constant controls. We prove that as the sampling mesh size tends to zero, the controlled state process converges weakly to the dynamics with coefficients aggregated according to the stochastic policy. We explicitly quantify the convergence rate based on the regularity of the coefficients and establish an optimal first-order convergence rate for sufficiently regular coefficients. Additionally, we show that the same convergence rates hold with high probability concerning the sampling noise, and further establish a $1/2$-order almost sure convergence when the volatility is not controlled. Building on these results, we analyze the bias and variance of various policy evaluation and policy gradient estimators based on discrete-time observations. Our results provide theoretical justification for the exploratory stochastic control framework in [H. Wang, T. Zariphopoulou, and X.Y. Zhou, J. Mach. Learn. Res., 21 (2020), pp. 1-34].
Key words and phrases:
policy execution, stochastic policy, continuous-time reinforcement learning, exploratory control, piecewise constant control, convergence rate
1. Introduction
Continuous-time RL
Reinforcement learning (RL) has catalyzed significant changes across various domains by attacking complex sequential decision-making challenges in a data-driven way. The theoretical framework for analyzing RL problems and designing RL algorithms has been predominantly based on the discrete-time Markov decision processes [5, 30]. However, many practical applications require real-time interaction with environments where time progresses continuously, such as autonomous driving, stock trading, and vehicle routing. Recent studies [33, 38, 20, 12] have shown that certain algorithms originally designed for discrete-time RL can become unstable when applied in high-frequency settings. This highlights the need for a new theoretical framework for designing RL algorithms that are robust to time discretization.
To address this issue, many suggest formulating the problem in a continuous-time framework from the outset and only introducing time discretization at the final implementation stage. Here, time discretization refers to the frequency at which data is collected and actions are updated, while the underlying system itself evolves continuously according to its intrinsic dynamics. Early studies on continuous-time RL have primarily focused on deterministic systems, where state variables follow ordinary differential equations; see, e.g., [1, 9, 33, 26, 22].
Exploratory control and stochastic policy
Recently, [36] proposed a novel exploratory control formulation with stochastic policies for continuous-time RL problems, which has served as the foundation for designing various RL algorithms; see, e.g., [18, 19, 20, 15, 8, 11, 37, 12, 29, 32].
This control formulation in [36] has three key features: (1) The agent incorporates stochastic policies into the control problem, which models exploration in RL. These policies are mappings from the state space to distributions over the action space, aligning with the notion of “relaxed control” in classical control theory. (2) The agent continuously samples actions from a stochastic policy, and the state process evolves according to a new diffusion process, with its coefficients integrated with respect to the sampling distribution specified by the stochastic policy. This auxiliary state dynamic is often referred to as the exploratory or aggregated dynamics. (3) The optimal policy remains stochastic due to the inclusion of an additional entropy regularizer in the cost objective, in contrast to the deterministic/strict policies typically found in standard control problems.
Execution of stochastic policies
However, despite its success in algorithm design, the use of stochastic policies in continuous-time RL problems introduces theoretical and practical gaps. Specifically, the connection between exploratory dynamics with aggregated coefficients and state dynamics controlled by frequent sampling from a stochastic policy remains unjustified. In fact, interacting with a diffusion process by continuously sampling from a stochastic policy requires generating uncountably many independent random variables, which raises measure-theoretical issues and renders the resulting controlled state dynamics ill-defined (see the end of Section 2 for more details).
To address this issue, [32] proposes approximating exploratory dynamics by sampling actions from a stochastic policy only at discrete time points. This approach naturally aligns with the practical implementation of stochastic policies. However, for algorithms designed based on aggregated dynamics, executing a stochastic policy in discrete time introduces a bias that depends on the sampling frequency. This leads to the following natural question:
What is the precise impact of sampling frequency on the performance of a stochastic policy and the corresponding learning algorithms?
To the best of our knowledge, this question has only been addressed in the context of linear-quadratic (LQ) RL problems with Gaussian policies [12, 32]. In this special setting, due to the LQ structure, the analysis is simplified by examining the discrepancy between the first and second moments of the aggregated dynamics and the controlled dynamics with sampled actions. Unfortunately, these techniques developed specifically for LQ RL problems are clearly not applicable to general RL problems. In particular, it is essential to quantify the difference between the entire distributions of two state processes.
Our contributions
This paper analyzes the accuracy of the discrete-time execution of stochastic policies for general diffusion processes and quantifies its impact on the associated learning algorithms.
  • We show that when actions are sampled from a stochastic policy at a discrete time grid $\mathscr{G}$ and executed as a piecewise constant control process, the resulting state process, referred to as the sampled dynamics, converges weakly to the aggregated dynamics as the time stepsize approaches zero. In particular, under sufficient regular coefficients, we show that the convergence rate is of first order $O(|\mathscr{G}|)$, with $|\mathscr{G}|$ being the time stepsize (Theorem 4.1). This matches the optimal convergence rate for standard piecewise constant control approximations [16]. We further quantify the precise weak convergence rate for cases with less regular coefficients (Theorem 4.2).
    The convergence results rigorously justify the exploratory control framework proposed in [36] (see Section 2). The weak convergence rate will be used to analyze the performance of learning algorithms based on observations from the sampled dynamics.
  • We prove that the sampled dynamics, in fact, converges almost surely with respect to the randomness in policy execution. Specifically, we show that the weak convergence order to the aggregated dynamics holds with high probability with respect to the randomness used to sample actions (Theorem 4.3). We further show that when the volatility is uncontrolled, the sampled dynamics converges almost surely to the aggregated dynamics with an order of $\mathcal{O}(|\mathscr{G}|^{1/2})$, for each fixed realization of the Brownian motion (Theorem 4.4). The analysis leverages concentration inequalities for bounded and sub-exponential martingale difference sequences to control the concentration behavior of the sampling errors in terms of the sampling frequency (see the discussions below Theorems 4.3 and 4.4).
    The almost sure convergence result indicates that the convergence of the sampled dynamics to the aggregated dynamics can be interpreted as a law of large numbers in time, i.e., the convergence can be achieved by refining the sampling frequency while fixing the realization of the system noise. Note that such an almost sure convergence fails when the volatility is controlled, as shown in Example 4.6.
  • We apply the weak convergence rate to quantify the error incurred when implementing continuous-time RL algorithms based on observations from sampled dynamics. In particular, we show that evaluating the reward involving aggregated dynamics by discrete sampling and piecewise constant execution has a bias of the order $O(|\mathscr{G}|)$. Calculating the martingale orthogonality conditions proposed in [18] using discretely sampled stochastic policies also introduces a bias of the same order, $O(|\mathscr{G}|)$. The policy gradient representation developed in [19] is an informal integration that involves the realized actions and the state variables. We rigorously show that using the forward Euler scheme to calculate such terms with discretely sampled stochastic policies is a biased estimator for the policy gradient, and the bias is $O(|\mathscr{G}|)$ and variance is bounded. The q-learning with quadratic variation penalty for the risk-sensitive RL proposed in [17] is also shown to have a similar order of discretization error.
Our approaches and most related works
This paper quantifies the accuracy of the discrete-time execution of a stochastic policy and its impact on continuous-time RL algorithms. The accuracy of policy execution has been studied in [14, 32] for LQ problems with Gaussian policies. However, as mentioned earlier, the techniques developed in [14, 32] are not applicable to RL problems with general state dynamics and stochastic policies considered in this paper.
For general state dynamics, the weak convergence of sampled dynamics has recently been studied by [4, 3]. They adopt a different definition of sampled dynamics, treating it as being driven by a random measure, and establish the weak convergence of the state process without quantifying the convergence order. In contrast to the time-varying control considered in [4, 3], we analyze sampled dynamics controlled by a piecewise constant control process. Moreover, we take a completely different approach and establish the weak convergence rate under various regularity conditions by leveraging the solution regularity of the associated partial differential equations (PDEs). Beyond weak convergence, we also establish convergence in the almost sure sense with respect to the policy execution noise. These strong convergence results require a careful analysis of the concentration properties of the state dynamics concerning the policy execution noise.
In addition to discrete sampling error, we also analyze the error incurred when approximating continuous-time RL algorithms using discrete-time observations. [18] analyzes a similar error but assumes that observations are generated from the aggregated dynamics. [2] analyzes both errors for LQ problems with uncontrolled volatility.
Our analysis of the policy execution error is fundamentally different from classical numerical analysis of time discretizations for stochastic differential equations (SDEs). In particular, numerical analysis of SDEs focuses on the error caused by discretizing the system noise, such as Brownian motion, while both the SDE and its time-discretized version have the same coefficients. In our setting, we treat the controlled SDE as a black-box environment evolving continuously and focus on the error introduced by the piecewise constant execution of a stochastic policy. As a result, the environment noise is not discretized, but the limiting SDE (the aggregated dynamics) and the sampled dynamics have different coefficients. The former involves aggregated coefficients, while the latter has random coefficients determined by the specific procedure used to sample the policy.
The rest of the paper is organized as follows. Section 2 briefly reviews the use of stochastic policies in continuous-time RL. Section 3 presents the problem setup and introduces the sampled dynamics rigorously. Section 4 presents the main convergence results of the sampled dynamics, and Section 5 applies them to analyze continuous-time RL algorithms. Finally, we conclude in Section 6. The proofs of all statements are included in Section 7.
2. Preliminary
In this section, we review the exploratory control formulation and the use of stochastic policies for continuous-time RL algorithms. The precise assumptions and setup can be found in Section 3.
2.1. Classical stochastic control
Recall that in a classical stochastic control problem, an agent aims to control the state process and optimize an objective functional through an adapted control process taking values in a Polish space $A$. Specifically, given an admissible control process $\{a_{t}\}_{t\geq 0}$, the corresponding state process $\{X_{t}\}_{t\geq 0}$ is governed by the following dynamics: [^1] To simplify the presentation, we assume that the state processes have a deterministic initial condition. However, our analysis and results extend naturally to cases with a random initial condition, provided it is independent of the Brownian motion $W$ and satisfies appropriate integrability conditions. We also assume that the Brownian motion’s dimension equals the state variable’s dimension; otherwise, we can introduce more Brownian motions or state variables to make them equal.
$\mathrm{d}X_{t}=b(t,X_{t},a_{t})\mathrm{d}t+\sigma(t,X_{t},a_{t})\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.1)
where $x_{0}\in{\mathbb{R}}^{d}$ is a given initial state, $b:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d\times d}$ are given drift and volatility coefficients, respectively, and $W$ is an $d$-dimensional Brownian motion supported on a probability space $(\Omega,\mathcal{F},\mathbb{P})$. The agent’s objective is to maximize the following objective
$\mathbb{E}\left[\int_{0}^{T}r(t,X_{t},a_{t})\mathrm{d}t+h(X_{T})\right]$ (2.2)
over all admissible controls $\{a_{t}\}_{t\geq 0}$, where $r:[0,T]\times{\mathbb{R}}^{d}\times A\to{\mathbb{R}}$ is a given instantaneous reward function, and $h:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a given terminal reward function.
It is known that under suitable conditions (see [39]), an optimal control for (2.2) exists and is given by a feedback form, i.e., there exists a measurable function $\bm{u}:[0,T]\times\mathbb{R}^{d}\to A$ such that the optimal control is obtained by $a_{t}=\bm{u}(t,X_{t}^{\bm{u}})$ for all $t\in[0,T]$, where $\{X_{t}^{\bm{u}}\}_{t\geq 0}$ satisfies the following dynamics:
$\mathrm{d}X_{t}=b(t,X_{t},\bm{u}(t,X_{t}))\mathrm{d}t+\sigma(t,X_{t},\bm{u}(t,X_{t}))\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.3)
The function can be determined using the coefficients $b,\sigma,r,$ and $h$. Once determined, the optimal control can be implemented by evaluating ${\bm{u}}$ at the current state of the system.
2.2. RL, stochastic policy and exploratory control
In the RL problem, the agent seeks to optimize the objective (2.2) by directly interacting with the dynamics (2.1), experimenting with different actions, and refining their strategies based on the observed responses. Unlike traditional control settings, the agent treats (2.1) as a black-box environment evolving in continuous time and learns an optimal policy solely from observations, without direct access to the coefficients $b,\sigma,r,$ and $h$.
To avoid getting trapped in local optima, a key feature of RL is explicitly exploring the environment through randomized actions. A common approach is to use stochastic policies $\bm{\pi}:[0,T]\times{\mathbb{R}}^{d}\to\mathcal{P}(A)$, which maps the current state to a probability distribution over the action space. Given a stochastic policy $\bm{\pi}$, the agent interacts with the environment by sampling noisy actions from $\bm{\pi}$ and observing the resulting system states. Over time, the agent refines the stochastic policy to optimize the objective.
To describe the performance of a stochastic policy $\bm{\pi}$, [36] suggests that the associated state process $\tilde{X}$ evolves according to the following dynamics (called the exploratory dynamics therein):
${\mathrm{d}}X_{t}=\tilde{b}^{\bm{\pi}}(t,X_{t})\mathrm{d}t+\tilde{\sigma}^{\bm{\pi}}(t,X_{t})\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.4)
with the coefficients $\tilde{b}^{\bm{\pi}}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ and $\tilde{\sigma}^{\bm{\pi}}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d\times d}$ given by
$\tilde{b}^{\bm{\pi}}(t,x)\coloneqq\int_{A}b(t,x,a)\bm{\pi}(\mathrm{d}a|t,x),\quad\tilde{\sigma}^{\bm{\pi}}(t,x)\coloneqq\sqrt{\int_{A}(\sigma\sigma^{\top})(t,x,a)\bm{\pi}({\mathrm{d}}a|t,x)},$ (2.5)
where $(\cdot)^{1/2}$ is the matrix square root of positive semidefinite matrices, and $(\cdot)^{\top}$ is the matrix transpose. Moreover, the associated objective function is given by:
$\mathbb{E}^{W}\left[\int_{0}^{T}\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t+h(\widetilde{X}_{T})\right],\quad\textnormal{with $\tilde{r}(t,x)\coloneqq\int_{A}r(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. }$ (2.6)
The dynamics (2.4) and the objective (2.6) are derived heuristically in [36] under the assumption that the agent interacts with the system (2.1) by continuously sampling independent random actions from the policy $\bm{\pi}$. As a result of these random actions, a “law-of-large-numbers” heuristic suggests that the infinitesimal change in the state at time $t$ is determined by the coefficients in (2.5), aggregated over the sampling distribution defined by $\bm{\pi}$. Likewise, the instantaneous reward is also aggregated with respect to this sampling distribution. In the sequel, we refer to (2.4) as the aggregated dynamics to highlight this aggregating effect.
However, the rigorous justification of the connection between (2.4) and the state dynamics resulting from the adoption of a stochastic policy $\bm{\pi}$ in (2.1), as well as the relationship between (2.6) and the original reward functional (2.2), remains to be established. Indeed, interacting with the SDE (2.1) by continuously sampling from a stochastic policy creates measure-theoretical issues. As already pointed out in [32, Remark 2.1], it is impossible to construct a family of non-constant random variables $(\xi_{t})_{t\in[0,1]}$ such that $(\xi_{t})_{t\in[0,1]}$ is (essentially) pairwise independent and $t\mapsto\xi_{t}$ is Lebesgue measurable. This implies that if one controls (2.1) by continuously generating independent actions, the resulting coefficients are not progressively measurable, rendering the conventional stochastic integral ill-defined; see [3] for more discussion. Moreover, it is practically infeasible to generate these independent samples continuously.
Therefore, both from theoretical and practical perspectives, evaluating the performance of a stochastic policy $\bm{\pi}$ requires discretely sampling actions from the policy and applying them to (2.1), as suggested by [32]. This work quantifies the accuracy of such a discrete execution of a stochastic policy and its impact on the corresponding RL algorithms.
3. Probabilistic Setup for Aggregated and Sampled Dynamics
This section introduces a rigorous probabilistic framework for the aggregated dynamics under a given stochastic policy and the associated sampled dynamics with a finite number of intervention times, as outlined in Section 2.
3.1. Aggregated dynamics
Fix a probability space $(\Omega^{W},\mathcal{F}^{W},\mathbb{P}^{W})$ on which a $d$-dimensional Brownian motion $\{W_{t}\}_{t\geq 0}$ is defined, and let ${\mathbb{F}}^{W}\coloneqq\{\mathcal{F}_{t}^{W}\}_{t\geq 0}$ be the $\mathbb{P}^{W}$-completion of the filtration generated by $W$, which satisfies the usual conditions of completeness and right-continuity. Throughout this paper, we impose the following regularity condition on the action space $A$, the drift and diffusion coefficients $b$ and $\sigma$, and the stochastic policy $\bm{\pi}$.
H. 1.
  1. (1)
    $({A},d_{A})$ is a Polish space equipped with the Borel $\sigma$-algebra $\mathcal{B}({A})$, and $a_{0}\in A$;
  2. (2)
    $b:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d\times d}$ are measurable functions. There exists $C\geq 0$ such that for all $\varphi\in\{b,\sigma\}$, $t\in[0,T]$, $x,x^{\prime}\in\mathbb{R}^{d}$ and $a\in{A}$,
    $|\varphi(t,x,a)-\varphi(t,x^{\prime},a)|\leq C|x-x^{\prime}|,\quad|\varphi(t,x,a)|$ $\leq C(1+|x|+d_{A}(a,a_{0}));$
  3. (3)
    $\bm{\pi}:[0,T]\times\mathbb{R}^{d}\times\mathcal{B}({A})\to[0,1]$ is a probability kernel. For all $p\geq 2$, there exists $C_{p}\geq 0$ such that for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$, $\int_{A}d_{A}(a,a_{0})^{p}\bm{\pi}({\mathrm{d}}a|t,x)\leq C_{p}(1+|x|^{p})$.
Condition (H.12) imposes the standard Lipschitz regularity on the system coefficients commonly used in the stochastic control literature (see, e.g., [10, 39]), and Condition (H.13) requires mild growth and integrability properties for the stochastic policy. Under (H.1), $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ in (2.5) are well-defined measurable functions and satisfy $|\tilde{b}^{\bm{\pi}}(t,x)|+|\tilde{\sigma}^{\bm{\pi}}(t,x)|\leq C(1+|x|)$ for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$.
We assume that $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are sufficiently regular so that the aggregated dynamic (2.4) is well-posed.
H. 2.
The functions $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are sufficiently regular so that (2.4) admits a unique strong solution $\tilde{X}$ on the filtered probability space $(\Omega^{W},\mathcal{F}^{W},{\mathbb{F}}^{W},\mathbb{P}^{W})$.
Here, we focus on strong solutions of (2.4) rather than weak solutions, which allows us to identify $W$ as the only driving noise in the state process $\tilde{X}$. This simplifies the construction of the sampled dynamics and the subsequent error estimates in terms of the sampling frequency. Under (H.1), Condition (H.2) holds if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are locally Lipschitz continuous in the state variable (cf. [27, Theorem 3.4]) and can be ensured via suitable regularity conditions on $\bm{\pi}$ (e.g., [19, Lemma 2]). Some typical sufficient conditions on $b,\sigma$ and $\bm{\pi}$ to guarantee these regularity conditions on $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are given in Examples 4.1, 4.2 and 4.3.
3.2. Sampled dynamics with randomized actions
We proceed to define the state process with random actions sampled according to the stochastic policy $\bm{\pi}$. Rather than focusing on a specific sampling procedure for $\bm{\pi}$, we adopt a more abstract formulation of the sampling process defined as follows.
Definition 3.1.
Assume the notation of (H.1). We say a tuple $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi},E,\xi,\phi)$ a sampling procedure of the policy $\bm{\pi}$ if $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$ is a complete probability space, $(E,\mathcal{B}(E))$ is a Borel space, $\xi:\Omega^{\xi}\to E$ is a random variable and $\phi:[0,T]\times{\mathbb{R}}^{d}\times E\to A$ is a measurable function such that for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$, $\Omega^{\xi}:\omega\mapsto\phi(t,x,\xi(\omega))\in A$ has the distribution $\bm{\pi}({\mathrm{d}}a|t,x)$ under the measure $\mathbb{P}^{\xi}$.
By Definition 3.1, $(E,\xi,\phi)$ provides a framework for executing the policy $\bm{\pi}$ by sampling a random action $a_{t}\coloneqq\phi(t,X_{t},\xi)$ from the distribution $\bm{\pi}({\mathrm{d}}a|t,X_{t})$ at a given time $t\in[0,T]$ and state $X_{t}$. This action, conditioned on the $\sigma$-algebra $\sigma\left\{(X_{s})_{s\in[0,t]}\right\}$, is independent of both $(X_{s})_{s\in[0,t]}$ and the underlying Brownian motion $W$.
Remark 3.1.
As $A$ is a Polish space, a sampling procedure of a probability kernel $\bm{\pi}$ always exists. Indeed, by [21, Lemma 2.22], it suffices to take $E=[0,1]$ and $\xi$ as a uniform random variable on $E$. This approach has been utilized in [11, 4, 3]. However, the corresponding function $\phi$ is guaranteed to exist only through a measure-theoretic argument, making it difficult to implement such a sampling procedure.
Here we work with a general Borel space $E$, instead of restricting $E$ to $[0,1]$. This generalization accommodates commonly used sampling procedures that generate actions using multiple independent normal random variables, as seen in applications such as multivariate Gaussian policies [32] or suitable Markov chain Monte Carlo methods. In Section 4, we establish the weak convergence of the control-randomized SDE under the mere measurability assumption of $\phi$. Additional regularity of the function $\phi$ will be imposed to obtain strong convergence results.
Now, fix a sampling procedure $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi},E,\xi,\phi)$ of $\bm{\pi}$, let ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$ and let $(\Omega^{\xi_{n}},\mathcal{F}^{\xi_{n}},\mathbb{P}^{\xi_{n}},\xi_{n})_{n\in{\mathbb{N}}_{0}}$ be independent copies of $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi},\xi)$. Consider a probability space of the following form:
$(\Omega,\mathcal{F},\mathbb{P})\coloneqq\bigg{(}\Omega^{W}\times\prod_{n=0}^{\infty}\Omega^{\xi_{n}},\mathcal{F}^{W}\otimes\bigotimes_{n=0}^{\infty}\mathcal{F}^{\xi_{n}},\mathbb{P}^{W}\otimes\bigotimes_{n=0}^{\infty}\mathbb{P}^{\xi_{n}}\bigg{)},$ (3.1)
where $(\Omega^{W},\mathcal{F}^{W},\mathbb{P}^{W})$ supports the Brownian motion $W$ in (2.4), and for each $n\in{\mathbb{N}}_{0}$, $(\Omega^{\xi_{n}},\mathcal{F}^{\xi_{n}},\mathbb{P}^{\xi_{n}})$ supports the random variable $\xi_{n}$ used to generate the random control at the grid point $t_{n}$. With a slight abuse of notation, we complete the probability space $(\Omega,\mathcal{F},\mathbb{P})$ by augmenting $\mathcal{F}$ with $\mathbb{P}$-null sets, and extend canonically any random variable on a marginal space to the whole space $\Omega$, e.g., $W(\omega^{W},\omega^{\xi})=W(\omega^{W})$ and $\widetilde{X}(\omega^{W},\omega^{\xi})=\widetilde{X}(\omega^{W})$ for all $\omega^{W}\in\Omega^{W}$ and $\omega^{\xi}\in\prod_{n=0}^{\infty}\Omega^{\xi_{n}}$.
Given a time grid $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ of $[0,T]$, we consider interacting with the state dynamics (2.1) by sampling actions at the grid points in $\mathscr{G}$ according to the policy $\bm{\pi}$. More precisely, we consider the dynamics such that for all $i=0,\ldots,n-1$ and all $t\in[t_{i},t_{i+1}]$,
$\begin{split}X_{t}&=X_{t_{i}}+\int_{t_{i}}^{t}b(s,X_{s},a_{t_{i}}){\mathrm{d}}s+\int_{t_{i}}^{t}\sigma(s,X_{s},a_{t_{i}}){\mathrm{d}}W_{s},\quad\textnormal{with $a_{t_{i}}=\phi(t_{i},X_{t_{i}},\xi_{i})$};\quad X_{0}=x_{0},\end{split}$ (3.2)
which will be referred to as the sampled dynamics [^2] This definition slightly differs from the definition of the “grid sampling SDE” in [3, Section 3] even though both rely on drawing finitely many samples on the grid point. In our definition, the action sequence $a_{t}$ is a piecewise constant and is given by a sampling function at the grid points $a_{t_{i}}=\phi(t_{i},X_{t_{i}},\xi_{i})$. By contrast, in the grid sampling SDE in [3], the action sequence is evolving as $a_{t}=\phi(t,X_{t},\xi_{t_{i}})$, where the random source $\xi_{t_{i}}$ is piecewise constant. in the subsequent analysis. For notational convenience, we write (3.2) in the following equivalent form:
$\mathrm{d}X_{t}=b(t,X_{t},a_{\delta(t)})\mathrm{d}t+\sigma(t,X_{t},a_{\delta(t)})\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0}$ (3.3)
with $\delta(t)\coloneqq t_{i}$ for $t\in[t_{i},t_{i+1})$, and $a_{t_{i}}$ given in (3.2). The dynamics (3.3) can be viewed as an SDE with random coefficients.
Remark 3.2.
Although the definition of the sampled dynamics depends on the sampling scheme $\phi$ and $\xi$, the weak convergence rate of the sampled dynamics to the aggregated dynamics depends only on the properties of the stochastic policy $\bm{\pi}$, rather than the specific choice of $\phi$ and $\xi$ (Sections 4.1 and 4.2). The properties of $\phi$ will be used to quantify the pathwise error between the sampled dynamics and the aggregated dynamics (Section 4.3).
Note that the sampled dynamics evolve continuously over time while the control process remains constant within each subinterval. In particular, a random action $a_{t_{i}}$ is generated at $t_{i}$ and applied to the system over the interval $[t_{i},t_{i+1})$ before being updated to the next action $a_{t_{i+1}}$. This aligns with the RL setting, where the agent observes a continuous-time system at discrete time points and adjusts their actions based on these observations.
The following lemma shows that the sampled dynamics (3.3) admits a unique strong solution $X^{\mathscr{G}}$ which is adapted to the filtration generated by both the Brownian motion $W$ and the execution noise $\xi$. The proof is given in Appendix A.
Lemma 3.1.
Suppose (H.1) holds. Let $x_{0}\in{\mathbb{R}}^{d}$, $(\Omega,\mathcal{F},{\mathbb{F}},\mathbb{P})$ be the probability space defined in (3.1), and ${\mathbb{F}}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration such that $\mathcal{F}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{i=0}^{\infty}\}\vee\mathcal{N}$, where $\mathcal{N}$ be the $\mathbb{P}$-null set. Then for any grid $\mathscr{G}$, (3.3) admits a unique ${\mathbb{F}}$-adapted strong solution $X^{\mathscr{G}}$ on the space $(\Omega,\mathcal{F},\mathbb{P})$. Moreover, for all $p\geq 2$, there exists a constant $C_{p}\geq 0$ such that for all grids $\mathscr{G}$ and all $t\in[0,T]$, $\mathbb{E}\left[|X^{\mathscr{G}}_{t}|^{p}\right]\leq C_{p}(1+|x_{0}|^{p})e^{C_{p}t}$, and $X^{\mathscr{G}}_{t}$ is adapted to $\mathcal{G}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{t_{i}<t}\}\vee\mathcal{N}$.
4. Error Estimates for the Sampled Dynamics
In this section, we analyze the convergence rate of the sampled dynamics (3.3) to the aggregated dynamics (2.4) under various criteria as the time grid $\mathscr{G}$ is refined. In particular, we quantify the weak error for quantities involving expectations over the system noise $W$ (but not necessarily over the execution noises $(\xi_{i})_{i=0}^{\infty}$), and the pathwise error $\tilde{X}-X^{\mathscr{G}}$. To this end, we define the mesh size of a time grid $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ by $|\mathscr{G}|=\max_{0\leq i\leq n-1}(t_{i+1}-t_{i})$.
4.1. Weak error with expectations over execution noise
We start by analyzing the weak convergence rate of $X^{\mathscr{G}}$ to $\tilde{X}$ under the measure $\mathbb{P}$. That is, for a suitable test function $f:{\mathbb{R}}^{d}\to{\mathbb{R}}$, we quantify the error $\mathbb{E}[f(\widetilde{X}_{s})]-\mathbb{E}[f(X^{\mathscr{G}}_{s})]$ in terms of the time grid $\mathscr{G}$, where the expectation is taken over both system noise $W$ and the execution noises $(\xi_{i})_{i=0}^{\infty}$ (cf.  (3.1)).
4.1.1. Weak convergence for regular coefficients
To highlight the key ideas without introducing needless technicalities, we begin by providing the analysis for sufficiently smooth coefficients and test functions. In this setting, we obtain the maximum convergence rate, which is of the order $\mathcal{O}(|\mathscr{G}|)$ as $|\mathscr{G}|$ tends to zero within the regular time grids $\mathscr{P}_{[0,T]}$.
To specify the precise regularity conditions, we introduce the following notation: for each Euclidean space $E$ and $k\in{\mathbb{N}}$, $p>0$, let $C^{k}_{p}([0,T]\times{\mathbb{R}}^{d};E)$ be the space of functions $u:[0,T]\times{\mathbb{R}}^{d}\to E$ such that for all $r\in{\mathbb{N}}_{0}$ and multi-indices $s$ satisfying $2r+|s|\leq k$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}u$ exists and is continuous for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$, and they all have polynomial growth in $x$:
$\|u\|_{C_{p}^{k}}:=\sum_{2r+|s|\leq k}\sup_{(t,x)\in[0,T]\times\mathbb{R}^{d}}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x)\right|}{1+|x|^{p}}<\infty.$ (4.1)
Let $C^{k,0}_{p}([0,T]\times{\mathbb{R}}^{d}\times A;E)$ be the space of functions $u:[0,T]\times{\mathbb{R}}^{d}\times{A}\to E$ such that for all $r\in{\mathbb{N}}_{0}$ and multi-indices $s$ satisfying $2r+|s|\leq k$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)$ exists, is continuous with respect to $t$ and $x$ for all $(t,x,a)\in[0,T]\times\mathbb{R}^{d}\times{A}$ and they all have polynomial growth in $x$ and $a$:
$\|u\|_{C_{p}^{k,0}}:=\sum_{2r+|s|\leq k}\sup_{(t,x,a)\in[0,T]\times\mathbb{R}^{d}\times A}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)\right|}{1+|x|^{p}+d_{A}(a,a_{0})^{p}}<\infty,$ (4.2)
where $a_{0}\in A$ is the fixed element in (H.1). For simplicity, we omit the range space $E$ when no confusion arises. Moreover, we denote by $\mathcal{L}$ the generator of the aggregated dynamics (2.4) satisfying for all $\phi\in C^{2}([0,T]\times{\mathbb{R}}^{d})$,
$\mathcal{L}\phi(t,x)=\tilde{b}^{\bm{\pi}}(t,x)\nabla_{x}\phi(t,x)+\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top})(t,x)\textrm{Hess}_{x}\phi(t,x)\right).$ (4.3)
The following assumption presents the regularity conditions to facilitate the analysis.
H. 3.
  1. (1)
    $b$ and $\sigma\sigma^{\top}$ are in the space ${C}^{2,0}_{p}([0,T]\times{\mathbb{R}}^{d}\times A)$;
  2. (2)
    $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}$ are sufficiently regular, such that for any $f\in C_{p}^{4}(\mathbb{R}^{d})$ and any $t^{\prime}\in[0,T]$, the PDE
    $(\partial_{t}v+\mathcal{L}v)(t,x)=0,\quad(t,x)\in[0,t^{\prime})\times{\mathbb{R}}^{d};\quad v(t^{\prime},x)=f(x),\quad x\in{\mathbb{R}}^{d}$ (4.4)
    has a unique solution $v_{f}(\cdot,\cdot;t^{\prime})\in C_{p}^{4}([0,t^{\prime}]\times\mathbb{R}^{d})$. Moreover, $\|v_{f}(\cdot,\cdot;t^{\prime})\|_{C_{p}^{4}}\leq C\|f\|_{C_{p}^{4}}$, where $C$ is a constant depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},p$.
Remark 4.1.
Condition (H.3) holds if the coefficients of the aggregated dynamics (2.4) are sufficiently regular. For instance, if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are in the class $C_{p}^{4}$ and have bounded derivatives, then (H.3) can be verified using the probabilistic argument in [28].
Theorem 4.1.
Suppose (H.1), (H.2) and (H.3) hold with $p\geq 2$. Then for any $f\in C_{p}^{4}(\mathbb{R}^{d})$, there exists a constant $C_{p}\geq 0$ depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},b,\sigma,p$ such that for all grids $\mathscr{G}$,
$\sup_{t\in[0,T]}\left|\mathbb{E}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|\leq C_{p}\|f\|_{C_{p}^{4}}|\mathscr{G}|.$ (4.5)
Proof.
See Section 7.1.1. ∎
For a given time grid $\mathscr{G}$, we execute the policy only at the nodes $t_{i}$, and use the same action over the time interval $[t_{i},t_{i+1})$ to obtain the sampled dynamics $X^{\mathscr{G}}$. Theorem 4.1 illustrates that with this policy execution method, for a given test function $f$, the error between $\mathbb{E}[f(X^{\mathscr{G}}_{t})]$ and the original aggregated dynamics $\mathbb{E}^{W}[f(\widetilde{X}_{t})]$ converges to 00 at the optimal first-order convergence $O(|\mathscr{G}|)$.
We illustrate several common cases in which the required assumptions in Theorem 4.1 hold.
Example 4.1 (General smooth coefficients).
Assume $\bm{\pi}(\mathrm{d}a|t,x)$ admits a density function with respect to a measure $\mu$ on $A$ (e.g., Lebesgue measure), i.e.,
$\bm{\pi}(\mathrm{d}a|t,x)=H(t,x,a)\mu(\mathrm{d}a).$ (4.6)
If it holds for any integer $r\geq 0$ and multi-index $s$ satisfying $2r+|s|\leq 4$ that
  1. ()
    For $\phi\in\{b,\sigma\}$, $(t,x,a)\mapsto\partial_{t}^{r}\partial_{x}^{s}\phi(t,x,a)$ is bounded and continuous.
  2. ()
    $(t,x)\mapsto\int_{A}|\partial_{t}^{r}\partial_{x}^{s}H(t,x,a)|\mu(\mathrm{d}a)$ is bounded.
  3. ()
    $\sigma\sigma^{\top}$ is uniformly elliptic.
Then $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are of class $C_{p}^{4}$ with bounded derivatives.
Example 4.2 (Gaussian policies).
Assume $A=\mathbb{R}^{m}$ and $\bm{\pi}(\cdot|t,x)$ is Gaussian with respect to the Lebesgue measure on $\mathbb{R}^{m}$, i.e.,
${\bm{\pi}(\mathrm{d}a|t,x)}=\varphi(a|\mu(t,x),\Sigma^{2}(t,x))\mathrm{d}a,$ (4.7)
where $\varphi(\cdot|\mu,\Sigma^{2})$ is the Gaussian density with mean $\mu$ and variance $\Sigma^{2}$. If the following conditions hold:
  • (1)
    For $\phi\in\{b,\sigma\}$, integer $r\geq 0$ and multi-index $s_{1},s_{2}$ satisfying that $0<2r+|s_{1}|+|s_{2}|\leq 4$, $(t,x,a)\mapsto\partial_{t}^{l}\partial_{x}^{s_{1}}\partial_{a}^{s_{2}}\phi(t,x,a)$ is bounded and continuous.
  • (2)
    $\mu,\Sigma\in C_{p}^{4}$ with bounded derivatives.
  • (3)
    $\sigma\sigma^{\top}$ is uniformly elliptic.
Then $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are of class $C_{p}^{4}$ with bounded derivatives.
Examples 4.1 and 4.2 assume that $\sigma\sigma^{\top}$ is uniformly elliptic, which implies the uniform ellipticity of $(t,x)\mapsto\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. This nondegeneracy condition ensures that after taking the square root, $(t,x)\mapsto\tilde{\sigma}^{\bm{\pi}}(t,x)$ has the same regularity as $(t,x)\mapsto\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. The existence of a sufficiently regular square root $\tilde{\sigma}^{\bm{\pi}}$ can be ensured by replacing the nondegeneracy condition with certain structural conditions on $\sigma$ and $\bm{\pi}$. Specifically, if we can express $\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)=\lambda(t,x)\lambda^{\top}(t,x)$ with a sufficiently smooth function $\lambda(t,x)$, then $\tilde{\sigma}^{\bm{\pi}}=\lambda$ will have the desired regularity, even if it is degenerate. We provide several examples to illustrate this.
Example 4.3 (Uncontrolled volatility).
Consider the case in which $\sigma$ is uncontrolled, i.e., $\sigma(t,x,a)=\sigma(t,x)$. In this case,
$\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)=\sigma\sigma^{\top}(t,x).$ (4.8)
Note that we only require $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}=\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$, which means that we can choose $\tilde{\sigma}^{\bm{\pi}}$ to be $\sigma$ instead of $\sqrt{\sigma\sigma^{\top}}$. Therefore, the conditions on $\sigma$ in Examples 4.1 and 4.2 can be relaxed so that $\sigma$ is of class $C_{p}^{4}$ with bounded derivatives and it can be degenerate.
Example 4.4 (Gaussian policies with linear volatility).
Consider the case where the volatility linearly depends on the action, i.e., $\sigma(t,x,a)=a\beta^{\top}$ with $\beta\in R^{d}$ and (4.7) holds with $m=d$ and $\mu(t,x)=0$ and, then
$\int_{A}\sigma\sigma^{\top}(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)=\int_{\mathbb{R}^{d}}\Sigma(t,x)u\beta^{\top}\beta u^{\top}\Sigma(t,x)^{\top}\varphi(u)\mathrm{d}u.$ (4.9)
We can choose $\tilde{\sigma}^{\bm{\pi}}(t,x)=\Sigma(t,x)\sqrt{\int_{\mathbb{R}^{d}}u\beta^{\top}\beta u^{\top}\varphi(u)\mathrm{d}u}$. Therefore, $\tilde{\sigma}^{\bm{\pi}}$ can degenerate; the assumption in this case is that $\Sigma\in C_{p}^{4}$ has bounded derivatives.
If $d=1$, then we can extend the second case to a general form. Assume that $\bm{\pi}(\cdot|t,x)$ is the location scaled distribution, i.e., there is a function $\Pi(t,x)$ such that $\bm{\pi}(\cdot|t,x)$ satisfies for any measurable function $f$, $\int_{A}f(a)\bm{\pi}(\mathrm{d}a|t,x)=\int_{B}f(\Pi(t,x)u)\mu(\mathrm{d}u)$, where $(B,\mathscr{F}_{B},\mu)$ is a measure space. Let $\sigma(t,x,a)=h(t,x)a$ for a bounded $h$. Moreover, assume that $h\geq 0$ and $\Pi\geq 0$. We have
$(\tilde{\sigma}^{\bm{\pi}}(t,x))^{2}=\int_{B}h^{2}(t,x)\Pi^{2}(t,x)u^{2}\mu(\mathrm{d}u).$ (4.10)
If we can choose $\tilde{\sigma}^{\bm{\pi}}(t,x)=h(t,x)\Pi(t,x)\sqrt{\int_{B}u^{2}\mu(\mathrm{d}u)}$, then $\tilde{\sigma}^{\bm{\pi}}$ can degenerate as long as $h\Pi\in C^{4}_{p}$ and has bounded derivatives.
4.1.2. Weak convergence with irregular coefficients
In the previous section, we established a first-order weak convergence of the sampled dynamics under the assumption that all coefficients are sufficiently regular so that the PDE associated with the aggregated dynamics has a four-times differentiable solution (see (H.3) and Remark 4.1). In this section, we relax the regularity conditions on the coefficients and quantify the weak convergence of the sampled dynamics based on their smoothness.
The following Hölder spaces will be used in the analysis.
Definition 4.1.
For $l\in(0,\infty)\setminus\mathbb{N}$, let $\mathcal{H}_{T}^{(l)}$ be the space of functions $u:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}$ such that for all integer $r\geq 0$ and multi-index $s$ satisfying $2r+|s|\leq l$, the partial derivatives $\partial_{t}^{r}\partial_{x}^{s}u$ are bounded and Hölder continuous with index $l-[l]$ in $x$ and with index $(l-2r-|s|)/2$ in $t$, i.e.,
$\|u\|_{T}^{(l)}:=$ $\sum_{2r+|s|\leq[l]}\sup_{t,x,a}|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)|$ (4.11)
$+\sum_{2r+|s|=[l]}\sup_{t,a,x\neq x^{\prime}}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)-\partial_{t}^{r}\partial_{x}^{s}u(t,x^{\prime},a)\right|}{\left|x-x^{\prime}\right|^{l-[l]}}$
$+\sum_{0<l-2r-|s|<2}\sup_{t\neq t^{\prime},x,a}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)-\partial_{t}^{r}\partial_{x}^{s}u(t^{\prime},x,a)\right|}{\left|t-t^{\prime}\right|^{(l-2r-|s|)/2}}<+\infty.$
Let $\mathcal{H}^{(l)}$ be the space of all functions $u\in\mathcal{H}_{T}^{(l)}$ that do not depend on $t$ and $a$, and let $\|\cdot\|^{(l)}$ be the corresponding norm.
The following regularity condition is imposed in this section.
H. 4.
Let $l\in(0,1)\cup(1,2)\cup(2,3)$, $b,\sigma\sigma^{\top}\in\mathcal{H}_{T}^{(l)}$, $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}\in\mathcal{H}_{T}^{(l)}$, and $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ is uniformly elliptic.
Under (H.4), standard regularity theory for parabolic PDEs (see, e.g., [24, Chapter 4, Theorem 5.1]) shows that for any $f\in\mathcal{H}^{(l+2)}$ and $t^{\prime}\in[0,T]$, the PDE (4.4) has a unique solution $v_{f}(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v_{f}(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where the constant $C$ depends only on $T,\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$.
Leveraging this PDE regularity, the following theorem quantifies the weak convergence order of the sampled dynamics in terms of the smoothness coefficient $l$.
Theorem 4.2.
Suppose that (H.1), (H.2) and (H.4) hold. Then for $f\in\mathcal{H}^{(l+2)}$, there exists a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for all grids $\mathscr{G}$,
$\sup_{t\in[0,T]}\left|\mathbb{E}\left[f(X^{\mathscr{G}}_{t})\right]-\mathbb{E}\left[f(\widetilde{X}_{t})\right]\right|\leq C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)},$ (4.12)
where the constant $\mathcal{X}(l)$ is defined by
$\mathcal{X}(l)\coloneqq\left\{\begin{aligned} &l/2,&l\in(0,1),\\ &1/(3-l),&l\in(1,2),\\ &1,&l\in(2,3).\end{aligned}\right.$ (4.13)
Proof.
See Section 7.1.2. ∎
Theorem 4.2 establishes the weak convergence order of the sampled dynamics while relaxing the regularity conditions on the coefficients in Theorem 4.1 and allowing for Hölder continuous coefficients. For clarity, we assume that the coefficients are bounded and that the volatility coefficients are uniformly elliptic, enabling us to leverage parabolic regularity theory to ensure the associated PDE admits a sufficiently regular classical solution. The result naturally extends to coefficients with polynomial growth and degenerate volatility, provided that the PDE (4.4) admits a classical solution with polynomial growth.
Here we emphasize that the analysis of policy execution error in Theorem 4.2 fundamentally differs from the convergence rate analysis of the classical piecewise constant control approximation in [16]. This distinction arises because the aggregated dynamics and the sampled dynamics have different coefficients and are thus associated with different PDEs—one governed by aggregated coefficients and the other by random coefficients influenced by execution noise. As a result, the viscosity solution arguments used in [16] are not directly applicable when the PDE (4.4) fails to admit a classical solution.
4.2. Weak error conditioned on execution noise
In this section, we establish high-probability bounds for the policy execution error with respect to the randomness used to sample the actions.
More precisely, let $\mathcal{F}^{\xi}:=\sigma\{\xi_{i}\mid{i\geq 0}\}$, and consider the error term $\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]$, where $f$ is a given test function, and $\mathbb{E}^{W}[\cdot]$ denotes the conditional expectation $\mathbb{E}[\cdot\mid\mathcal{F}^{\xi}]$. Note that the error $\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]$ is a random variable since it depends on the realization of the execution noises $\xi$. We have shown in Section 4.1 that the error converges to zero in expectation as $|\mathscr{G}|\to 0$. In the sequel, we provide a high-probability bound of this error.
H. 5.
For every $a\in A$, the functions $b(\cdot,\cdot,a)$ and $\sigma\sigma^{\top}(\cdot,\cdot,a)$ are of class $C_{p}^{2}([0,T]\times\mathbb{R}^{d})$ for some $p\geq 2$. Moreover, for $\varphi_{1}\in\{b,\sigma\}$, there is a constant $C\geq 0$ such that
$\left|\varphi_{1}(t,x,a)\right|\leq C\left(1+|x|\right),\quad\forall(t,x,a)\in[0,T]\times\mathbb{R}^{d}\times{A},$ (4.14)
and $\sup_{a\in A}\|b(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$ and $\sup_{a\in A}\|\sigma\sigma^{\top}(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$.
Remark 4.2.
Compared with (H.3), (H.5) imposes the additional condition on the uniform boundedness $\sup_{a\in A}\|b(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, and $\sup_{a\in A}\|\sigma\sigma^{\top}(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, which allows for controlling the concentration behavior of the sampled dynamics with respect to the execution noise $\xi$. It is easy to see that (H.4) implies (H.5).
The following theorem shows that the weak convergence rate in expectation, established in Section 4.1, also holds in a high probability sense with respect to the sampling noises. It extends a similar result established in [32] for LQ problems with Gaussian policies to general state dynamics and stochastic policies.
Theorem 4.3.
Suppose (H.1), (H.2), (H.3) and (H.5) hold. Then for the test function $f$ in Theorem 4.1, there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$, such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|>C_{p}\|f\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right]\leq\varrho.$ (4.15)
Assume further that (H.1), (H.2) and (H.4) hold. Then for the test function $f$ in Theorem 4.2, there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|>C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right]\leq\varrho,$ (4.16)
with $\mathcal{X}(l)$ defined by (4.13).
Proof.
See Section 7.1.3. ∎
The proof of Theorem 4.3, provided in Section 7.1.3, is based on establishing a pointwise upper bound for the weak error on each subinterval in terms of the grid size, and subsequently controlling the overall error using concentration inequalities for bounded martingale difference sequences.
4.3. Strong convergence
In this subsection, we quantify the pathwise difference between the sampled dynamics and the aggregated dynamics in terms of the sampling frequency for a fixed realization of the Brownian motion. To this end, we assume the volatility $\sigma$ is uncontrolled, i.e., $\sigma(t,x,a)=\sigma(t,x)$, and impose regularity conditions on the function $\phi$ used for sampling the actions (recall that $a_{t_{i}}=\phi(t_{i},X^{\mathscr{G}}_{t_{i}},\xi_{i})$ as in (3.2)).
H. 6.
  1. (1)
    There exists $C\geq 0$ such that for all $(t_{1},x_{1},a_{1}),(t_{2},x_{2},a_{2})\in[0,T]\times\mathbb{R}^{d}\times A$, $u\in E$,
    $\left|b(t_{1},x_{1},a_{1})-b(t_{2},x_{2},a_{2})\right|+d_{A}(\phi(t_{1},x_{1},u),\phi(t_{2},x_{2},u))\leq C\left(\left|t_{1}-t_{2}\right|^{\frac{1}{2}}+\left|x_{1}-x_{2}\right|+d_{A}(a_{1},a_{2})\right),$ (4.17)
    and for all $x,y\in\mathbb{R}^{d}$ and $t\in[0,T]$,
    $\left|\sigma(t,x)-\sigma(t,y)\right|\leq C\left|x-y\right|.$ (4.18)
  2. (2)
    There exists $(t_{0},x_{0},a_{0})\in[0,T]\times\mathbb{R}^{d}\times A$, such that $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential  [^3] We say a random variable $X$ is sub-exponential if there exists a constant $C>0$ such that $\mathbb{E}[|X|>t]\leq 2e^{-Ct}$ for every $t\geq 0$..
Example 4.5.
If $\bm{\pi}(\cdot|t,x)$ satisfies (4.7) with $\Sigma(t,x)=\Sigma$, then we can write $a=\phi(t,x,\xi)=\mu(t,x)+\Sigma\xi$. In this case, the assumptions for $\phi$ in (H.6) hold. This corresponds to the commonly used Gaussian policies considered in [32] and many others.
Theorem 4.4.
Suppose (H.2) and (H.6) hold. For all $p\geq 2$ and all grids $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ with $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}<\infty$,
$\mathbb{E}\left[\sup_{0\leq t\leq T}\left|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}\right|^{p}\right]\leq C|\mathscr{G}|^{\frac{p}{2}},$ (4.19)
where the constant $C$ depends only on $b,\sigma,\phi,T,p$ and $C_{G}$.
Consequently, if $\mathscr{G}=\{t_{i}\}_{i=0}^{n}$ with $t_{i}=i\frac{T}{n}$, then for every $\varepsilon>0$ and almost every $\omega\in\Omega$, there is an integer $N_{\varepsilon}(\omega)$ such that for every $n\geq N_{\varepsilon}(\omega)$,
$\sup_{0\leq t\leq T}\left|X^{\mathscr{G}}_{t}(\omega)-\widetilde{X}_{t}(\omega)\right|\leq n^{-\frac{1}{2}+\varepsilon}.$ (4.20)
The proof of Theorem 4.4 is given in Section 7.1.4. It involves carefully estimating the Orlicz norm of the pathwise difference between the aggregated and sampled dynamics on each subinterval and subsequently controlling the overall error using concentration inequalities for sub-exponential martingale difference sequences.
Theorem 4.4 indicates that the convergence of the sampled dynamics to the aggregated dynamics can be interpreted as a law of large numbers in time. Specifically, when the volatility is uncontrolled, the $1/2$-order convergence is achieved by increasing the sampling frequency for any fixed realization of the system noise.
Unfortunately, such a strong convergence fails when the volatility is controlled, as shown in the following example.
Example 4.6.
Let $d=1$, $A\subset{\mathbb{R}}$, $b(t,x,a)=0$ and $\sigma(t,x,a)=a$. Assume $\bm{\pi}(\mathrm{d}a|t,x)=\mu(\mathrm{d}a)$ for some $\mu\in\mathcal{P}(A)$ and $Q:=\int_{A}a^{2}\mu(\mathrm{d}a)<+\infty$. Assume that $\{\xi_{i}\}_{i=0}^{\infty}$ are i.i.d. random variables with distribution $\mu$. Let $a_{t_{i}}=\xi_{i}$ for all $i$, then we have
$X^{\mathscr{G}}_{t}-\widetilde{X}_{t}=\int_{0}^{t}\left(a_{\delta(s)}-Q^{1/2}\right)\mathrm{d}W_{s}.$ (4.21)
Note that $W$ is a Brownian motion under the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ with $\mathcal{F}_{t}:=\sigma\{a_{\delta(s)}:s\leq t\}\lor\mathcal{F}^{W}_{t}$. It follows from Itô’s isometry that
$\mathbb{E}\left[|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}|^{2}\right]=\mathbb{E}\left[\int_{0}^{t}(a_{\delta(s)}-Q^{1/2})^{2}\mathrm{d}s\right]=t\mathbb{E}\left[(\xi_{1}-Q^{1/2})^{2}\right],$ (4.22)
which does not converge to zero as $|\mathscr{G}|\rightarrow 0$, except that $\xi_{1}$ is a constant (or equivalently $\mu$ is a Dirac distribution).
5. Applications
This section builds on the convergence results from Section 4 to quantify the accuracy of various estimators used in continuous-time RL algorithms. The proofs of all statements are given in Section 7.2.
5.1. Accuracy of the Monte-Carlo evaluation
Recall that in RL, the agent has the objective function (2.2), and she needs to estimate its value under a given policy $\bm{\pi}$ with data that is generated by executing this policy. This problem is known as the (on-policy) policy evaluation.
The agent can only discretely sample the actions at the pre-specified time grid and observe the state variable at the same time grid $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$. Therefore, the agent can sample the process $X^{\mathscr{G}}_{s}$. For simplicity, we assume that the agent can collect and observe the accumulated running reward $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ [^4] If only the instantaneous reward $r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})$ can be observed at the grid point, then the accumulated reward need to be further approximated as $R_{t_{i}}^{\mathscr{G}}\approx r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})(t_{i}-t_{i-1})$ as in the usual Euler scheme. Under suitable assumptions on the regularity conditions on the function $r$, the order of our analysis does not change. (cf. [18]) Hence, the expected reward collected by the agent is
$J^{\mathscr{G}}\coloneqq\mathbb{E}\left[\int_{0}^{T}r(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\mathrm{d}s+h(X^{\mathscr{G}}_{T})\right]=\mathbb{E}\left[\sum_{i=0}^{n-1}R_{t_{i}}^{\mathscr{G}}+h(X^{\mathscr{G}}_{T})\right].$ (5.1)
In, the expected reward with a continuous-time execution of the policy is given by (2.6):
$\tilde{J}\coloneqq\mathbb{E}\left[\int_{0}^{T}\tilde{r}(s,\widetilde{X}_{s})\mathrm{d}s+h(\widetilde{X}_{T})\right]=\mathbb{E}\left[\int_{0}^{T}\int_{A}r(s,\widetilde{X}_{s},a)\bm{\pi}(\mathrm{d}a|s,\widetilde{X}_{s})\mathrm{d}s+h(\widetilde{X}_{T})\right].$ (5.2)
The following corollary characterizes the difference between (5.2) and (5.1), that is, the bias of using $\int_{0}^{T}r(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\mathrm{d}s+h(X^{\mathscr{G}}_{T})$ as an estimate for $\tilde{J}$.
Proposition 5.1.
Assume (H.1) and (H.2) hold.
  1. (1)
    If $r\in C^{2,0}_{p},\tilde{r}\in C^{2}_{p}$ and (H.3) holds with the terminal function $f=h,\tilde{r}(t,\cdot),0\leq t\leq T$, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$ such that
    $\left|J^{\mathscr{G}}-\tilde{J}\right|\leq C_{p}\left(\|h\|_{C_{p}^{4}}+\|r\|_{C_{p}^{2,0}}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|_{C_{p}^{4}}+\|\tilde{r}\|_{C_{p}^{2}}\right)|\mathscr{G}|.$ (5.3)
  2. (2)
[…]
    $\left|J^{\mathscr{G}}-\tilde{J}\right|\leq C\left(\|h\|^{(l+2)}+\|r\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|^{(l+2)}+\|\tilde{r}\|_{T}^{(l)}\right)|\mathscr{G}|^{\mathcal{X}(l)},$ (5.4)
    with $\mathcal{X}(l)$ defined by (4.13).
Proposition 5.1 characterizes the (unconditional) bias of the estimated value of a given policy by discrete sampling and piecewise constant execution. Given this result, one may understand the required simulation budget in order to obtain an estimate with desired accuracy. Next, we use the notion of the conditional weak error to highlight that the required simulation budget can be refined to those conditioned on the policy sampler. More precisely, consider the value conditional on $\xi$,
$J^{\mathscr{G}}(\xi)\coloneqq\mathbb{E}^{W}\left[\int_{0}^{T}r(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\mathrm{d}s+h(X^{\mathscr{G}}_{T})\right].$ (5.5)
Combining the proof of Proposition 5.1 and Theorem 4.3, we obtain the following result for the large deviation bound for the value function.
Proposition 5.2.
Assume (H.1) and (H.2) hold. If (H.5) and the conditions in Proposition 5.1(1) hold, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$ such that for every $\varrho\in(0,1)$,
$\mathbb{P}\left[|J^{\mathscr{G}}(\xi)-\tilde{J}|>C_{p}\left(\|h\|_{C_{p}^{4}}+\|r\|_{C_{p}^{2,0}}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|_{C_{p}^{4}}+\|\tilde{r}\|_{C_{p}^{2}}\right)\sqrt{|\mathscr{G}|}(1+\sqrt{\ln(2/\varrho)})\right]\leq\varrho.$ (5.6)
If conditions in Proposition 5.1(2) holds, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for every $\varrho\in(0,1)$,
$\mathbb{P}\left[|J^{\mathscr{G}}(\xi)-\tilde{J}|>C\left(\|h\|^{(l+2)}+\|r\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|^{(l+2)}+\|\tilde{r}\|_{T}^{(l)}\right)|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}(1+\sqrt{\ln(2/\varrho)})\right]\leq\varrho$ (5.7)
with $\mathcal{X}(l)$ defined by (4.13).
5.2. Accuracy of Temporal-Difference learning
The temporal difference (TD) plays a crucial role in many RL algorithms as the policy evaluation methods. In contrast to the Monte-Carlo evaluation, TD learning often aims to find an approximation of the value function associated with a given policy $\bm{\pi}$, defined as
$\tilde{J}(t,x;\bm{\pi})=\mathbb{E}\left[\int_{t}^{T}\tilde{r}(s,\widetilde{X}_{s})\mathrm{d}s+h(\widetilde{X}_{T})\Big{|}\widetilde{X}_{t}=x\right],$ (5.8)
within a suitable parametric family of functions, and it makes use of the increment of the value function process.
In particular, suppose $V^{\theta}$ is a family of parameterized functions of $(t,x)$ with $\theta$ a finite-dimensional parameter. TD learning concerns the increment in the form “$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” and the instantaneous reward “$r(t,\widetilde{X}_{t})\mathrm{d}t$”. In the continuous-time framework, these two increments have different natures. “$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” is merely an informal notation that shall be combined with a suitable integrand as in the definition of Itô’s integral, and the latter “$r(t,\widetilde{X}_{t})\mathrm{d}t$” is the usual Lebesgue integral with respect to time. In theory, [19] identify the continuous-time counterpart of TD learning as solving the martingale orthogonality condition. More precisely, if $\theta^{*}$ is the root to
$\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left[\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right]=0,$ (5.9)
for all admissible test functions $S(t,x)$, then the $V^{\theta^{*}}(t,x)=\tilde{J}(t,x;\bm{\pi})$ is the desired value function associated with a given policy $\bm{\pi}$. Therefore, various RL algorithms are based on using stochastic approximation (cf. [7]) to find the root $\theta^{*}$ to the martingale orthogonality condition (5.9). The most popular choice is the so-called “TD(0)” algorithm, which corresponds to taking $S(t,x)=\frac{\partial V^{\theta}}{\partial\theta}(t,x)$.
For a given test function $S$ and parameter $\theta$, applying the stochastic approximation algorithm means approximating this expectation with random samples, i.e., an estimator of this expectation. The typical convergence conditions of the stochastic approximation algorithm are threefold [^5] Besides these conditions on the accuracy, the convergence also depends on the landscape of (5.9) as a function of $\theta$. However, this is jointly determined by the parametrization $\theta\mapsto V^{\theta}$ and the properties of the aggregated dynamics, which is not affected by time discretization.: first, the estimator has diminishing bias; second, the estimator has non-exploding variances; third, the learning rates should be suitably chosen according to the bias and variance of the estimator. Therefore, it is crucial to understand the magnitude of bias and variance and how they depend on the grid size.
In continuous time, however, due to the discretely sampled actions and the nature of the stochastic integral inside (5.9), the agent only has access to the sampled dynamics and uses the discrete sum (we drop the superscript $\theta$ in the following to illustrate the results for a generic function $V$)
$S\circ V:=\sum_{i=0}^{n-1}S(t_{i},X^{\mathscr{G}}_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R_{t_{i}}^{\mathscr{G}}\right),$
with the observed accumulated reward $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ as an approximation for the expectation in (5.9). We characterize the bias and variance of such an approximation in Proposition 5.3.
Proposition 5.3.
Suppose (H.1) and (H.2) hold. Assume that $V\in C_{p}^{2}$. Denote
$g(t,x,a)=\partial_{t}V(t,x)+b^{\top}(t,x,a)\partial_{x}V(t,x)+\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t,x,a)\textrm{Hess}_{x}V(t,x)\right)+r(t,x,a)=:\mathcal{L}^{a}V(t,x).$ (5.10)
For the variance of $S\circ V$, if $S\in C_{p}^{0}$ and $r\in C^{0,0}_{p}$, then
$\mathrm{Var}\left(S\circ V\right)\leq C_{p}\|S\|_{C_{p}^{0}}^{2}(\|V\|_{C_{p}^{2}}+\|r\|_{C_{p}^{0,0}})^{2},$ (5.11)
where $C_{p}$ is a constant depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$.
For the bias of $S\circ V$, we have the following holds.
  1. (1)
    If $g\in C_{p}^{0,0},S\in C^{2}_{p}$ and $gS$ satisfies the condition (1) in Proposition 5.1 as the running term, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$ such that
    $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left[\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right]\right|$ (5.12)
    $\leq C_{p}\left(\|S\|_{C_{p}^{2}}\|g\|_{C_{p}^{0,0}}+\|Sg\|_{C_{p}^{2,0}}+\|\widetilde{Sg}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|_{C_{p}^{4}}\right)|\mathscr{G}|,$
    where $\widetilde{Sg}(t,x):=\int_{A}Sg(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$.
  2. (2)
    If $g\in\mathcal{H}_{T}^{(0)},S\in\mathcal{H}_{T}^{(l)}$ and $gS$ satisfies the condition (2) in Proposition 5.1 as the running term, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that
    $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left[\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right]\right|$ (5.13)
    $\leq C\left(\|S\|_{T}^{(l)}\|g\|_{T}^{(0)}+\|Sg\|^{(l)}_{T}+\|\widetilde{Sg}\|^{(l)}_{T}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|^{(l+2)}\right)|\mathscr{G}|^{\mathcal{X}(l)}.$
Proposition 5.3 implies that we can obtain a more accurate stochastic approximation with a refined grid. In addition, the variance of such approximate is uniformly bounded in the grid size. It means that with a suitable schedule to refine the grid, we may expect the typical convergence conditions in the stochastic approximation algorithm will be satisfied.
5.3. Accuracy of policy gradient
The policy gradient is one of the most commonly used methods in policy optimization in RL. The well-celebrated (model-free) policy gradient representation (e.g., [31]) gives an unbiased estimator for the gradient of the value function with respect to certain parameters in the policy. For example, the stochastic policy $\bm{\pi}^{\psi}$ is parameterized by $\psi$, a finite-dimensional parameter, and it admits a density function with respect to the Lebesgue measure. We denote the density function by $H^{\psi}(t,x,a)=\frac{\bm{\pi}^{\psi}(\mathrm{d}a|t,x)}{\mu(\mathrm{d}a)}$. Then we know the value function under the policy $\bm{\pi}^{\psi}$ (defined in (5.8)) will depend on $\psi$ implicitly, denoted by $\tilde{J}(t,x;\bm{\pi}^{\psi})$. The policy gradient concerns about the derivative of $\tilde{J}(t,x;\bm{\pi}^{\psi})$ with respect to $\psi$, i.e., $\frac{\partial\tilde{J}(t,x;\bm{\pi}^{\psi})}{\partial\psi}$. For simplicity, we focus on the initial time-state gradient and suppress the time-state pair’s dependence.
A heuristic derivation in [19] shows that the gradient $\frac{\partial\tilde{J}(\bm{\pi}^{\psi})}{\partial\psi}$ can be represented by
$\mathbb{E}\left[\int_{0}^{T}\frac{\partial\log H^{\psi}}{\partial\psi}(a_{t}|t,X_{t})\left[\mathrm{d}\tilde{J}(t,X_{t};\bm{\pi}^{\psi})+r(t,X_{t},a_{t})\mathrm{d}t\right]\right]$
$=\mathbb{E}^{W}\left[\int_{0}^{T}\int_{A}\frac{\partial\log H^{\psi}}{\partial\psi}(a|t,\widetilde{X}_{t})\mathcal{L}^{a}\tilde{J}(t,\widetilde{X}_{t};\bm{\pi}^{\psi})\bm{\pi}(\mathrm{d}a|t,\widetilde{X}_{t})\mathrm{d}t\right].$ (5.14)
The expectation (5.14) can be approximated by the forward-Euler-type of finite sums using sampled dynamics, i.e.,
$\sum_{i=0}^{n-1}\frac{\partial\log H^{\psi}}{\partial\psi}(a_{t_{i}}|t_{i},X^{\mathscr{G}}_{t_{i}})\left[\tilde{J}(t_{i+1},X^{\mathscr{G}}_{t_{i+1}};\bm{\pi}^{\psi})-\tilde{J}(t_{i},X^{\mathscr{G}}_{t_{i}};\bm{\pi}^{\psi})+R_{t_{i}}^{\mathscr{G}}\right].$
Naturally, the convergence of gradient-based algorithms depends on whether we have an accurate gradient estimate. To analyze the accuracy of the policy gradient estimator, we consider the following general expression:
$S\circ V:=\sum_{i=0}^{n-1}S(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R_{t_{i}}^{\mathscr{G}}\right),$ (5.15)
where $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ is the observed accumulated reward over $[t_{i},t_{i+1}]$.
The form (5.15) also appears in the continuous-time q-learning algorithm proposed in [20], where $S$ is the derivative of the q-function that needs to be learned. The difference between (5.15) and (5.9) is that now the function $S$ also depends on the sampled action $a_{t_{i}}$. Thus, (5.15) is not a straightforward substitution and discretization of any integration along the aggregated dynamics and can only be well-defined using the sampled dynamics. Note that given $X^{\mathscr{G}}_{t_{i}}$, the value of $X^{\mathscr{G}}_{t_{i+1}}$ not only depends on the environmental noise (the underlying Brownian motions) but also depends on $a_{t_{i}}$. The subtle difference in this measurability condition will affect the limiting value of (5.15).
We characterize the limit of a general estimator in the form (5.15) and quantify the magnitude of its bias and variance in the next proposition. The proof is similar to that of Proposition 5.3, so we omit it.
Proposition 5.4.
Suppose (H.1) and (H.2) hold. Assume $V\in C_{p}^{2}$. Define $g$ by (5.10). For the variance of $S\circ V$, if $S\in C_{p}^{0,0}$ and $r\in C_{p}^{0,0}$, then
$\mathrm{Var}\left(S\circ V\right)\leq C_{p}\|S\|_{C_{p}^{0,0}}^{2}(\|V\|_{C_{p}^{2}}+\|r\|_{C_{p}^{0,0}})^{2},$ (5.16)
where $C_{p}$ is a constant depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$.
For the bias of $S\circ V$, we have the following holds.
  1. (1)
    If $g\in C_{p}^{0,0},S\in C^{2,0}_{p}$ and $gS$ satisfies the condition (1) in Proposition 5.1 as the running term, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$ such that
    $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}\widetilde{Sg}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right|$ (5.17)
    $\leq C_{p}\left(\|S\|_{C_{p}^{2,0}}\|g\|_{C_{p}^{0,0}}+\|Sg\|_{C_{p}^{2,0}}+\|\widetilde{Sg}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|_{C_{p}^{4}}\right)|\mathscr{G}|,$
    where $\widetilde{Sg}(t,x):=\int_{A}S(t,x,a)g(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$.
    In particular, if the conditions in [19] are satisfied so that the policy gradient has the representation (5.14), then by setting $V(\cdot,\cdot)=\tilde{J}(\cdot,\cdot;\bm{\pi}^{\psi})$, and $S(t,x,a)=\frac{\partial\log H^{\psi}}{\partial\psi}(a|t,x)$, we have
    $\left|\mathbb{E}\left[S\circ V\right]-\frac{\partial\tilde{J}(\bm{\pi}^{\psi})}{\partial\psi}\right|\leq C_{p}\left(\|S\|_{C_{p}^{2,0}}\|g\|_{C_{p}^{0,0}}+\|Sg\|_{C_{p}^{2,0}}+\|\widetilde{Sg}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|_{C_{p}^{4}}\right)|\mathscr{G}|.$
  2. (2)
    If $g\in\mathcal{H}_{T}^{(0)},S\in\mathcal{H}_{T}^{(l)}$ and $gS$ satisfies the condition (2) in Proposition 5.1 as the running term, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that
    $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}\widetilde{Sg}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right|$ (5.18)
    $\leq C\left(\|S\|_{T}^{(l)}\|g\|_{T}^{(0)}+\|Sg\|^{(l)}_{T}+\|\widetilde{Sg}\|^{(l)}_{T}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|^{(l+2)}\right)|\mathscr{G}|^{\mathcal{X}(l)}.$
    In particular, if the conditions in [19] are satisfied so that the policy gradient has the representation (5.14), then by setting $V(\cdot,\cdot)=\tilde{J}(\cdot,\cdot;\bm{\pi}^{\psi})$, and $S(t,x,a)=\frac{\partial\log H^{\psi}}{\partial\psi}(a|t,x)$, we have
    $\left|\mathbb{E}\left[S\circ V\right]-\frac{\partial\tilde{J}(\bm{\pi}^{\psi})}{\partial\psi}\right|\leq C\left(\|S\|_{T}^{(l)}\|g\|_{T}^{(0)}+\|Sg\|^{(l)}_{T}+\|\widetilde{Sg}\|^{(l)}_{T}+\sup_{t\in[0,T]}\|\widetilde{Sg}(t,\cdot)\|^{(l+2)}\right)|\mathscr{G}|^{\mathcal{X}(l)}.$
Proposition 5.4 implies that with a refined grid, we can obtain a more accurate estimate for the policy gradient. In addition, the variance of such an estimator is uniformly bounded in the grid size. These properties are crucial for adjusting the sampling frequency to ensure the convergence of the overall policy gradient algorithm. [^6] The typical conditions of convergence of the stochastic gradient descent algorithm also require $\frac{\partial\tilde{J}(\bm{\pi}^{\psi})}{\partial\psi}$ to satisfy certain conditions as a function of $\psi$, but it is not affected by time discretization. Note that in a typical actor-critic algorithm, one does not have an exact estimate of $\tilde{J}(\cdot,\cdot;\bm{\pi}^{\psi})$, and can only obtain its approximation $V^{\theta}\approx\tilde{J}(\cdot,\cdot;\bm{\pi}^{\psi})$ via policy evaluation. Therefore, the convergence of a general actor-critic algorithm is more complex and open for future research.
5.4. Accuracy of risk-sensitive q-learning
So far, all the discussions have been restricted to the additive functional, which is often regarded as a risk-neutral problem. In continuous-time risk-sensitive problems (e.g., [6]), one encounters the objective function in the exponential form:
$\frac{1}{\epsilon}\log\mathbb{E}\left[\exp\left\{\epsilon\left[\int_{0}^{T}r(s,X_{s},a_{s})\mathrm{d}s+h(X_{T})\right]\right\}\right].$
[17] establishes the risk-sensitive q-learning theory and reveals that it only differs from the non-risk-sensitive counterpart by a penalty on the quadratic variation of the value function. Hence, the RL algorithm for the non-risk-sensitive problems can be easily modified to account for risk sensitivity by adding an integral with respect to a suitable quadratic variation process, which can be approximated by the sampled dynamics as
$S\Box V:=\sum_{i=0}^{n-1}S(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R_{t_{i}}^{\mathscr{G}}\right)^{2}.$
Proposition 5.5.
Suppose (H.1) and (H.2) hold. Assume $V\in C_{p}^{2}$. Define $g_{1}$ by (5.10) and $g_{2}(t,x,a)=\left|\sigma^{\top}\partial_{x}V(t,x,a)\right|^{2}$. For the variance of $S\circ V$, if $S,g_{1}\in C_{p}^{0,0}$, then
$\mathrm{Var}\left(S\Box V\right)\leq C_{p}\|S\|_{C_{p}^{0,0}}^{2}\left(\|g_{1}\|_{C_{p}^{0,0}}^{2}+\|V\|_{C_{p}^{2}}^{2}\right)^{2},$ (5.19)
where $C_{p}$ is a constant depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$.
For the bias of $S\Box V$, we have the following holds.
  1. (1)
    If $S,g_{1},\partial_{x}V,\sigma\in C^{2,0}_{p},g_{2}\in C_{p}^{0,0}$ and $Sg_{2}$ satisfies the condition (1) on Proposition 5.1 as the running term, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},T,p$ such that
    $\left|\mathbb{E}[S\Box V]-\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg_{2}}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right|$ (5.20)
    $\leq C_{p}\left(\|S\|_{C_{p}^{2,0}}(\|g_{1}\|_{C_{p}^{2,0}}^{2}+\|g_{2}\|_{C_{p}^{0,0}})+\|g_{1}\|_{C_{p}^{2,0}}\|V\|_{C_{p}^{2}}+\|Sg_{2}\|_{C_{p}^{2,0}}+\|\widetilde{Sg_{2}}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|_{C_{p}^{4}}\right)|\mathscr{G}|.$
  2. (2)
    If $S,g_{1},\partial_{x}V,\sigma\in\mathcal{H}^{(l)}_{T},g_{2}\in\mathcal{H}_{T}^{(0)}$ and $Sg_{2}$ satisfies the condition (2) on Proposition 5.1 as the running term, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that
    $\left|\mathbb{E}[S\Box V]-\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg_{2}}(t,\widetilde{X}_{t})\mathrm{d}t\right]\right|$ (5.21)
    $\leq C\left(\|S\|_{T}^{(l)}((\|g_{1}\|_{T}^{(l)})^{2}+\|g_{2}\|_{T}^{(0)})+\|g_{1}\|_{T}^{(l)}\|V\|_{T}^{(l)}+\|Sg_{2}\|_{T}^{(l)}+\|\widetilde{Sg_{2}}\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|^{(l+2)}\right)|\mathscr{G}|^{\mathcal{X}(l)}.$
In particular, if $S$ does not depend on $a$, then
$\mathbb{E}\left[\int_{0}^{T}\int_{A}\widetilde{Sg_{2}}(t,\widetilde{X}_{t})\mathrm{d}t\right]=\mathbb{E}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\mathrm{d}\left<V(\cdot,\widetilde{X}_{\cdot})\right>_{t}\right],$
where $t\mapsto\left<V(\cdot,\widetilde{X}_{\cdot})\right>_{t}$ is the quadratic variation of the process $t\mapsto V(t,\widetilde{X}_{t})$.
6. Conclusion
This paper addresses a fundamental problem in the continuous-time RL framework proposed in [36] by answering the relation between the aggregated dynamics (referred to as the “exploratory dynamics” therein) and the actual sampling process. In terms of SDE theory, we provide a rigorous treatment of the “law-of-large-numbers” heuristics used in [36] to demonstrate how the original dynamics lead to the aggregated dynamics. Our analysis is particularly crucial for understanding how discrete sampling relates to the continuous-time RL problems. We establish the convergence and convergence rate of discretely sampled stochastic policies, i.e., the sampled dynamics, to the aggregated dynamics in different senses. Furthermore, we illustrate the bias and variance of the estimate for several important quantities in analyzing continuous-time RL algorithms when using sampled dynamics, including values, temporal differences, policy gradients, and quadratic variations. These bias and variance analyses are critical for the convergence of stochastic approximation algorithms, and they provide guidance on properly tuning the interaction frequency to foster faster convergence of RL algorithms.
The results of this paper also stimulate more interesting open questions for future research. For example, other sampling methods that could achieve higher accuracy in the RL context, any exact simulation methods for controlled SDEs under stochastic policies if the coefficients are all known/unknown, and a full account for the convergence of the gradient-based continuous-time RL algorithm.
7. Proofs of Main Results
7.1. Proofs of Theorems 4.1, 4.2, 4.3, and 4.4
7.1.1. Proof of Theorem 4.1
Proof.
Without loss of generality, assume $t^{\prime}=t_{j}$, where $1\leq j\leq n$. Consider the following PDE
$\frac{\partial}{\partial t}v(t,x)+\mathcal{L}v(t,x)=0,\quad t\in[0,t_{j}]\times{\mathbb{R}}^{d};\quad v(t_{j},x)=f(x),\quad x\in{\mathbb{R}}^{d}.$ (7.1)
Let $v_{f}(t,x;t_{j})$ be the solution of (7.1). For simplicity, we dorp the $f$ and $t_{j}$ and use the notation $v(t,x)$. By (H. 3), $v(\cdot,\cdot)$ is smooth and $\|v(\cdot,\cdot)\|_{C_{p}^{4}}<+\infty$.
Noting that if $v$ is the solution of (7.1), then by the Feynman-Kac formula, $\mathbb{E}[f(\widetilde{X}_{t_{j}})]=v(0,x_{0})$. With this relation, we have
$\mathbb{E}\left[f(X^{\mathscr{G}}_{t_{j}})-f(\widetilde{X}_{t_{j}})\right]$ $=\mathbb{E}\left[v(t_{j},X^{\mathscr{G}}_{t_{j}})-v(0,x_{0})\right]$ (7.2)
$=\mathbb{E}\left[\sum_{i=1}^{j}\left(v(t_{i},X^{\mathscr{G}}_{t_{i}})-v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right]=:\sum_{i=1}^{j}\mathbb{E}\left[e_{i}\right].$
It suffices to estimate every $\mathbb{E}[e_{i}]$. For simplicity, denote $\partial_{t}:=\partial/\partial t$ and $\partial_{x}:=\partial/\partial x$. It follows from Itô’s lemma (see the dynamic (3.3) of $X^{\mathscr{G}}$) that
$\mathbb{E}\left[e_{i}\right]=$ $\mathbb{E}\left[v(t_{i},X^{\mathscr{G}}_{t_{i}})-v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right]$
$=$ $\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})+b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}v(s,X^{\mathscr{G}}_{s})\right)\mathrm{d}s$
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})\right)\mathrm{d}s\bigg{]},$
with $a_{t_{i-1}}\sim\bm{\pi}(\cdot|t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$, where in the second equality, we use the fact that the Itô integral is a martingale, given that $\sigma$ and $\partial_{x}v$ have polynomial growth and the moments of $X^{\mathscr{G}}_{s}$ are finite (by Lemma 3.1). Noting that $\xi_{i-1}$ is independent of $X^{\mathscr{G}}_{t_{i-1}}$, we have
$\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$ (7.3)
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
$=$ $\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+(\tilde{b}^{\bm{\pi}})^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}})^{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}=0,$
where the first equality follows form taking conditional expectation $\mathbb{E}[\cdot|X^{\mathscr{G}}_{t_{i-1}}]$ and the second equality follows from (7.1). Therefore,
$\mathbb{E}\left[e_{i}\right]$ $=\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})-\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$ (7.4)
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}v(s,X^{\mathscr{G}}_{s})-\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
For $h=\partial_{t}v,\partial_{x}v,\partial_{x}^{2}v,b(\cdot,\cdot,a_{t_{i-1}})\ \text{or}\ \frac{1}{2}\sigma\sigma^{\top}(\cdot,\cdot,a_{t_{i-1}})$, we again use Itô’s lemma to obtain
$h$ $(s,X^{\mathscr{G}}_{s})-h(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (7.5)
$=\int_{t_{i-1}}^{s}\left(\partial_{t}h(u,X^{\mathscr{G}}_{u})+b^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}h(u,X^{\mathscr{G}}_{u})\right)\mathrm{d}u$
$\quad+\int_{t_{i-1}}^{s}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}^{2}h(u,X^{\mathscr{G}}_{u})\right)\mathrm{d}u+\int_{t_{i-1}}^{s}(\partial_{x}h(u,X^{\mathscr{G}}_{u}))^{\top}\sigma(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\mathrm{d}W_{u}.$
By (H.3), Lemma 3.1 and due to $b,\sigma\sigma^{\top}\in C_{p}^{2,0}$, the integrands on the right hand side of the above equation have finite moments. Therefore, if we substitute (7.5) to the first and last two terms of (7.4), then every time integral term is bounded by $C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}$ for constant $C_{p}>0$ depends only on $b,\sigma,\bm{\pi},T,p$, and the Itô integral terms vanish. For the second and third terms of (7.4), let $h_{1}=b(\cdot,\cdot,a_{t_{i-1}})\ \text{or}\ \frac{1}{2}\sigma\sigma^{\top}(\cdot,\cdot,a_{t_{i-1}})$ and $h_{2}=\partial_{x}v$. We have
$\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}h_{1}(s,X^{\mathscr{G}}_{s})\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right]$ (7.6)
$=$ $\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right]$
$+\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left(h_{1}(s,X^{\mathscr{G}}_{s})-h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right].$
For the first term of (7.6), we again use the Itô’s Lemma to obtain the upper bound $C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}$. Note that by Hölder’s inequality and (7.5), the second term of (7.6) is bounded by
$\sqrt{\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left|h_{1}(s,X^{\mathscr{G}}_{s})-h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right|^{2}\mathrm{d}s\right]\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left|h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right|^{2}\mathrm{d}s\right]}\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}.$
Thus $\mathbb{E}\left[e_{i}\right]\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$ where $C_{p}$ is a constant depending only on $b,\sigma,\bm{\pi},T,p$. The desired result follows from $\sum_{i=1}^{j}(t_{i}-t_{i-1})^{2}\leq T|\mathscr{G}|$. ∎
7.1.2. Proof of Theorem 4.2
The following lemma is similar to [23, Lemma 14.1.6]. One can combine the method in Theorem 4.1 to prove it, so we omit the proof here.
Lemma 7.1.
Suppose (H.1) holds, $l\in(0,1)\cup(1,2)\cup(2,3)$ and $g\in\mathcal{H}_{T}^{(l)}$. Then there exist a constant $C$, such that for all $s\in[0,T]$,
$\left|\mathbb{E}\left[g(s,X^{\mathscr{G}}_{s},a_{\delta(s)})-g(\delta(s),X^{\mathscr{G}}_{\delta(s)},a_{\delta(s)})\right]\right|\leq C\|g\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)},$ (7.7)
where $\mathcal{X}(l)$ satisfies (4.13).
Based on Lemma 7.1, we prove Theorem 4.2.
Proof.
It suffices to prove $|\mathbb{E}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}[f(\widetilde{X}_{t^{\prime}})]|\leq C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}$ for any $t^{\prime}\in[0,T]$. Under (H. 4), by [24, Chapter 4, Theorem 5.1], for any $f\in\mathcal{H}^{(l+2)}$, the PDE (4.4) has a unique solution $v(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where $C$ is a constant depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}$. Note that $\mathcal{H}_{T}^{(l)}\subset\mathcal{H}_{s}^{(l)}$ for every $0\leq t^{\prime}\leq T$. Without loss of generality, suppose that $t^{\prime}=T$. For simplicity, we drop the notation $t^{\prime}$ and use $v$ to represent the solution of (4.4) with $t^{\prime}=T$. Recall (7.2),
$\mathbb{E}\left[f(X^{\mathscr{G}}_{T})-f(\widetilde{X}_{T})\right]$ $=\mathbb{E}\left[v(T,X^{\mathscr{G}}_{T})-v(0,x_{0})\right]=\sum_{i=1}^{n}\mathbb{E}\left[e_{i}\right].$
Note that (7.3) still holds in this situation. Therefore, we have
$\mathbb{E}$ $\left[e_{i}\right]=\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})-\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$ (7.8)
$+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}v(s,X^{\mathscr{G}}_{s})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
$+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
Note that $\partial_{t}v,b^{\top}\partial_{x}v$ and $\frac{1}{2}\operatorname*{Tr}(\sigma\sigma^{\top}\partial_{x}^{2}v)$ belong to $\mathcal{H}_{T}^{(l)}$. By Lemma 7.1, the three terms on the right hand side of above equation are all bounded by $C\|v\|^{(l+2)}_{T}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1})$ with a constant $C$ depending only on $b,\sigma,T$. As a result, we have
$\bigg{|}\mathbb{E}\left[f(X^{\mathscr{G}}_{T})-f(\widetilde{X}_{T})\right]\bigg{|}\leq\sum_{i=1}^{n}\left|\mathbb{E}\left[e_{i}\right]\right|\leq\sum_{i=1}^{n}C\|v\|^{(l+2)}_{T}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1})\leq CT\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.9)
This proves the desired result. ∎
7.1.3. Proof of Theorem 4.3
Recall $\mathcal{F}^{\xi}:=\sigma\{\xi_{i}\mid{i\geq 0}\}$ as defined in Section 4.2. To facilitate the proof, we define $\mathcal{F}^{\xi}_{i}:=\sigma\{\xi_{j}\mid j\leq i-1\}$ for all $i\geq 0$, and define the filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with ~$\widetilde{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W}\vee\mathcal{F}^{\xi}$. Note that $\{W_{s}\}_{s\geq 0}$ is still a Brownian motion on $(\Omega,\mathcal{F},{\mathbb{P}})$ with filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ and we have $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}\big{[}\cdot\big{|}\widetilde{\mathcal{F}}_{0}\big{]}$.
The following lemma shows that $X^{\mathscr{G}}$ has finite moments conditional on $\mathcal{F}^{\xi}$. The proof is given in Appendix A.
Lemma 7.2.
Under condition (4.14) of (H.5), for $p\geq 2$, there is a constant $C_{p}$ such that
$\mathbb{E}^{W}\left[|X^{\mathscr{G}}_{s}|^{p}\right]\leq C_{p}(1+|x_{0}|^{p})e^{C_{p}s},\quad s\in[0,T].$ (7.10)
Based on Lemma 7.2, we prove Theorem 4.3.
Proof of Theorem 4.3.
Without loss of generality, assume that $t=t_{j}$ for some $1\leq j\leq n$. To prove (4.15), consider the following PDE:
$\frac{\partial}{\partial t}v(t,x)+\mathcal{L}v(t,x)=0,\quad(t,x)\in[0,t_{j}]\times{\mathbb{R}}^{d};\quad v(t_{j},x)=f(x),\quad x\in{\mathbb{R}}^{d}.$ (7.11)
Under the conditions of this theorem, the solution $v$ of (7.11) is smooth and bounded. By the Feynman-Kac formula, $\mathbb{E}^{W}[f(\widetilde{X}_{t_{j}})]=v(0,x_{0})$. With this relation, we have
$\mathbb{E}^{W}\left[f(X^{\mathscr{G}}_{t_{j}})-f(\widetilde{X}_{t_{j}})\right]$ $=\mathbb{E}^{W}\left[v(t_{j},X^{\mathscr{G}}_{t_{j}})-v(0,x_{0})\right]$
$=\mathbb{E}^{W}\left[\sum_{i=1}^{j}\left(v(t_{i},X^{\mathscr{G}}_{t_{i}})-v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right]=:\sum_{i=1}^{j}\mathbb{E}^{W}\left[e_{i}\right].$
As in the proof of Theorem 4.1, noting that the Itô integral in (7.5) is mean zero under $\mathbb{P}^{W}$, we can decompose $\mathbb{E}^{W}\left[e_{i}\right]$ into two terms and obtain
$\mathbb{E}^{W}\left[f(X^{\mathscr{G}}_{T})\right]-\mathbb{E}^{W}\left[f(\widetilde{X}_{T})\right]=\sum_{i=1}^{j}A_{i}+\sum_{i=1}^{j}B_{i},$ (7.12)
where $A_{i}$ is defined by
$A_{i}:=$ $\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
and $B_{i}$ is defined by
$B_{i}$ $\coloneqq\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})-\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$ (7.13)
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}v(s,X^{\mathscr{G}}_{s})-\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right)\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
We again use the Itô’s lemma to analysis every difference term of (7.13). Note that every term in $B_{j}$ can be written as double integrals with respect to time as in the proof of Theorem 4.1, and by conditions (H.5) and (H.3), the integand is bounded by $C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{s}|^{m_{p}})$ for some $m_{p}$ and $C_{p}$ that depend on $p$. Applying Lemma 7.2, we obtain
$|B_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$
where $C_{p}$ is a constant depending on $b,\sigma,\bm{\pi},T,p$. As a result,
$\sum_{i=1}^{j}|B_{i}|\leq\sum_{i=1}^{j}C_{P}\|v\|_{C_{p}^{4}}|\mathscr{G}|(t_{i}-t_{i-1})\leq TC_{p}\|v\|_{C_{p}^{4}}|\mathscr{G}|.$ (7.14)
As for $A_{i}$, note that $\{(A_{i},\mathcal{F}^{\xi}_{i})\}_{i}$ is a martingale difference sequence. In fact, denote
$S_{i}:=\bigg{(}\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ $+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (7.15)
$+\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\bigg{)}(t_{i}-t_{i-1}).$
Observing that $S_{i}$ is independent of $\xi_{i},\xi_{i+1},\dots$, we have $A_{i}=\mathbb{E}^{W}\left[S_{i}\right]=\mathbb{E}[S_{i}\mid\mathcal{F}^{\xi}_{i}]$ is $\mathcal{F}^{\xi}_{i}$-measurable and for $i\geq 2$,
$\mathbb{E}[A_{i}\mid\mathcal{F}^{\xi}_{i-1}]$ $=\mathbb{E}\left[\mathbb{E}^{W}\left[S_{i}\right]\big{|}\xi_{0},\dots,\xi_{i-2}\right]=\mathbb{E}\left[\mathbb{E}\left[S_{i}\big{|}\xi_{0},\dots,\xi_{i-2},\xi_{i-1}\right]\big{|}\xi_{0},\dots,\xi_{i-2}\right]$ (7.16)
$=\mathbb{E}\left[\mathbb{E}\left[S_{i}\big{|}X_{t_{i-1}},\xi_{0},\dots,\xi_{i-2}\right]\big{|}\xi_{0},\dots,\xi_{i-2}\right]=0,$
where the last equality follows from the same result in (7.3).
By conditions (H.3) and (H.5), there is a constant $C_{p}$ and an integer $m_{p}$ such that $|S_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{t_{i-1}}|^{m_{p}})$. Again using Lemma 7.2, we obtain
$|A_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1}).$ (7.17)
By [35, Corollary 2.20], we have
$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq t\bigg{)}\leq 2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right).$ (7.18)
Let $\varrho=2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right)$, we obtain
$\mathbb{P}\bigg{(}\bigg{|}\sum_{j=1}^{i}A_{i}\bigg{|}\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}\leq\varrho.$ (7.19)
It follows from (7.14) that
$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t^{\prime}})]\right|>(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right)$
$\quad\leq\mathbb{P}\bigg{(}\bigg{|}\sum_{j=1}^{i}A_{i}\bigg{|}\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\right)$
$\quad\leq\varrho+0=\varrho.$
The desired result follows from $\|v\|_{C_{p}^{4}}\leq C\|f\|_{C_{p}^{4}}$.
For the proof of (4.16), we can start from (7.12). Note that $\{A_{i}\}_{i}$ is still a martingale difference sequence. By (H.4), we have
$|A_{i}|\leq C\|v\|_{t_{j}}^{(l+2)}(t_{i}-t_{i-1})$ (7.20)
It suffices to analyze (7.13). Applying Lemmas 7.2 and 7.1, one can use a similar method in the proof of Theorem 4.2 to obtain
$|B_{i}|\leq C\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1}).$ (7.21)
As a result,
$\sum_{i=1}^{j}|B_{i}|\leq CT\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.22)
By [35, Corollary 2.20], we have
$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq t\bigg{)}\leq 2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right).$ (7.23)
Let $\varrho=2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right)$, we obtain
$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}\leq\varrho.$ (7.24)
It follows from (7.22) that
$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t^{\prime}})]\right|>(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right)$
$\quad\leq\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}\right)$
$\quad\leq\varrho+0=\varrho.$
The desired result follows from $\|v\|_{t_{j}}^{(l+2)}\leq C\|f\|^{(l+2)}$. ∎
7.1.4. Proof of Theorem 4.4
We introduce the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$. Let $\psi$ be a convex, nondecreasing function with $\psi(0)=0$. For a random variable $X(\omega):\Omega^{\xi}\rightarrow\mathbb{R}^{m},m\in\mathbb{N}_{+}$, the Orlicz norm is defined as
$\|X\|_{\psi}=\inf\left\{C>0:\mathbb{E}\left[\psi\left(\frac{|X|}{C}\right)\right]\leq 1\right\},$ (7.25)
where ||$\cdot$|| ⋅ | is the Euclid norm on $\mathbb{R}^{m}$, and the corresponding Orlicz space is $L_{\psi}$. Note that the $L_{p}$-norm $\|\cdot\|_{p}$ is the corresponding Orlicz norm with the function $\psi(x)=x^{p}$. Denote $\psi_{\alpha}=e^{x^{\alpha}}-1,\alpha>0$ (linear near the origin when $10<\alpha<$ 1 in order for $\psi_{\alpha}$ to be convex). We say $X$ is sub-exponential if and only if $X\in L_{\psi_{1}}$. It follows from $|x|^{p}/p!\leq e^{|x|}-1$ that
$\|X\|_{p}\leq\sqrt[p]{p!}\|X\|_{\psi_{1}}.$ (7.26)
Combining the above inequality and [34, Lemma 2.2.2], we derive the following lemma.
Lemma 7.3.
For $p\geq 1$, if $\{X_{i}\}_{i=1}^{n}$ are sub-exponential random variables, then there is a constant $C_{p}$, such that
$\mathbb{E}\left[\max_{1\leq i\leq n}|X_{i}|^{p}\right]\leq C_{p}\log^{p}(1+n)\max_{i}\|X_{i}\|_{\psi_{1}}^{p}.$ (7.27)
Note that under condition (H.6), $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ is globally Lipschitz with respect to the state $x$. By classical SDE results, e.g. [13, Theorem 7.9], we have the following lemma.
Lemma 7.4.
Assume (H.6) holds. For every $p\geq 2$, there is a constant $C_{p}$, such that
$\mathbb{E}^{W}\left[\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p}\right]\leq C_{p},\quad\mathbb{E}\left[\sup_{0\leq t\leq T}|X^{\mathscr{G}}_{t}|^{p}\right]\leq C_{p}.$ (7.28)
Moreover,
$\mathbb{E}^{W}\left[\sup_{0\leq s\leq T}|\widetilde{X}_{s}-\widetilde{X}_{\delta(s)}|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}},\quad\mathbb{E}\left[\sup_{0\leq s\leq T}|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}.$ (7.29)
Proof of Theorem 4.4.
By the Lipschitz condition of $\phi$, $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential if and only if $d_{A}(\phi(0,0,\xi),a_{0})$ is sub-exponential by the triangular inequality and the condition (1) of (H.6). we take $(t_{0},x_{0})=(0,0)$ for simplicity. Recall the notation $i(t):=\max\{i:t_{i}\leq t\}$. Note that
$X^{\mathscr{G}}_{t}$ $=x_{0}+\int_{0}^{t}b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))\mathrm{d}s+\int_{0}^{t}\sigma(s,X^{\mathscr{G}}_{s})\mathrm{d}W_{s},$ (7.30)
$\widetilde{X}_{t}$ $=x_{0}+\int_{0}^{t}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,\widetilde{X}_{s})\mathrm{d}W_{s}.$
Denote $\Delta X_{t}:=X^{\mathscr{G}}_{t}-\widetilde{X}_{t}$. We have
$\sup_{0\leq u\leq t}|\Delta X_{u}|$ $=\sup_{0\leq u\leq t}\bigg{|}\int_{0}^{u}\left(b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))-b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\right)\mathrm{d}s$ (7.31)
$\quad+\int_{0}^{u}\left(b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\right)\mathrm{d}s+\int_{0}^{u}\left(\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right)\mathrm{d}W_{s}\bigg{|}.$
It follows from the Lipschitz conditions and the Burkholder-Davis-Gundy inequality that for $p\geq 2$,
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}\left(\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right)\mathrm{d}W_{s}\right|^{p}\right]$ $\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}\left|\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right|^{2}\mathrm{d}s\right)^{\frac{p}{2}}\right]$
$\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}|\Delta X_{u}|^{2}\mathrm{d}u\right)^{\frac{p}{2}}\right]\leq C_{p}\mathbb{E}\left[\int_{0}^{t}|\Delta X_{u}|^{p}\mathrm{d}u\right].$
Thus, we have
$\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]\leq$ $C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\left|\Delta X_{s}\right|^{p}+\left|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}\right|^{p}\right]\mathrm{d}s$ (7.32)
$+C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\mathrm{d}s-\int_{0}^{u}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s\right|^{p}\right].$
We focus on the third term of the above equation. Denote $U_{s}:=b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})$. Under (H.6), we have
$|U_{s}|=\left|b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\right|\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|+d_{A}(\phi(0,0,\xi_{i(s)}),a_{0})\right).$ (7.33)
Denote $\mathbb{E}_{\xi}\coloneqq\mathbb{E}\left[\cdot|\mathcal{F}^{W}\right]$ and note that $\xi$ is independent of $\mathcal{F}^{W}$.
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\leq C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]+C_{p}\mathbb{E}\left[\mathbb{E}_{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right].$ (7.34)
It follows from (7.33) and Lemma 7.3 that
$\mathbb{E}\left[\mathbb{E}_{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right]$ $\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\mathbb{E}\left[\max_{1\leq j\leq n}\left(d_{A}(\phi(0,0,\xi_{j}),a_{0})\right)^{p}\right]\right)|\mathscr{G}|^{p}\right]$ (7.35)
$\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\log^{p}(1+n)\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}^{p}\right)|\mathscr{G}|^{p}\right]$
$\leq C_{p}|\mathscr{G}|^{\frac{p}{2}},$
where in the last inequality, we used the fact that $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}$.
As for the first term on the right hand side of (7.34), denote $M_{j}:=\int_{t_{j}}^{t_{j+1}}U_{s}\mathrm{d}s$ and $\mathcal{F}_{i}:=\sigma(\xi_{j}:0\leq j\leq i-1)\lor\mathcal{F}^{W}$. Note that $\left\{(\sum_{j=0}^{i-1}M_{j},\mathcal{F}_{i})\right\}_{i=1}^{n}$ is a martingale. It follows from Doob’s inequality that
$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]=\mathbb{E}\left[\max_{1\leq i\leq n}\Big{|}\sum_{j=0}^{i-1}M_{j}\Big{|}^{p}\right]\leq C_{p}\mathbb{E}\left[\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}^{p}\right],$ (7.36)
where $C_{p}$ is a constant depend on $p$.
By (H.6) and relation (7.33), $M_{j}$ is sub-exponential conditional on $\mathcal{F}^{W}$ and with a little abuse of notations,
$\|M_{j}\|_{\psi_{1}}\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\|1\|_{\psi_{1}}+\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}\right)(t_{j+1}-t_{j}),$ (7.37)
where $\|\cdot\|_{\psi_{1}}$ is the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$. It follows from [25, Theorem 6.21] that
$\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\leq C\left(\mathbb{E}_{\xi}\left[\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}\right]+\|\max_{j}|M_{j}|\|_{\psi_{1}}\right).$ (7.38)
For the first term on the right hand side of the above equation, using the fact that $\mathbb{E}_{\xi}\left[M_{i}^{\top}M_{j}\right]=0,i\neq j$, it follows from relation (7.33) and $\psi(0,0,\xi)$ is square integrable that
$\mathbb{E}_{\xi}\bigg{[}\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}\bigg{]}\leq\sqrt{\mathbb{E}_{\xi}\bigg{[}\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}^{2}\bigg{]}}=\sqrt{\sum_{j=0}^{n-1}\mathbb{E}_{\xi}\left[|M_{j}|^{2}\right]}\leq C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}.$ (7.39)
As for the second term, it follows from [34, Lemma 2.2.2] and (7.37) that
$\|\max_{j}|M_{j}|\|_{\psi_{1}}\leq C\log(1+n)\max_{j}\|M_{j}\|_{\psi_{1}}\leq C\log(1+n)\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|.$ (7.40)
In conclusion, we have
$\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\leq C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}.$ (7.41)
Conditional on $\mathcal{F}^{W}$, we have
$\mathbb{P}_{\xi}\left(\bigg{|}\sum_{j=0}^{n-1}M_{j}\bigg{|}^{p}\geq t\right)$ $=\mathbb{P}_{\xi}\left(\psi_{1}\bigg{(}\bigg{|}\sum_{j=0}^{n-1}M_{j}\bigg{|}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\leq\psi_{1}\bigg{(}t^{\frac{1}{p}}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\right)$ (7.42)
$\leq 1/\psi_{1}\bigg{(}t^{\frac{1}{p}}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\leq 2\exp\left\{-\frac{t^{\frac{1}{p}}}{C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}}\right\}.$
Integrating the above equation with respect to $t$ from 00 to +$\infty$+ ∞, we have
$\mathbb{E}\left[\left|\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{j}))\mathrm{d}s-\int_{t_{i}}^{t_{i+1}}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s\right|^{p}\bigg{|}\mathcal{F}^{W}\right]\leq C_{p}(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p})|\mathscr{G}|^{\frac{p}{2}}.$ (7.43)
Combining (7.32), (7.34), (7.35) and (7.43), using Lemma 7.4 to bound the time difference of $X^{\mathscr{G}}$ and $\widetilde{X}$ and the moments of $\sup_{t}|\widetilde{X}_{t}|$, we have
$\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]$ $\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}\left|\Delta X_{u}\right|^{p}\right]+\mathbb{E}\left[|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}|^{p}\right]\mathrm{d}s$ (7.44)
$\quad\quad+C_{p}|\mathscr{G}|^{\frac{p}{2}}\int_{0}^{t}\mathbb{E}[(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p})]\mathrm{d}s$
$\leq C_{p}\left|\mathscr{G}\right|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}|\Delta X_{u}|\right]\mathrm{d}s.$
It follows from the Grönwall’s inequality that
$\mathbb{E}\left[\sup_{0\leq s\leq T}\left|\Delta X_{s}\right|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}.$ (7.45)
Now we prove the second part of this theorem. If $|\mathscr{G}|=\frac{T}{n}$, then for $p\geq 2$, $t\geq 0$,
$\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>t\right)=\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|^{p}>t^{p}\right)\leq t^{-p}\mathbb{E}\left[\sup_{0\leq t\leq T}|\Delta X_{t}|^{p}\right]\leq C_{p}n^{-\frac{p}{2}}t^{-p}.$ (7.46)
Let $t=n^{-\frac{1}{2}+\varepsilon}$,
$\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}n^{-\varepsilon p}.$ (7.47)
By choosing $p=\frac{1}{\varepsilon}+2$, we have
$\sum_{n=1}^{\infty}\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}\sum_{n=1}^{\infty}n^{-1-2\varepsilon}<+\infty.$ (7.48)
The desired result follows from the Borel-Cantelli lemma. ∎
7.2. Proofs of Propositions 5.1, 5.3 and 5.5
Proof of Proposition 5.1.
For the first part (1), note that
$J^{\mathscr{G}}-\tilde{J}=\left(\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right)+\sum_{i=0}^{n-1}\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}(r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s}))\mathrm{d}s\right]$ (7.49)
By Theorem 4.1, there is a constant $C_{p}$, such that
$\left|\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right|\leq C_{p}\|h\|_{C_{p}^{4}}|\mathscr{G}|.$ (7.50)
As for the second term on the right hand side of (7.49), note that if $t_{i}\leq s<t_{i+1}$, then
$\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]$ $=\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]+\mathbb{E}\left[r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]$ (7.51)
$=\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]+\mathbb{E}\left[\tilde{r}(t_{i},X^{\mathscr{G}}_{t_{i}})-\tilde{r}(t_{i},\widetilde{X}_{t_{i}})\right]$
$\quad\quad+\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right].$
Since $r\in C^{2,0}_{p}$, it follows from Itô’s lemma that
$|\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]|\leq C_{p}\|r\|_{C_{p}^{2,0}}|\mathscr{G}|.$ (7.52)
Set $f=\tilde{r}(t_{i},\cdot)$ in Theorem 4.1, we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},X^{\mathscr{G}}_{t_{i}})-\tilde{r}(t_{i},\widetilde{X}_{t_{i}})\right]\right|\leq C_{p}\|\tilde{r}(t_{i},\cdot)\|_{C_{p}^{4}}|\mathscr{G}|.$ (7.53)
Using Itô’s formula, we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]\right|\leq C_{p}\|\tilde{r}\|_{C_{p}^{2}}|\mathscr{G}|.$ (7.54)
Combining these inequalities, we prove the desired result.
For the second part (2), we start at (7.49). By Theorem 4.2, there is a constant $C$, such that
$\left|\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right|\leq C\|h\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.55)
As for the second term on the right hand side of (7.49), we still have (7.51). It follows from Lemma 7.1 that
$|\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]|\leq C\|r\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.56)
Set $f=\tilde{r}(t_{i},\cdot)$ in Theorem 4.2, we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},X^{\mathscr{G}}_{t_{i}})-\tilde{r}(t_{i},\widetilde{X}_{t_{i}})\right]\right|\leq C\|\tilde{r}(t_{i},\cdot)\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}\leq C\sup_{0\leq t\leq T}\|\tilde{r}(t,\cdot)\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.57)
Note that $\tilde{r}\in\mathcal{H}^{(l)}_{T}$, by the results of [23, Lemma 14.1.6], we have
$\left|\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]\right|\leq C\|\tilde{r}\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.58)
Combining these inequalities, we prove the desired result. ∎
Proof of Proposition 5.3.
For the variance part, note that
$\mathbb{E}\left[(S\circ V)^{2}\right]$ (7.59)
$=$ $\mathbb{E}\bigg{[}\bigg{(}\int_{0}^{T}S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t+\int_{0}^{T}S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})(\partial_{x}V(t,X^{\mathscr{G}}_{t}))^{\top}\sigma(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}W_{t}\bigg{)}^{2}\bigg{]}$
$\leq$ $2\int_{[0,T]^{2}}\mathbb{E}\left[S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})S(\delta(s),X^{\mathscr{G}}_{\delta(s)},a_{\delta(s)})g(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\right]\mathrm{d}s\mathrm{d}t$
$+2\int_{0}^{T}\mathbb{E}\left[\left|S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})(\partial_{x}V(t,X^{\mathscr{G}}_{t}))^{\top}\sigma(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\right|^{2}\right]\mathrm{d}t.$
The desired result follows from Lemma 3.1.
For the bias part, it follows from Itô’s formula that
$\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left(\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right)\right]$ (7.60)
$=\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left(\partial_{t}V+(\tilde{b}^{\bm{\pi}})^{\top}\partial_{x}V+\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}})^{2}\partial_{x}^{2}V\right)+\tilde{r}\right)(t,\widetilde{X}_{t})\mathrm{d}t\right]$
and
$\mathbb{E}[(S,V)]$ $=\sum_{i=0}^{n-1}\mathbb{E}\left[S(t_{i},X^{\mathscr{G}}_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R^{\mathscr{G}}_{t_{i}}\right)\right]$ (7.61)
$=\mathbb{E}\bigg{[}\int_{0}^{T}S(\delta(t),X^{\mathscr{G}}_{\delta(t)})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\bigg{]}$
$=\mathbb{E}\left[\int_{0}^{T}S(t,X^{\mathscr{G}}_{t})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right]-\mathbb{E}\left[\int_{0}^{T}\Delta S_{t}g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right],$
where $\Delta S_{t}:=S(t,X^{\mathscr{G}}_{t})-S(\delta(t),X^{\mathscr{G}}_{\delta(t)})$. We have
$\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left(\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right)\right]\right|$ (7.62)
$\leq\left|\mathbb{E}\left[\int_{0}^{T}Sg(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right]-\mathbb{E}\left[\int_{0}^{T}\int_{A}Sg(t,\widetilde{X}_{t},a)\bm{\pi}(\mathrm{d}a|t,\widetilde{X}_{t})\mathrm{d}t\right]\right|+\left|\mathbb{E}\left[\int_{0}^{T}\Delta S_{t}g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right]\right|.$
The second term on the right hand side of above equation is bounded by $C_{p}\|S\|_{C_{p}^{2,0}}\|g\|_{C_{p}^{0,0}}|\mathscr{G}|$ and $C\|S\|_{T}^{(l)}\|g\|_{T}^{(0)}|\mathscr{G}|^{\mathcal{X}(l)}$ using the methods of Theorem 4.1 and Theorem 4.2, respectively. The desired results follow from the application of Proposition 5.1. ∎
Proof of Proposition 5.5.
For the variance part, note that
$\mathrm{Var}\left(S\Box V\right)$
$\leq\mathbb{E}\left[(S\Box V)^{2}\right]$
$=\mathbb{E}\left[\left(\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\left(\int_{t_{i-1}}^{t_{i}}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}s+\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\right)^{2}\right)^{2}\right]$
$\leq C_{p}\|S\|_{C_{p}^{0,0}}^{2}\left(\|g_{1}\|_{C_{p}^{0,0}}^{2}+\|V\|_{C_{p}^{2}}^{2}\right)^{2}.$
For the bias part, it follows from Itô’s formula that
$S\Box V$ $=\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\left(V(t_{i},X^{\mathscr{G}}_{t_{i}})-V(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+R^{\mathscr{G}}_{t_{i-1}}\right)^{2}$ (7.63)
$=\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\left(\int_{t_{i-1}}^{t_{i}}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}s+\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\right)^{2}$
$=\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\bigg{(}\bigg{(}\int_{t_{i-1}}^{t_{i}}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}s\bigg{)}^{2}$
$\quad+2\int_{t_{i-1}}^{t_{i}}\big{(}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-g_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\big{)}\mathrm{d}s\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}$
$\quad+2\Delta t_{i}g(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}$
$\quad+\bigg{(}\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\bigg{)}^{2}\bigg{)}.$
Note that the first and second terms are bounded by $C_{p}(\|S\|_{C_{p}^{2,0}}\|g_{1}\|_{C_{p}^{2,0}}^{2}+\|g_{1}\|_{C_{p}^{2,0}}\|V\|_{C_{p}^{2}})|\mathscr{G}|$ and $C(\|S\|_{T}^{(l)}(\|g_{1}\|_{T}^{(l)})^{2}+\|g_{1}\|_{T}^{(l)}\|V\|_{T}^{(l)})|\mathscr{G}|^{\mathcal{X}(l)}$ under different types of conditions, and third term is mean zero. For the last term, we have
$\mathbb{E}\left[\sum_{i=0}^{n-1}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\left(\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\right)^{2}\right]$ (7.64)
$=\mathbb{E}\left[\int_{0}^{T}S(\delta(t),X^{\mathscr{G}}_{\delta(t)},a_{\delta(t)})g_{2}(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right].$
It converges to $\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg}_{2}(t,\widetilde{X}_{t})\mathrm{d}t\right]$ with error bounds $C_{p}(\|S\|_{C_{p}^{2,0}}\|g_{2}\|_{C_{p}^{0,0}}+\|Sg_{2}\|_{C_{p}^{2,0}}+\|\widetilde{Sg_{2}}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|_{C_{p}^{4}})|\mathscr{G}|$ and $C(\|S\|_{T}^{(l)}\|g_{2}\|_{T}^{(0)}+\|Sg_{2}\|_{T}^{(l)}+\|\widetilde{Sg_{2}}\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|^{(l)})|\mathscr{G}|^{\mathcal{X}(l)}$ under different types of conditions by the results of Proposition 5.1. In particular, if $S$ do not depend on $a$, then
$\mathbb{E}\left[\int_{0}^{T}\int_{A}S\partial_{x}V^{\top}\sigma\sigma^{\top}\partial_{x}V(t,\widetilde{X}_{t},a)\bm{\pi}(\mathrm{d}a|t,\widetilde{X}_{t})\mathrm{d}t\right]$ $=\mathbb{E}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\partial_{x}V(t,\widetilde{X}_{t})^{\top}(\tilde{\sigma}^{\bm{\pi}})^{2}(t,\widetilde{X}_{t})\partial_{x}V(t,\widetilde{X}_{t})\mathrm{d}t\right]$
$=\mathbb{E}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\mathrm{d}\left<V(\cdot,\widetilde{X}_{\cdot})\right>_{t}\right].$
This proves the desired results. ∎
7.3. Proofs of Examples 4.1 and 4.2
Proof of Example 4.1.
It suffices to work on $\tilde{\sigma}^{\bm{\pi}}$. Note that
$\partial_{x}^{s}(\tilde{\sigma}^{\bm{\pi}})^{2}(t,x)=\sum_{s_{1}+s_{2}=s}\int_{A}\partial_{x}^{s_{1}}\sigma\sigma^{\top}(t,x,a)\partial_{x}^{s_{2}}H(t,x,a)\mu(\mathrm{d}a)$ (7.65)
is bounded. The desired result follows from (7.70) and $\sigma\sigma^{\top}$ is uniformly elliptic. ∎
Proof of Example 4.2.
For simplicity, we only prove that $\tilde{\sigma}^{\bm{\pi}}$ has bounded derivatives with respect to $x$ for the case $d=m=1$. Note that
$\int_{\mathbb{R}}\sigma^{2}(t,x,a)\varphi(a|\mu(t,x),\Sigma(t,x))\mathrm{d}a=\int_{\mathbb{R}}\sigma^{2}(t,x,\mu(t,x)+\Sigma(t,x)u)\varphi(u)\mathrm{d}u:=\lambda(t,x),$ (7.66)
where $\varphi$ is the standard Gaussian density. Note that
$\partial_{x}\lambda^{\frac{1}{2}}$ $=\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}\lambda$ (7.67)
$\partial_{x}^{2}\lambda^{\frac{1}{2}}$ $=-\frac{1}{4}\lambda^{-\frac{3}{2}}(\partial_{x}\lambda)^{2}+\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}^{2}\lambda$
$\partial_{x}^{3}\lambda^{\frac{1}{2}}$ $=\frac{3}{8}\lambda^{-\frac{5}{2}}(\partial_{x}\lambda)^{3}-\frac{3}{4}\lambda^{-\frac{3}{2}}\partial_{x}\lambda\partial_{x}^{2}\lambda+\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}^{3}\lambda$
$\partial_{x}^{4}\lambda^{\frac{1}{2}}$ $=-\frac{15}{16}\lambda^{-\frac{7}{2}}(\partial_{x}\lambda)^{4}+\frac{9}{4}\lambda^{-\frac{5}{2}}(\partial_{x}\lambda)^{2}\partial_{x}^{2}\lambda-\frac{3}{4}\lambda^{-\frac{3}{2}}(\partial_{x}^{2}\lambda)^{2}-\lambda^{-\frac{3}{2}}\partial_{x}\lambda\partial_{x}^{3}\lambda+\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}^{4}\lambda$
Under the conditions of this proposition, we can exchange the order of integration and differentiation, therefore,
$\partial_{x}\lambda=\partial_{x}\int_{\mathbb{R}}\sigma^{2}\varphi(u)\mathrm{d}u=\int_{\mathbb{R}}2\sigma(\partial_{x}\sigma+\partial_{a}\sigma(\partial_{x}\mu+\partial_{x}\Sigma u))\varphi(u)\mathrm{d}u.$ (7.68)
It follows from Hölder’s inequality that
$|\partial_{x}\lambda|\leq 2\sqrt{\int_{\mathbb{R}}\sigma^{2}\varphi(u)\mathrm{d}u}\sqrt{\int_{\mathbb{R}}(\partial_{x}\sigma+\partial_{a}\sigma(\partial_{x}\mu+\partial_{x}\Sigma u))^{2}\varphi(u)\mathrm{d}u}\leq C\lambda^{\frac{1}{2}}.$ (7.69)
One can use the same method to obtain that for $s\leq 4$, there is a constant $C$ such that
$|\partial_{x}^{s}\lambda|\leq C(1+\lambda^{\frac{1}{2}}).$ (7.70)
Note that $\sigma\sigma^{\top}$ is uniformly elliptic. There is a $\varepsilon>0$ such that $\lambda^{-1}\leq\varepsilon^{-1}$. The desired result follows from (7.70) and (7.67). ∎
Appendix A Proofs of Lemmas 3.1 and 7.2
Proof of Lemma 3.1.
Throughout this proof, we denote by $C\geq 0$ a generic constant independent of $t_{i}$, which may take a different value at each occurrence. Note that $X^{\mathscr{G}}_{0}$ is deterministic and hence is $p$-th power integrable for all $p\geq 2$. Suppose that for some $i=0,\ldots,n-2$, (3.3) admits a unique strong solution $X^{\mathscr{G}}$ on $[0,t_{i}]$ satisfying ${\mathbb{E}}[\sup_{t\in[0,t_{i}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. Observe that by (H.1) and the Cauchy-Schwarz inequality, for all $p\geq 2$,
${\mathbb{E}}\left[\left(\int_{t_{i}}^{t_{i+1}}|b(t,0,a_{t_{i}})|{\mathrm{d}}t\right)^{p}\right]$ $\leq C(1+{\mathbb{E}}\left[d_{A}(a_{t_{i}},a_{0})^{p}\right])=C(1+{\mathbb{E}}\left[{\mathbb{E}}\left[d_{A}(a_{t_{i}},a_{0})^{p}\mid X^{\mathscr{G}}_{t_{i}}\right]\right])$
$=C\left(1+{\mathbb{E}}\left[\int_{A}d_{A}(a,a_{0})^{p}\bm{\pi}(da|t_{i},X^{\mathscr{G}}_{t_{i}})\right]\right)\leq C\left(1+{\mathbb{E}}\left[|X^{\mathscr{G}}_{t_{i}}|^{p}\right]\right)<\infty.$
Similar arguments show that
${\mathbb{E}}\left[\left(\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{2}{\mathrm{d}}t\right)^{p/2}\right]\leq{\mathbb{E}}\left[\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{p}{\mathrm{d}}t\right]<\infty.$
Since $x\mapsto b(t,x,a)$ and $x\mapsto\sigma(t,x,a)$ are Lipschitz continuous uniformly in $(t,a)$, by [40, Theorems 3.3.1 and 3.4.3], (3.3) with the initial state $X^{\mathscr{G}}_{t_{i}}$ at $t_{i}$ admits a unique strong solution $X^{\mathscr{G}}$ on $[t_{i},t_{i+1}]$ satisfying ${\mathbb{E}}[\sup_{t\in[t_{i},t_{i+1}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. This along with an inductive argument shows the desired well-posedness result.
Apply the notation $\delta(u):=t_{i}$ for $u\in[t_{i},t_{i+1})$, then it satisfies that
$X^{\mathscr{G}}_{s}=x_{0}+\int_{t}^{s}b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}u+\int_{t}^{s}\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}W_{u}.$ (A.1)
Given that $b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})$ and $\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})$ have finite moments, there is a constant $C_{p}>0$, depending only on $p$, such that
$\mathbb{E}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]$ $\leq C_{p}\left|x_{0}\right|^{p}+C_{p}\int_{t}^{s}\mathbb{E}\left[\left|b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]\mathrm{d}u$ (A.2)
$+C_{p}\int_{t}^{s}\mathbb{E}\left[\left|\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]\mathrm{d}u.$
Noting that by (H.1), for $h=b$ or $\sigma$, it satisfies that
$\mathbb{E}\left[\left|h(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]$ $\leq\mathbb{E}\left[C_{p}\left(1+\left|X^{\mathscr{G}}_{u}\right|^{p}+d_{A}(a_{0},a_{\delta(u)})^{p}\right)\right]$ (A.3)
$\leq C_{p}\left(1+\mathbb{E}\left[\left|X^{\mathscr{G}}_{u}\right|^{p}\right]+\mathbb{E}\left[\left|X^{\mathscr{G}}_{\delta(u)}\right|^{p}\right]\right),$
where the second inequality follows from taking conditional expectation $\mathbb{E}[\cdot|X^{\mathscr{G}}_{\delta(u)}]$ and using Condition (H.13). Substituting (A.3) into (A.2), we obtain
$\mathbb{E}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]\leq C_{p}\left|x_{0}\right|^{p}+C_{p}\int_{t}^{s}\left(1+\mathbb{E}\left[\left|X^{\mathscr{G}}_{u}\right|^{p}\right]+\mathbb{E}\left[\left|X^{\mathscr{G}}_{\delta(u)}\right|^{p}\right]\right)\mathrm{d}u.$ (A.4)
Choosing $s=\delta(s)$ in (A.4), and adding it with (A.4), we obtain
$\mathbb{E}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]$ $+\mathbb{E}\left[\left|X^{\mathscr{G}}_{\delta(s)}\right|^{p}\right]$
$\leq$ $2C_{p}\left|x_{0}\right|^{p}+2C_{p}\int_{t}^{s}\left(1+\mathbb{E}\left[\left|X^{\mathscr{G}}_{u}\right|^{p}\right]+\mathbb{E}\left[\left|X^{\mathscr{G}}_{\delta(u)}\right|^{p}\right]\right)\mathrm{d}u.$
The desired result follows from the Grönwall’s inequality. ∎
Proof of Lemma 7.2.
We first prove $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. Observe that by (H.5), for all $p\geq 2$,
$\mathbb{E}^{W}\left[\left(\int_{t_{i}}^{t_{i+1}}|b(t,0,a_{t_{i}})|\mathrm{d}t\right)^{p}\right]\leq C<\infty.$ (A.5)
Similarly, we have
$\mathbb{E}^{W}\left[\left(\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{2}\mathrm{d}t\right)^{p/2}\right]\leq\mathbb{E}^{W}\left[\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{p}\mathrm{d}t\right]\leq C<\infty.$ (A.6)
Note that $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$, and all Itô integrals appear here equal to 0 when taking the conditional expectation $\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$. Analogous to the proofs of Lemma 3.1 and [40, Theorem 3.4.3], we obtain that $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
Next, we aim to prove that the moments of $X^{\mathscr{G}}_{t}$ is independent of $|\mathscr{G}|$ and obtain (7.10). Use the notation $\delta(u):=t_{i}$ for $u\in[t_{i},t_{i+1})$, then it satisfies that
$X^{\mathscr{G}}_{s}=x_{0}+\int_{0}^{s}b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}u+\int_{0}^{s}\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}W_{u}.$ (A.7)
Note that $b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})$ and $\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})$ have finite moments by (H.5). There is a constant $C_{p}>0$, depending only on $p$, such that
$\mathbb{E}^{W}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]$ $\leq C_{p}\left|x_{0}\right|^{p}+C_{p}\int_{0}^{s}\mathbb{E}^{W}\left[\left|b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]\mathrm{d}u$ (A.8)
$+C_{p}\int_{0}^{s}\mathbb{E}^{W}\left[\left|\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]\mathrm{d}u.$
Noting that by condition (H.5), for $h=b$ or $\sigma$, it satisfies that
$\mathbb{E}^{W}\left[\left|h(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\right|^{p}\right]$ $\leq\mathbb{E}^{W}\left[C_{p}\left(1+\left|X^{\mathscr{G}}_{u}\right|^{p}\right)\right].$ (A.9)
Substituting (A.9) into (A.8), we obtain
$\mathbb{E}^{W}\left[\left|X^{\mathscr{G}}_{s}\right|^{p}\right]\leq C_{p}\left|x_{0}\right|^{p}+C_{p}\int_{t}^{s}\left(1+\mathbb{E}^{W}\left[\left|X^{\mathscr{G}}_{u}\right|^{p}\right]\right)\mathrm{d}u.$ (A.10)
The desired result follows from the Grönwall’s inequality. ∎

### v2

@@
 Abstract.
-Stochastic policies are widely used in continuous-time reinforcement learning algorithms.
+Stochastic policies (also known as relaxed controls) are widely used in continuous-time reinforcement learning algorithms.
 However, executing a stochastic policy and evaluating its performance in a continuous-time environment remain open challenges.
@@
 We explicitly quantify the convergence rate based on the regularity of the coefficients and establish an optimal first-order convergence rate for sufficiently regular coefficients.
-Additionally, we show that the same convergence rates hold with high probability concerning the sampling noise, and further establish a $1/2$-order almost sure convergence when the volatility is not controlled.
+Additionally, we prove a $1/2$-order weak convergence rate that holds uniformly over the sampling noise with high probability, and establish a $1/2$-order pathwise convergence for each realization of the system noise in the absence of volatility control.
 Building on these results, we analyze the bias and variance of various policy evaluation and policy gradient estimators based on discrete-time observations.
@@
 Reinforcement learning (RL) has catalyzed significant changes across various domains by attacking complex sequential decision-making challenges in a data-driven way.
-The theoretical framework for analyzing RL problems and designing RL algorithms has been predominantly based on the discrete-time Markov decision processes [5, 30].
+The theoretical framework for analyzing RL problems and designing RL algorithms has been predominantly based on the discrete-time Markov decision processes [35].
 However, many practical applications require real-time interaction with environments where time progresses continuously, such as autonomous driving, stock trading, and vehicle routing.
@@
 Early studies on continuous-time RL have primarily focused on deterministic systems, where state variables follow ordinary differential equations; see, e.g., [1, 10, 38, 29, 24].
-Exploratory control and stochastic policy
-Recently, [36] proposed a novel exploratory control formulation with stochastic policies for continuous-time RL problems, which has served as the foundation for designing various RL algorithms; see, e.g., [18, 19, 20, 15, 8, 11, 37, 12, 29, 32].
-This control formulation in [36] has three key features: (1) The agent incorporates stochastic policies into the control problem, which models exploration in RL.
-These policies are mappings from the state space to distributions over the action space, aligning with the notion of “relaxed control” in classical control theory.
-(2) The agent continuously samples actions from a stochastic policy, and the state process evolves according to a new diffusion process, with its coefficients integrated with respect to the sampling distribution specified by the stochastic policy.
-This auxiliary state dynamic is often referred to as the exploratory or aggregated dynamics.
-(3) The optimal policy remains stochastic due to the inclusion of an additional entropy regularizer in the cost objective, in contrast to the deterministic/strict policies typically found in standard control problems.
+Stochastic policy in continuous-time RL
+Recently, [42] introduced a novel exploratory control framework for continuous-time RL that fundamentally relies on the use of stochastic policies.
+Stochastic policies are mappings from the state space to distributions over the action space, aligning with the notion of “relaxed control” in classical control theory.
+By sampling actions from these stochastic policies, the agent assigns positive probabilities to different actions at each state, unlike the deterministic/strict policies typically employed in standard control problems.
+This approach effectively captures the concept of exploration in RL, which is essential for the agents to learn about the environment.
+[^1] Exploration in RL broadly refers to deliberately acting suboptimally to gather information about the environment.
+Stochastic policies achieve this by randomizing actions, but exploration can also be driven by methods such as adding exploration bonuses to encourage actions in less-certain outcomes; see [35, Chapter 2].
+Since then, the use of stochastic policies and the corresponding exploratory control framework have become a foundational basis for developing a wide range of RL algorithms for continuous-time problems; see, e.g., [20, 21, 22, 16, 9, 12, 43, 47, 13, 37, 34].
 Execution of stochastic policies
-However, despite its success in algorithm design, the use of stochastic policies in continuous-time RL problems introduces theoretical and practical gaps.
-Specifically, the connection between exploratory dynamics with aggregated coefficients and state dynamics controlled by frequent sampling from a stochastic policy remains unjustified.
-In fact, interacting with a diffusion process by continuously sampling from a stochastic policy requires generating uncountably many independent random variables, which raises measure-theoretical issues and renders the resulting controlled state dynamics ill-defined (see the end of Section 2 for more details).
-To address this issue, [32] proposes approximating exploratory dynamics by sampling actions from a stochastic policy only at discrete time points.
-This approach naturally aligns with the practical implementation of stochastic policies.
-However, for algorithms designed based on aggregated dynamics, executing a stochastic policy in discrete time introduces a bias that depends on the sampling frequency.
-This leads to the following natural question:
-What is the precise impact of sampling frequency on the performance of a stochastic policy and the corresponding learning algorithms?
-To the best of our knowledge, this question has only been addressed in the context of linear-quadratic (LQ) RL problems with Gaussian policies [12, 32].
+However, despite its popularity in algorithm design, the use of stochastic policies in continuous-time RL presents both theoretical and practical challenges.
+In particular, the behavior of the state dynamics under frequent sampling from a stochastic policy has not been rigorously analyzed.
+In [42], the authors assume that agents continuously sample from a stochastic policy and heuristically argue that the resulting dynamics follow a new diffusion process, with coefficients aggregated according to the sampling distribution of the policy.
+This aggregated dynamics offers a mathematically convenient framework for designing continuous-time RL algorithms.
+However, it requires generating uncountably many independent random variables, which is computationally infeasible and raises measure-theoretical concerns, rendering the controlled dynamics ill-defined.
+As a consequence, such an aggregated dynamics are not directly implementable.
+(see footnote 3 for details).
+In this paper, we consider a natural and practical approach to implement stochastic policies.
+Rather than sampling actions continuously, we sample them at discrete time points and apply them to the continuous-time state system.
+As a result, the state dynamics are governed by a piecewise constant control process, whose evolution differs from the aggregated dynamics studied in [42].
+For algorithms developed under the aggregated dynamics, such discrete-time execution of stochastic policies introduces a bias that must be carefully quantified.
+In summary, this paper aims to address the following two questions:
+Q1:
+Does the aggregated dynamics in [42] capture the state dynamics induced by frequently sampling a stochastic policy? If so, in what sense, and what is the resulting error?
+Q2:
+What is the precise impact of the sampling frequency of a stochastic policy on the performance of the corresponding learning algorithms?
+To the best of our knowledge, these questions have only been studied in the context of linear-quadratic (LQ) RL problems with Gaussian policies [13, 37].
 In this special setting, due to the LQ structure, the analysis is simplified by examining the discrepancy between the first and second moments of the aggregated dynamics and the controlled dynamics with sampled actions.
@@
 Our contributions
-This paper analyzes the accuracy of the discrete-time execution of stochastic policies for general diffusion processes and quantifies its impact on the associated learning algorithms.
+This paper analyzes the convergence of the discrete-time execution of stochastic policies for general diffusion processes and quantifies its impact on the associated learning algorithms.
   • We show that when actions are sampled from a stochastic policy at a discrete time grid $\mathscr{G}$ and executed as a piecewise constant control process, the resulting state process, referred to as the sampled dynamics, converges weakly to the aggregated dynamics as the time stepsize approaches zero.
-In particular, under sufficient regular coefficients, we show that the convergence rate is of first order $O(|\mathscr{G}|)$, with $|\mathscr{G}|$ being the time stepsize (Theorem 4.1).
-This matches the optimal convergence rate for standard piecewise constant control approximations [16].
+In particular, under sufficiently regular coefficients, we show that the convergence rate is of first order $O(|\mathscr{G}|)$, with $|\mathscr{G}|$ being the time stepsize (Theorem 4.1).
+This matches the optimal convergence rate for standard piecewise-constant control approximations with smooth coefficients [18].
 We further quantify the precise weak convergence rate for cases with less regular coefficients (Theorem 4.2).
@@
   • We prove that the sampled dynamics, in fact, converges almost surely with respect to the randomness in policy execution.
-Specifically, we show that the weak convergence order to the aggregated dynamics holds with high probability with respect to the randomness used to sample actions (Theorem 4.3).
-We further show that when the volatility is uncontrolled, the sampled dynamics converges almost surely to the aggregated dynamics with an order of $\mathcal{O}(|\mathscr{G}|^{1/2})$, for each fixed realization of the Brownian motion (Theorem 4.4).
-The analysis leverages concentration inequalities for bounded and sub-exponential martingale difference sequences to control the concentration behavior of the sampling errors in terms of the sampling frequency (see the discussions below Theorems 4.3 and 4.4).
-    The almost sure convergence result indicates that the convergence of the sampled dynamics to the aggregated dynamics can be interpreted as a law of large numbers in time, i.e., the convergence can be achieved by refining the sampling frequency while fixing the realization of the system noise.
-Note that such an almost sure convergence fails when the volatility is controlled, as shown in Example 4.6.
+Specifically, we show that the sampled dynamics converge at order $\mathcal{O}(|\mathscr{G}|^{1/2})$, averaged over the system noise but uniformly across the realizations of the sampling noise with high probability (Theorem 4.3).
+We further show that when the volatility is uncontrolled, the sampled dynamics converge at order $\mathcal{O}(|\mathscr{G}|^{1/2})$ for almost every realization of the system noise (Theorem 4.4).
+We show in Example 4.7 that the condition of uncontrolled volatility is necessary for the convergence of the sampled dynamics for each fixed system noise.
+    These almost sure convergence results allow for interpreting the convergence of the sampled dynamics to the aggregated dynamics as a law of large numbers in time, i.e., convergence is achieved by increasing the sampling frequency without averaging over multiple realizations of the sampling noises at given time points.
+See Corollary 4.1 and Remark 4.3 for its implications on the improved sample complexity of associated Monte Carlo estimators.
   • We apply the weak convergence rate to quantify the error incurred when implementing continuous-time RL algorithms based on observations from sampled dynamics.
@@
 The q-learning with quadratic variation penalty for the risk-sensitive RL proposed in [19] is also shown to have a similar order of discretization error.
-Our approaches and most related works
-This paper quantifies the accuracy of the discrete-time execution of a stochastic policy and its impact on continuous-time RL algorithms.
-The accuracy of policy execution has been studied in [14, 32] for LQ problems with Gaussian policies.
-However, as mentioned earlier, the techniques developed in [14, 32] are not applicable to RL problems with general state dynamics and stochastic policies considered in this paper.
-For general state dynamics, the weak convergence of sampled dynamics has recently been studied by [4, 3].
-They adopt a different definition of sampled dynamics, treating it as being driven by a random measure, and establish the weak convergence of the state process without quantifying the convergence order.
-In contrast to the time-varying control considered in [4, 3], we analyze sampled dynamics controlled by a piecewise constant control process.
-Moreover, we take a completely different approach and establish the weak convergence rate under various regularity conditions by leveraging the solution regularity of the associated partial differential equations (PDEs).
-Beyond weak convergence, we also establish convergence in the almost sure sense with respect to the policy execution noise.
-These strong convergence results require a careful analysis of the concentration properties of the state dynamics concerning the policy execution noise.
-In addition to discrete sampling error, we also analyze the error incurred when approximating continuous-time RL algorithms using discrete-time observations.
-[18] analyzes a similar error but assumes that observations are generated from the aggregated dynamics.
-[2] analyzes both errors for LQ problems with uncontrolled volatility.
-Our analysis of the policy execution error is fundamentally different from classical numerical analysis of time discretizations for stochastic differential equations (SDEs).
-In particular, numerical analysis of SDEs focuses on the error caused by discretizing the system noise, such as Brownian motion, while both the SDE and its time-discretized version have the same coefficients.
-In our setting, we treat the controlled SDE as a black-box environment evolving continuously and focus on the error introduced by the piecewise constant execution of a stochastic policy.
-As a result, the environment noise is not discretized, but the limiting SDE (the aggregated dynamics) and the sampled dynamics have different coefficients.
-The former involves aggregated coefficients, while the latter has random coefficients determined by the specific procedure used to sample the policy.
-The rest of the paper is organized as follows.
+Our approaches and key technical challenges
+Our convergence rate analysis for the discrete-time execution of stochastic policies requires addressing several technical issues beyond those encountered in existing literature.
+These challenges arise because the sampled dynamics contain two distinct sources of randomness: one from control randomization and the other from the underlying system noise.
+Specifically, the sampled dynamics have random coefficients due to control randomization, and they converge to a limiting diffusion process with a different set of deterministic coefficients, resulting from aggregation over the stochastic policy.
+Quantifying this convergence rate has not been addressed in the literature, and it is fundamentally different from the classical analysis of time discretizations for stochastic differential equations (SDEs) (see, e.g., [25]).
+Specifically, standard numerical analysis of SDEs focuses on the error caused by discretizing the system noise, while both the SDE and its time-discretized version share the same coefficients.
+In our setting, the environment noise is not discretized; instead, the error arises from the control randomization.
+Moreover, we quantify the convergence rate of the sampled dynamics under different criteria, including the weak error with respect to both system and sampling noises, the weak error conditioned on the sampling noise, and almost sure convergence with respect to both noises.
+Identifying precise conditions to ensure convergence under these criteria requires carefully distinguishing the roles of system and sampling noises and employing different analytical techniques, including the regularity theory of partial differential equations (PDEs) and various martingale concentration inequalities.
+In particular, when quantifying the weak error conditioned on the sampling noise (Theorem 4.3), we leverage the averaging over the system noise and apply the Azuma–Hoeffding inequality for bounded martingales.
+For the almost sure convergence with respect to both noises (Theorem 4.4), more precise estimates involving conditional Orlicz norms of sub-exponential martingales are required to carefully characterize the concentration behavior with respect to both system and sampling noises.
+To the best of our knowledge, this is the first work that applies concentration inequalities to the convergence rate analysis of sampling schemes for SDEs.
+Most related works
+This paper quantifies the accuracy of the discrete-time execution of a stochastic policy and the performance of continuous-time RL algorithms based on discrete-time observations.
+The accuracy of policy execution has been studied in [15, 37] for LQ problems with Gaussian policies; however, their techniques do not extend to the RL problems with general state dynamics and stochastic policies as considered here.
+Regarding the performance of continuous-time RL algorithms using discrete-time observations, most existing works assume that the observations are generated directly from the desired continuous-time state dynamics, independent of time discretizations (see, e.g., [2, 20, 48, 49]).
+This assumption no longer holds in the present setting with stochastic policies, since the agent observes the sampled dynamics, which differ from their continuous-time limit.
+For general state dynamics and general stochastic policies, the weak convergence of sampled dynamics has recently been studied in [4, 3].
+In their setting, the policy is evaluated continuously at the current state, and weak convergence is established by interpreting the sampled action as a random measure, though without providing a convergence rate.
+In contrast, we consider a more practical setting where the policy is evaluated only at discrete time points, so that the sampled dynamics are governed by a piecewise constant control process.
+Moreover, we establish not only the weak convergence rate, leveraging the regularity theory of PDEs, but also the high-probability bound and almost sure convergence rate with respect to the policy execution noise.
+For state dynamics with drift controls only, [7] establishes the strong convergence of time-discretized sampled dynamics under the assumption that the drift coefficient is bounded.
+This boundedness condition is restrictive, as it rules out the linear dynamics commonly encountered in many RL problems [2, 15, 37].
+Our result removes this limitation by allowing drift coefficients with linear growth.
+To address the challenges posed by unbounded coefficients, we employ advanced sub-exponential concentration inequalities with conditional Orlicz norms to quantify the convergence rate.
+Organization of the paper
 Section 2 briefly reviews the use of stochastic policies in continuous-time RL.
@@
 Section 4 presents the main convergence results of the sampled dynamics, and Section 5 applies them to analyze continuous-time RL algorithms.
-Finally, we conclude in Section 6.
-The proofs of all statements are included in Section 7.
+Section 6 presents the proofs of the main results.
+Appendix A contains the proofs of useful Lemmas and intermediary results.
 2. Preliminary
-In this section, we review the exploratory control formulation and the use of stochastic policies for continuous-time RL algorithms. The precise assumptions and setup can be found in Section 3.
+In this section, we review the use of stochastic policies in continuous-time RL problems, while avoiding unnecessary technical details. The precise assumptions and setup are given in Section 3.
 2.1. Classical stochastic control
-Recall that in a classical stochastic control problem, an agent aims to control the state process and optimize an objective functional through an adapted control process taking values in a Polish space $A$.
-Specifically, given an admissible control process $\{a_{t}\}_{t\geq 0}$, the corresponding state process $\{X_{t}\}_{t\geq 0}$ is governed by the following dynamics: [^1] To simplify the presentation, we assume that the state processes have a deterministic initial condition.
-However, our analysis and results extend naturally to cases with a random initial condition, provided it is independent of the Brownian motion $W$ and satisfies appropriate integrability conditions.
-We also assume that the Brownian motion’s dimension equals the state variable’s dimension; otherwise, we can introduce more Brownian motions or state variables to make them equal.
+Recall that in a classical stochastic control problem, an agent aims to control the state process and optimize an objective functional through a non-anticipative control process taking values in a Polish space $A$.
+Specifically, given a non-anticipative control process $\{a_{t}\}_{t\geq 0}$, the corresponding state process $\{X_{t}\}_{t\geq 0}$ is governed by the following dynamics: [^2] To simplify the presentation, we assume that the state processes have a deterministic initial condition.
+The analysis and results extend naturally to cases with a random initial condition that is independent of the Brownian motion $W$ and satisfies appropriate integrability conditions.
+We also assume without loss of generality that the Brownian motion and the state variable have the same dimensions.
 $\mathrm{d}X_{t}=b(t,X_{t},a_{t})\mathrm{d}t+\sigma(t,X_{t},a_{t})\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.1)
-where $x_{0}\in{\mathbb{R}}^{d}$ is a given initial state, $b:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d\times d}$ are given drift and volatility coefficients, respectively, and $W$ is an $d$-dimensional Brownian motion supported on a probability space $(\Omega,\mathcal{F},\mathbb{P})$.
+where $x_{0}\in{\mathbb{R}}^{d}$ is a given initial state, $b:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times A\to\mathbb{R}^{d\times d}$ are given drift and volatility coefficients, respectively, and $W$ is a $d$-dimensional Brownian motion supported on a probability space $(\Omega,\mathcal{F},\mathbb{P})$.
 The agent’s objective is to maximize the following objective
 $\mathbb{E}\left[\int_{0}^{T}r(t,X_{t},a_{t})\mathrm{d}t+h(X_{T})\right]$ (2.2)
-over all admissible controls $\{a_{t}\}_{t\geq 0}$, where $r:[0,T]\times{\mathbb{R}}^{d}\times A\to{\mathbb{R}}$ is a given instantaneous reward function, and $h:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a given terminal reward function.
-It is known that under suitable conditions (see [39]), an optimal control for (2.2) exists and is given by a feedback form, i.e., there exists a measurable function $\bm{u}:[0,T]\times\mathbb{R}^{d}\to A$ such that the optimal control is obtained by $a_{t}=\bm{u}(t,X_{t}^{\bm{u}})$ for all $t\in[0,T]$, where $\{X_{t}^{\bm{u}}\}_{t\geq 0}$ satisfies the following dynamics:
+over all non-anticipative controls, where $r:[0,T]\times{\mathbb{R}}^{d}\times A\to{\mathbb{R}}$ is a given instantaneous reward function, and $h:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a given terminal reward function.
+We refer the reader to [45, Chapter 2, Section 4] for a more detailed formulation of stochastic control problems.
+It is known that under mild regularity conditions (see e.g., [32, 26]), an optimal control of (2.2) exists and can be expressed in feedback form.
+That is, there exists a function $\bm{u}:[0,T]\times\mathbb{R}^{d}\to A$, called a policy, such that the optimal control is obtained by $a_{t}=\bm{u}(t,X_{t}^{\bm{u}})$ for all $t\in[0,T]$, where $\{X_{t}^{\bm{u}}\}_{t\geq 0}$ satisfies the following dynamics:
 $\mathrm{d}X_{t}=b(t,X_{t},\bm{u}(t,X_{t}))\mathrm{d}t+\sigma(t,X_{t},\bm{u}(t,X_{t}))\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.3)
-The function can be determined using the coefficients $b,\sigma,r,$ and $h$. Once determined, the optimal control can be implemented by evaluating ${\bm{u}}$ at the current state of the system.
-2.2. RL, stochastic policy and exploratory control
+The policy $\bm{u}$} can be determined using the coefficients $b,\sigma,r,$ and $h$. Once determined, the optimal control can be implemented by evaluating $\bm{u}$} at the current state of the system.
+2.2. Stochastic policy in RL and the associated state dynamics
 In the RL problem, the agent seeks to optimize the objective (2.2) by directly interacting with the dynamics (2.1), experimenting with different actions, and refining their strategies based on the observed responses.
 Unlike traditional control settings, the agent treats (2.1) as a black-box environment evolving in continuous time and learns an optimal policy solely from observations, without direct access to the coefficients $b,\sigma,r,$ and $h$.
-To avoid getting trapped in local optima, a key feature of RL is explicitly exploring the environment through randomized actions.
+To avoid getting trapped in local optima, a key feature of RL is explicitly exploring the environment.
 A common approach is to use stochastic policies $\bm{\pi}:[0,T]\times{\mathbb{R}}^{d}\to\mathcal{P}(A)$, which maps the current state to a probability distribution over the action space.
-Given a stochastic policy $\bm{\pi}$, the agent interacts with the environment by sampling noisy actions from $\bm{\pi}$ and observing the resulting system states.
+Given a stochastic policy $\bm{\pi}$, the agent interacts with the state by sampling noisy actions from $\bm{\pi}$ and observing the resulting state transitions.
 Over time, the agent refines the stochastic policy to optimize the objective.
-To describe the performance of a stochastic policy $\bm{\pi}$, [36] suggests that the associated state process $\tilde{X}$ evolves according to the following dynamics (called the exploratory dynamics therein):
+A crucial step in designing learning algorithms with stochastic policies is to characterize the performance of a stochastic policy $\bm{\pi}$.
+[42] suggests that given the policy $\bm{\pi}$, the associated state process $\widetilde{X}$ evolves according to the following dynamics:
 ${\mathrm{d}}X_{t}=\tilde{b}^{\bm{\pi}}(t,X_{t})\mathrm{d}t+\tilde{\sigma}^{\bm{\pi}}(t,X_{t})\mathrm{d}W_{t},\quad t\in[0,T];\quad X_{0}=x_{0},$ (2.4)
@@
 where $(\cdot)^{1/2}$ is the matrix square root of positive semidefinite matrices, and $(\cdot)^{\top}$ is the matrix transpose. Moreover, the associated objective function is given by:
-$\mathbb{E}^{W}\left[\int_{0}^{T}\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t+h(\widetilde{X}_{T})\right],\quad\textnormal{with $\tilde{r}(t,x)\coloneqq\int_{A}r(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$.
-}$ (2.6)
+$\mathbb{E}\left[\int_{0}^{T}\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t+h(\widetilde{X}_{T})\right],\quad\textnormal{with $\tilde{r}(t,x)\coloneqq\int_{A}r(t,x,a)\bm{\pi}(\mathrm{d}a|t,x)$. }$ (2.6)
+In the sequel, we call (2.4) the aggregated dynamics to emphasize that its coefficients are obtained by averaging the coefficients of (2.1) with respect to the sampling distribution defined by $\bm{\pi}$.
 The dynamics (2.4) and the objective (2.6) are derived heuristically in [42] under the assumption that the agent interacts with the system (2.1) by continuously sampling independent random actions from the policy $\bm{\pi}$.
-As a result of these random actions, a “law-of-large-numbers” heuristic suggests that the infinitesimal change in the state at time $t$ is determined by the coefficients in (2.5), aggregated over the sampling distribution defined by $\bm{\pi}$.
-Likewise, the instantaneous reward is also aggregated with respect to this sampling distribution.
-In the sequel, we refer to (2.4) as the aggregated dynamics to highlight this aggregating effect.
-However, the rigorous justification of the connection between (2.4) and the state dynamics resulting from the adoption of a stochastic policy $\bm{\pi}$ in (2.1), as well as the relationship between (2.6) and the original reward functional (2.2), remains to be established.
-Indeed, interacting with the SDE (2.1) by continuously sampling from a stochastic policy creates measure-theoretical issues.
-As already pointed out in [32, Remark 2.1], it is impossible to construct a family of non-constant random variables $(\xi_{t})_{t\in[0,1]}$ such that $(\xi_{t})_{t\in[0,1]}$ is (essentially) pairwise independent and $t\mapsto\xi_{t}$ is Lebesgue measurable.
-This implies that if one controls (2.1) by continuously generating independent actions, the resulting coefficients are not progressively measurable, rendering the conventional stochastic integral ill-defined; see [3] for more discussion.
-Moreover, it is practically infeasible to generate these independent samples continuously.
-Therefore, both from theoretical and practical perspectives, evaluating the performance of a stochastic policy $\bm{\pi}$ requires discretely sampling actions from the policy and applying them to (2.1), as suggested by [32].
-This work quantifies the accuracy of such a discrete execution of a stochastic policy and its impact on the corresponding RL algorithms.
+Under this assumption, the authors determine the state coefficients (2.5) and the instantaneous reward (2.6) by examining the first and second moments of the infinitesimal state change and applying a “law-of-large-numbers” heuristic.
+However, the aggregated dynamics (2.4) differs from the practically observed state process, since realizing it would require continuously sampling independent actions.
+[^3] In fact, interacting with the SDE (2.1) by continuously sampling from a stochastic policy creates measure-theoretical issues.
+As pointed out in [37, Remark 2.1], it is impossible to construct a family of non-constant random variables $(\xi_{t})_{t\in[0,1]}$ such that $(\xi_{t})_{t\in[0,1]}$ is (essentially) pairwise independent and $t\mapsto\xi_{t}$ is Lebesgue measurable.
+This implies that if one controls (2.1) by continuously generating independent actions, the resulting coefficients are not progressively measurable, rendering the conventional stochastic integral ill-defined.
+A more natural and practical implementation of a stochastic policy is to sample actions only at discrete time points and apply them to the state process (2.1) (see [13, 37]).
+The resulting state dynamics is governed by (2.1) with a piecewise constant control process, whose evolution differs from the aggregated dynamics (2.4).
+In the sequel, we show that these state dynamics converge to the aggregated dynamics (2.4) as the sampling frequency increases, and we quantify the corresponding convergence rates.
+We also analyze the impact of this approximation on algorithms designed based on (2.4).
 3. Probabilistic Setup for Aggregated and Sampled Dynamics
-This section introduces a rigorous probabilistic framework for the aggregated dynamics under a given stochastic policy and the associated sampled dynamics with a finite number of intervention times, as outlined in Section 2.
+This section rigorously formulates the aggregated dynamics under a given stochastic policy and the associated sampled dynamics with a finite number of intervention times, as outlined in Section 2.
 3.1. Aggregated dynamics
 Fix a probability space $(\Omega^{W},\mathcal{F}^{W},\mathbb{P}^{W})$ on which a $d$-dimensional Brownian motion $\{W_{t}\}_{t\geq 0}$ is defined, and let ${\mathbb{F}}^{W}\coloneqq\{\mathcal{F}_{t}^{W}\}_{t\geq 0}$ be the $\mathbb{P}^{W}$-completion of the filtration generated by $W$, which satisfies the usual conditions of completeness and right-continuity.
-Throughout this paper, we impose the following regularity condition on the action space $A$, the drift and diffusion coefficients $b$ and $\sigma$, and the stochastic policy $\bm{\pi}$.
+Let the action set $A$ be a nonempty Polish space equipped with the metric $d_{A}$, and let $\mathcal{B}({A})$ be the Borel $\sigma$-algebra on $A$.
+Throughout this paper, we impose the following regularity condition on the drift and diffusion coefficients $b$ and $\sigma$, and the stochastic policy $\bm{\pi}$.
 H. 1.
   1. (1)
-    $({A},d_{A})$ is a Polish space equipped with the Borel $\sigma$-algebra $\mathcal{B}({A})$, and $a_{0}\in A$;
-  2. (2)
     $b:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d}$ and $\sigma:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}^{d\times d}$ are measurable functions.
-There exists $C\geq 0$ such that for all $\varphi\in\{b,\sigma\}$, $t\in[0,T]$, $x,x^{\prime}\in\mathbb{R}^{d}$ and $a\in{A}$,
+There exists $C\geq 0$ such that for a fixed $a_{0}\in A$, any $\varphi\in\{b,\sigma\}$, $t\in[0,T]$, $x,x^{\prime}\in\mathbb{R}^{d}$ and $a\in{A}$,
     $|\varphi(t,x,a)-\varphi(t,x^{\prime},a)|\leq C|x-x^{\prime}|,\quad|\varphi(t,x,a)|$ $\leq C(1+|x|+d_{A}(a,a_{0}));$
@@
     $\bm{\pi}:[0,T]\times\mathbb{R}^{d}\times\mathcal{B}({A})\to[0,1]$ is a probability kernel.
-For all $p\geq 2$, there exists $C_{p}\geq 0$ such that for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$, $\int_{A}d_{A}(a,a_{0})^{p}\bm{\pi}({\mathrm{d}}a|t,x)\leq C_{p}(1+|x|^{p})$.
+For all $p\geq 2$, there exists $C_{p}\geq 0$ such that for all $(t,x)\in[0,T]\times{\mathbb{R}}^{d}$, $\int_{A}d_{A}(a,a_{0})^{p}\bm{\pi}({\mathrm{d}}a|t,x)\leq C_{p}(1+|x|^{p})$, with a fixed $a_{0}\in A$.
 Condition (H.11) imposes the standard Lipschitz regularity on the system coefficients commonly used in the stochastic control literature (see, e.g., [11, 45]), and Condition (H.12) requires mild growth and integrability properties for the stochastic policy.
@@
 Under (H.1), Condition (H.2) holds if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are locally Lipschitz continuous in the state variable (cf.
-[27, Theorem 3.4]) and can be ensured via suitable regularity conditions on $\bm{\pi}$ (e.g., [19, Lemma 2]).
+[31, Theorem 3.4]) and can be ensured via suitable regularity conditions on $\bm{\pi}$ (e.g., [21, Lemma 2]).
 Some typical sufficient conditions on $b,\sigma$ and $\bm{\pi}$ to guarantee these regularity conditions on $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are given in Examples 4.1, 4.2 and 4.3.
@@
 As $A$ is a Polish space, a sampling procedure of a probability kernel $\bm{\pi}$ always exists.
-Indeed, by [21, Lemma 2.22], it suffices to take $E=[0,1]$ and $\xi$ as a uniform random variable on $E$.
+Indeed, by [23, Lemma 2.22], it suffices to take $E=[0,1]$ and $\xi$ as a uniform random variable on $E$.
 This approach has been utilized in [12, 4, 3].
@@
 In our definition, the action sequence $a_{t}$ is a piecewise constant and is given by a sampling function at the grid points $a_{t_{i}}=\phi(t_{i},X_{t_{i}},\xi_{i})$.
-By contrast, in the grid sampling SDE in [3], the action sequence is evolving as $a_{t}=\phi(t,X_{t},\xi_{t_{i}})$, where the random source $\xi_{t_{i}}$ is piecewise constant.
+In contrast, for the grid sampling SDE in [3], the action sequence is evolving as $a_{t}=\phi(t,X_{t},\xi_{t_{i}})$, where the random source $\xi_{t_{i}}$ is piecewise constant.
 in the subsequent analysis.
@@
 Note that the sampled dynamics evolve continuously over time while the control process remains constant within each subinterval.
-In particular, a random action $a_{t_{i}}$ is generated at $t_{i}$ and applied to the system over the interval $[t_{i},t_{i+1})$ before being updated to the next action $a_{t_{i+1}}$.
+In particular, a random action $a_{t_{i}}$ is generated at $t_{i}$ and applied to the system over the interval $t_{i},t_{i+1})$ before being updated to the next action $a_{t_{i+1}}$.
 This aligns with the RL setting, where the agent observes a continuous-time system at discrete time points and adjusts their actions based on these observations.
@@
 Suppose (H.1) holds.
-Let $x_{0}\in{\mathbb{R}}^{d}$, $(\Omega,\mathcal{F},{\mathbb{F}},\mathbb{P})$ be the probability space defined in (3.1), and ${\mathbb{F}}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration such that $\mathcal{F}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{i=0}^{\infty}\}\vee\mathcal{N}$, where $\mathcal{N}$ be the $\mathbb{P}$-null set.
+Let $x_{0}\in{\mathbb{R}}^{d}$, $(\Omega,\mathcal{F},{\mathbb{F}},\mathbb{P})$ be the probability space defined in (3.1), and ${\mathbb{F}}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration such that $\mathcal{F}_{t}\coloneqq\sigma\{(W_{s})_{s\leq t},(\xi_{i})_{i=0}^{\infty}\}\vee\mathcal{N}$, where $\mathcal{N}$ is the $\sigma$-algebra of $\mathbb{P}$-null sets.
 Then for any grid $\mathscr{G}$, (3.3) admits a unique $\mathbb{F}$}-adapted strong solution $X^{\mathscr{G}}$ on the space $(\Omega,\mathcal{F},\mathbb{P})$.
@@
 To highlight the key ideas without introducing needless technicalities, we begin by providing the analysis for sufficiently smooth coefficients and test functions.
-In this setting, we obtain the maximum convergence rate, which is of the order $\mathcal{O}(|\mathscr{G}|)$ as $|\mathscr{G}|$ tends to zero within the regular time grids $\mathscr{P}_{[0,T]}$.
-To specify the precise regularity conditions, we introduce the following notation: for each Euclidean space $E$ and $k\in{\mathbb{N}}$, $p>0$, let $C^{k}_{p}([0,T]\times{\mathbb{R}}^{d};E)$ be the space of functions $u:[0,T]\times{\mathbb{R}}^{d}\to E$ such that for all $r\in{\mathbb{N}}_{0}$ and multi-indices $s$ satisfying $2r+|s|\leq k$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}u$ exists and is continuous for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$, and they all have polynomial growth in $x$:
+In this setting, we obtain the maximum convergence rate of the order $\mathcal{O}(|\mathscr{G}|)$ as $|\mathscr{G}|$ tends to zero.
+To specify the precise regularity conditions, we introduce the following notation: for each Euclidean space $E$ and $k\in{\mathbb{N}}$, $p\in{\mathbb{N}}_{0}$, let $C^{k}_{p}([0,T]\times{\mathbb{R}}^{d};E)$ be the space of functions $u:[0,T]\times{\mathbb{R}}^{d}\to E$ such that for all $r\in{\mathbb{N}}_{0}$ and multi-indices $s$ satisfying $2r+|s|\leq k$, the partial derivative $\partial_{t}^{r}\partial_{x}^{s}u$ exists and is continuous for all $(t,x)\in[0,T]\times\mathbb{R}^{d}$, and they all have polynomial growth in $x$:
 $\|u\|_{C_{p}^{k}}:=\sum_{2r+|s|\leq k}\sup_{(t,x)\in[0,T]\times\mathbb{R}^{d}}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x)\right|}{1+|x|^{p}}<\infty.$ (4.1)
@@
   1. (1)
-    $b$ and $\sigma\sigma^{\top}$ are in the space ${C}^{2,0}_{p}([0,T]\times{\mathbb{R}}^{d}\times A)$;
+    $b$ and $\sigma\sigma^{\top}$ are in the space ${C}^{2,0}_{p}([0,T]\times{\mathbb{R}}^{d}\times A)$ for some $p\in{\mathbb{N}}_{0}$;
   2. (2)
@@
 Remark 4.1.
-Condition (H.3) holds if the coefficients of the aggregated dynamics (2.4) are sufficiently regular.
+To obtain maximal first-order weak convergence, Condition (H.3) requires that both the state coefficients and the solution to the associated Kolmogorov PDE possess sufficient regularity.
+This regularity requirement is stronger than Condition (H.6) assumed for half-order strong convergence.
+This distinction is consistent with the fact that establishing first-order weak convergence of Euler–Maruyama schemes for SDEs typically requires stronger regularity conditions than the Lipschitz continuity that ensures strong convergence (see, e.g., [25, Theorem 14.1.5] and [17] for the conditions of weak and strong convergence of Euler–type schemes, respectively).
+However, as emphasized in Section 1, our convergence result does not follow from existing results on Euler–Maruyama schemes, since the sampled dynamics (3.3) have random coefficients and evolve continuously, and consequently do not arise from a standard Euler–Maruyama discretization of the aggregated dynamics (2.4).
+The PDE solution regularity in Condition (H.3) holds when the coefficients of the aggregated dynamics (2.4) are sufficiently regular.
 For instance, if $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ are in the class $C_{p}^{4}$ and have bounded derivatives, then (H.3) can be verified using the probabilistic argument in [33].
+Alternatively, if the diffusion matrix $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ is uniformly elliptic and $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ are bounded functions in $C^{3}_{p}$ with bounded derivatives, then the desired solution regularity follows from the standard regularity theory of parabolic PDEs; see e.g., [30, Lemma 2.2] and [27, Chapter 4, Theorem 5.1].
+The details are left to the interested reader.
+The following theorem establishes the optimal first-order weak convergence of the sampled dynamics to the aggregated dynamics.
 Theorem 4.1.
@@
 Then for any $f\in C_{p}^{4}(\mathbb{R}^{d})$, there exists a constant $C_{p}\geq 0$ depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},\bm{\pi},b,\sigma,p$ such that for all grids $\mathscr{G}$,
-$\sup_{t\in[0,T]}\left|\mathbb{E}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|\leq C_{p}\|f\|_{C_{p}^{4}}|\mathscr{G}|.$ (4.5)
+$\sup_{t\in[0,T]}\left|\mathbb{E}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}[f(\widetilde{X}_{t})]\right|\leq C_{p}\|f\|_{C_{p}^{4}}|\mathscr{G}|.$ (4.5)
 Proof.
 See Section 6.1.1. ∎
-For a given time grid $\mathscr{G}$, we execute the policy only at the nodes $t_{i}$, and use the same action over the time interval $[t_{i},t_{i+1})$ to obtain the sampled dynamics $X^{\mathscr{G}}$.
-Theorem 4.1 illustrates that with this policy execution method, for a given test function $f$, the error between $\mathbb{E}[f(X^{\mathscr{G}}_{t})]$ and the original aggregated dynamics $\mathbb{E}^{W}[f(\widetilde{X}_{t})]$ converges to 00 at the optimal first-order convergence $O(|\mathscr{G}|)$.
+By Theorem 4.1, for a given function $f\in C_{p}^{4}(\mathbb{R}^{d})$, $f(X^{\mathscr{G}}_{T})$ is a biased estimator of $\mathbb{E}[f(\widetilde{X}_{T})]$, and the bias vanishes at first order, as the sampling mesh size $|\mathscr{G}|$ tends to zero.
+Moreover, by Lemma 3.1, the variance of the estimator $f(X^{\mathscr{G}}_{T})$ is uniformly bounded, independent of the mesh size $|\mathscr{G}|$.
+The computational cost of generating one trajectory of $X^{\mathscr{G}}_{T}$ increases linearly with the number of grid points in $\mathscr{G}$.
+For further details on its implications for the sample complexity of Monte Carlo estimators of $\mathbb{E}[f(\widetilde{X}_{t})]$, we refer to Remark 4.3.
 We illustrate several common cases in which the required assumptions in Theorem 4.1 hold.
@@
 Definition 4.1.
-For $l\in(0,\infty)\setminus\mathbb{N}$, let $\mathcal{H}_{T}^{(l)}$ be the space of functions $u:[0,T]\times\mathbb{R}^{d}\times{A}\to\mathbb{R}$ such that for all integer $r\geq 0$ and multi-index $s$ satisfying $2r+|s|\leq l$, the partial derivatives $\partial_{t}^{r}\partial_{x}^{s}u$ are bounded and Hölder continuous with index $l-[l]$ in $x$ and with index $(l-2r-|s|)/2$ in $t$, i.e.,
-$\|u\|_{T}^{(l)}:=$ $\sum_{2r+|s|\leq[l]}\sup_{t,x,a}|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)|$ (4.11)
+For $l\in(0,\infty)\setminus\mathbb{N}$, let $\mathcal{H}_{T}^{(l)}$ be the space of functions $u:[0,T]\times\mathbb{R}^{d}\times{A}\to{\color[rgb]{0,0,0}E}$ such that for all integer $r\geq 0$ and multi-index $s$ satisfying $2r+|s|\leq l$, the partial derivatives $\partial_{t}^{r}\partial_{x}^{s}u$ are bounded and Hölder continuous with index l[l]l-[l] in $x$ and with index (l2r|s|)/2(l-2r-|s|)/2 in $t$, i.e.,
+$\|u\|_{T}^{(l)}$= $\sum_{2r+|s|\leq[l]}\sup_{t,x,a}|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)|$ (4.11)
 $+\sum_{2r+|s|=[l]}\sup_{t,a,x\neq x^{\prime}}\frac{\left|\partial_{t}^{r}\partial_{x}^{s}u(t,x,a)-\partial_{t}^{r}\partial_{x}^{s}u(t,x^{\prime},a)\right|}{\left|x-x^{\prime}\right|^{l-[l]}}$
@@
 H. 4.
-Let $l\in(0,1)\cup(1,2)\cup(2,3)$, $b,\sigma\sigma^{\top}\in\mathcal{H}_{T}^{(l)}$, $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}\in\mathcal{H}_{T}^{(l)}$, and $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ is uniformly elliptic.
-Under (H.4), standard regularity theory for parabolic PDEs (see, e.g., [24, Chapter 4, Theorem 5.1]) shows that for any $f\in\mathcal{H}^{(l+2)}$ and $t^{\prime}\in[0,T]$, the PDE (4.4) has a unique solution $v_{f}(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v_{f}(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where the constant $C$ depends only on $T,\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$.
+Let $l\in(0,1)\cup(1,2)\cup(2,3)$, $b,\sigma\sigma^{\top}\in\mathcal{H}_{T}^{(l)}$, $\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}\in\mathcal{H}_{T}^{(l)}$, and $\tilde{\sigma}^{\bm{\pi}}(\tilde{\sigma}^{\bm{\pi}})^{\top}$ is uniformly elliptic.
+Under (H.4), standard regularity theory for parabolic PDEs (see, e.g., [27, Chapter 4, Theorem 5.1]) shows that for any $f\in\mathcal{H}^{(l+2)}$ and $t^{\prime}\in[0,T]$, the PDE (4.4) has a unique solution $v_{f}(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v_{f}(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where the constant $C$ depends only on $T,\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$.
 Leveraging this PDE regularity, the following theorem quantifies the weak convergence order of the sampled dynamics in terms of the smoothness coefficient $l$.
@@
 Compared with (H.3), (H.5) imposes the additional condition on the uniform boundedness $\sup_{a\in A}\|b(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, and $\sup_{a\in A}\|\sigma\sigma^{\top}(\cdot,\cdot,a)\|_{C_{p}^{2}}<\infty$, which allows for controlling the concentration behavior of the sampled dynamics with respect to the execution noise $\xi$.
-It is easy to see that (H.4) implies (H.5).
-The following theorem shows that the weak convergence rate in expectation, established in Section 4.1, also holds in a high probability sense with respect to the sampling noises.
-It extends a similar result established in [32] for LQ problems with Gaussian policies to general state dynamics and stochastic policies.
+It is easy to see that (H.4) includes the uniform boundedness assumption on $a$, thus, if (H.4) holds, then we do not need (H.5).
+The following theorem establishes that the sampled dynamics converge weakly at half-order, averaged over the system noise, uniformly over the realizations of the sampling noises with high probability.
+It extends [37, Theorem 2.2] established for LQ problems with Gaussian policies to general state dynamics and stochastic policies.
 Theorem 4.3.
 Suppose (H.1), (H.2), (H.3) and (H.5) hold.
-Then for the test function $f$ in Theorem 4.1, there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$, such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
-$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|>C_{p}\|f\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right]\leq\varrho.$ (4.15)
-Assume further that (H.1), (H.2) and (H.4) hold.
-Then for the test function $f$ in Theorem 4.2, there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
-$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}^{W}[f(\widetilde{X}_{t})]\right|>C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right]\leq\varrho,$ (4.16)
+Then for all $f\in C^{4}_{p}({\mathbb{R}}^{d})$, there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$, such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
+$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}[f(\widetilde{X}_{t})]\right|>C_{p}\|f\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big(1+\sqrt{\ln(2/\varrho)}\big)\right]\leq\varrho.$ (4.15)
+Suppose instead that (H.1), (H.2) and (H.4) hold.
+Then for all $f\in\mathcal{H}^{(l+2)}$ with $l$ in (H.4), there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for all grids $\mathscr{G}$ and all $\varrho\in(0,1)$,
+$\sup_{t\in[0,T]}\mathbb{P}\left[\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t})]-\mathbb{E}[f(\widetilde{X}_{t})]\right|>C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big(1+\sqrt{\ln(2/\varrho)}\big)\right]\leq\varrho,$ (4.16)
 with $\mathcal{X}(l)$ defined by (4.13).
-Proof.
-See Section 7.1.3. ∎
 The proof of Theorem 4.3, provided in Section 6.1.3, is based on establishing a pointwise upper bound for the weak error on each subinterval in terms of the grid size, and subsequently controlling the overall error using concentration inequalities for bounded martingale difference sequences.
+Note that Theorem 4.3 shows that the convergence of the sampled dynamics to the aggregated dynamics can be achieved simply by increasing the sampling frequency, without the need to average over multiple realizations of the sampling noise at fixed time points.
+This observation significantly reduces the sample complexity of estimators for the aggregated dynamics.
+To see it, consider estimating $\mathbb{E}[f(\widetilde{X}_{T})]$ using the sampled dynamics (3.2).
+Let $\mathscr{G}_{n}=\{i\frac{T}{n}\}_{i=0}^{n}$, $n\in{\mathbb{N}}$, be a uniform time grid, and let $(W_{k})_{k\in{\mathbb{N}}}$ be a collection of independent $d$-dimensional Brownian motions, which are also independent of $\xi$.
+For each $m\in{\mathbb{N}}$, define the estimator $\hat{I}_{m,n}$ of $\mathbb{E}[f(\widetilde{X}_{T})]$ by
+$\hat{I}_{m,n}\coloneqq\frac{1}{m}\sum_{k=1}^{m}f(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)),$ (4.17)
+where $X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)$ denotes the terminal state of the sampled dynamics (3.2), driven by the Brownian motion $W_{k}$ and the common sampling noises $\xi$ on the grid $\mathscr{G}_{n}$.
+We define the complexity of $\hat{I}_{m,n}$ as the total number of random variables, including the Brownian motions $(W_{k})_{k\in{\mathbb{N}}}$ and the sampling noises $(\xi_{n})_{n\in{\mathbb{N}}_{0}}$, required to evaluate a single realization of $\hat{I}_{m,n}$.
+Corollary 4.1.
+Suppose (H.1), (H.2), (H.3) and (H.5) hold.
+Let $f\in C^{4}_{p}({\mathbb{R}}^{d})$ be bounded.
+There is a constant $C\geq 0$ such that for all $\varepsilon\in(0,1)$, by setting $m_{\varepsilon}=n_{\varepsilon}=\lceil\varepsilon^{-2}\left(1+\ln(2/\varepsilon)\right)\rceil$, it holds with probability at least $1-\varepsilon$ that $|\hat{I}_{m_{\varepsilon},n_{\varepsilon}}-\mathbb{E}[f(\widetilde{X}_{T})]|\leq C\varepsilon$.
+Moreover, the sample complexity of $\hat{I}_{m_{\varepsilon},n_{\varepsilon}}$ is $2\lceil\varepsilon^{-2}\left(1+\ln(2/\varepsilon)\right)\rceil$.
+The proof of Corollary 4.1 is given in Section 6.2.
+Similar results hold when the objective involves a sufficiently regular running reward of the aggregated process, as in the cases analyzed in Section 5.
+Remark 4.3.
+Note that in the idealized scenario where one can directly sample trajectories of the aggregated dynamics, a Monte Carlo estimator of $\mathbb{E}[f(\widetilde{X}_{T})]$ achieves a sample complexity of order $\mathcal{O}(\varepsilon^{-2})$.
+As shown in Corollary 4.1, the estimator (4.17) attains this optimal sample complexity up to a logarithmic factor.
+This is due to the fact that the estimator (4.17) does not regenerate sampling noises for different Brownian motion trajectories.
+In contrast, the weak convergence results in Theorems 4.1 and 4.2 suggest considering the naive Monte Carlo estimator
+$\tilde{I}_{m,n}\coloneqq\frac{1}{m}\sum_{k=1}^{m}f(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi_{k})),$ (4.18)
+where $(\xi_{k})_{k\in{\mathbb{N}}}$ are independent copies of the sampling noises $\xi$.
+Under the same conditions as in Corollary 4.1, the standard bias–variance decomposition shows that for all $m,n\in{\mathbb{N}}$,
+${\mathbb{E}}[|\tilde{I}_{m,n}-\mathbb{E}[f(\widetilde{X}_{T})]|^{2}]^{1/2}\leq C\left(\frac{1}{\sqrt{m}}+\frac{1}{n}\right),$
+for a constant $C\geq 0$ independent of $m$ and $n$.
+To achieve a total error $\varepsilon$, we choose $m_{\varepsilon}=\mathcal{O}(\varepsilon^{-2})$ and $n_{\varepsilon}=\mathcal{O}(\varepsilon^{-1})$.
+However, since the estimator (4.18) requires regenerating sampling noises for each trajectory, the overall complexity is $\mathcal{O}(m_{\varepsilon}n_{\varepsilon})=\mathcal{O}(\varepsilon^{-3})$, which is significantly higher than that of the estimator (4.17).
 4.3. Strong convergence
@@
   1. (1)
-    There exists $C\geq 0$ such that for all $(t_{1},x_{1},a_{1}),(t_{2},x_{2},a_{2})\in[0,T]\times\mathbb{R}^{d}\times A$, $u\in E$,
-    $\left|b(t_{1},x_{1},a_{1})-b(t_{2},x_{2},a_{2})\right|+d_{A}(\phi(t_{1},x_{1},u),\phi(t_{2},x_{2},u))\leq C\left(\left|t_{1}-t_{2}\right|^{\frac{1}{2}}+\left|x_{1}-x_{2}\right|+d_{A}(a_{1},a_{2})\right),$ (4.17)
-    and for all $x,y\in\mathbb{R}^{d}$ and $t\in[0,T]$,
-    $\left|\sigma(t,x)-\sigma(t,y)\right|\leq C\left|x-y\right|.$ (4.18)
+    There exists $C\geq 0$ such that for all $x,y\in\mathbb{R}^{d}$, $a_{1},a_{2}\in A$, and $t\in[0,T]$,
+    $\left|b(t,x,a_{1})-b(t,y,a_{2})\right|+\left|\sigma(t,x)-\sigma(t,y)\right|$ $\leq C\left(\left|x-y\right|+d_{A}(a_{1},a_{2})\right),$
+    $\left|b(t,0,0)\right|+\left|\sigma(t,0)\right|$ $\leq C.$
   2. (2)
-    There exists $(t_{0},x_{0},a_{0})\in[0,T]\times\mathbb{R}^{d}\times A$, such that $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential  [^3] We say a random variable $X$ is sub-exponential if there exists a constant $C>0$ such that $\mathbb{E}[|X|>t]\leq 2e^{-Ct}$ for every $t\geq 0$..
+    There exists $(t_{0},x_{0},a_{0})\in[0,T]\times\mathbb{R}^{d}\times A$, such that $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential.
+[^5] We say a random variable $X$ is sub-exponential if there exists a constant $C>0$ such that $\mathbb{E}[|X|>t]\leq 2e^{-Ct}$ for every $t\geq 0$.
+Moreover, there exists $C\geq 0$ such that for all $(t_{1},x_{1}),(t_{2},x_{2})\in[0,T]\times\mathbb{R}^{d}$, and $u\in E$,
+    $d_{A}(\phi(t_{1},x_{1},u),\phi(t_{2},x_{2},u))\leq C\left(\left|t_{1}-t_{2}\right|^{\frac{1}{2}}+\left|x_{1}-x_{2}\right|\right).$ (4.19)
+Assumption (H.62) holds for commonly used Gaussian policies considered in [37] and many others.
 Example 4.5.
-If $\bm{\pi}(\cdot|t,x)$ satisfies (4.7) with $\Sigma(t,x)=\Sigma$, then we can write $a=\phi(t,x,\xi)=\mu(t,x)+\Sigma\xi$.
-In this case, the assumptions for $\phi$ in (H.6) hold.
-This corresponds to the commonly used Gaussian policies considered in [32] and many others.
+Let $A={\mathbb{R}}^{k}$, and $\bm{\pi}(\cdot|t,x)$ be a Gaussian policy of the form (see e.g., [37]):
+${\bm{\pi}(\mathrm{d}a|t,x)}=\varphi(a|\mu(t,x),\Sigma^{2})\mathrm{d}a,$ (4.20)
+where $\varphi(\cdot|\mu,\Sigma^{2})$ denotes the Gaussian density with mean $\mu$ and variance $\Sigma^{2}$.
+In this case, one can define the sampling function $\phi(t,x,\xi)=\mu(t,x)+\Sigma\xi$, with a standard normal random variable $\xi$.
+If $\mu$ is Lipschitz continuous in $x$ and $1/2$-Hölder continuous in $t$, then $\phi$ satisfies (H.62).
+More generally, Assumption (H.62) holds for stochastic policies with sufficiently regular densities.
+Example 4.6.
+Let $A=[\alpha,\beta]$ for $\alpha<\beta$, and $\bm{\pi}({\mathrm{d}}a|t,x)=h(t,x,a){\mathrm{d}}a$ with a continuous density $h:[0,T]\times{\mathbb{R}}^{d}\times A\to{\mathbb{R}}$.
+In this case, by the inverse transform method, one can sample from $\bm{\pi}({\mathrm{d}}a|t,x)$ by $\phi(t,x,\xi)$, where $\xi$ is a uniform random variable on $[0,1]$, and $0,1]\ni u\mapsto\phi(t,x,u)\in A$ is the inverse function of the map $A\ni r\mapsto H(t,x,r)\coloneqq\int_{\alpha}^{r}h(t,x,a){\mathrm{d}}a\in[0,1]$.
+If there exists $C,c>0$ such that for all $(t,x),(t^{\prime},x^{\prime})\in[0,T]\times{\mathbb{R}}^{d}$ and $a\in A$,
+$h(t,x,a)\geq c,\quad|h(t,x,a)-h(t^{\prime},x^{\prime},a)|\leq C(|t-t^{\prime}|^{\frac{1}{2}}+|x-x^{\prime}|),$ (4.21)
+then $\phi$ satisfies (H.62).
+Indeed, $|\phi(t_{0},x_{0},\xi)|$ is sub-exponential since $A$ is bounded. Moreover, since $\phi$ is the inverse of $H$,
+$H(t,x,\phi(t,x,u))-H(t,x,\phi(t^{\prime},x^{\prime},u))+H(t,x,\phi(t^{\prime},x^{\prime},u))-H(t^{\prime},x^{\prime},\phi(t^{\prime},x^{\prime},u))$
+$=H(t,x,\phi(t,x,u))-H(t^{\prime},x^{\prime},\phi(t^{\prime},x^{\prime},u))=u-u=0,$
+which along with the mean value theorem and the definition of $H$ yields that
+$\phi(t,x,u)-\phi(t^{\prime},x^{\prime},u)=\frac{\int_{\alpha}^{\phi(t^{\prime},x^{\prime},u)}(h(t^{\prime},x^{\prime},a)-h(t,x,a)){\mathrm{d}}a}{\partial_{r}H(t,x,\tilde{a})}\leq\frac{\int_{\alpha}^{\beta}|h(t^{\prime},x^{\prime},a)-h(t,x,a)|{\mathrm{d}}a}{h(t,x,\tilde{a})}$
+for some $\tilde{a}$ between $\phi(t,x,u)$ and $\phi(t^{\prime},x^{\prime},u)$. Using (4.21) shows that $\phi$ satisfies (4.19).
+Similar arguments extend to multidimensional action sets $A$ using the conditional inverse sampling method (see e.g., [8, Chapter 6.3]).
+Now we prove the strong convergence of the sampled dynamics under Assumption (H.6).
+Note that (H.6) implies (H.1) and (H.2), which ensure that the aggregated state process $\widetilde{X}$ and the sampled state process $X^{\mathscr{G}}$ are $p$-th power integrable for all $p\geq 2$ (see Lemma 6.5).
 Theorem 4.4.
-Suppose (H.2) and (H.6) hold. For all $p\geq 2$ and all grids $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ with $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}<\infty$,
+Suppose (H.6) holds. For all $p\geq 2$ and all grids $\mathscr{G}=\{0=t_{0}<\ldots<t_{n}=T\}$ with $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}<\infty$,
 $\mathbb{E}\left[\sup_{0\leq t\leq T}\left|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}\right|^{p}\right]\leq C|\mathscr{G}|^{\frac{p}{2}},$ (4.22)
@@
 $\sup_{0\leq t\leq T}\left|X^{\mathscr{G}}_{t}(\omega)-\widetilde{X}_{t}(\omega)\right|\leq n^{-\frac{1}{2}+\varepsilon}.$ (4.23)
+Proof.
+See Section 6.1.4. ∎
 The proof of Theorem 4.4 is given in Section 6.1.4.
-It involves carefully estimating the Orlicz norm of the pathwise difference between the aggregated and sampled dynamics on each subinterval and subsequently controlling the overall error using concentration inequalities for sub-exponential martingale difference sequences.
+The analysis involves carefully estimating the Orlicz norm of the pathwise difference between the aggregated and sampled dynamics on each subinterval by conditioning on the realization of the Brownian motion, and then controlling the overall error using concentration inequalities for conditional sub-exponential martingale difference sequences.
 Theorem 4.4 indicates that the convergence of the sampled dynamics to the aggregated dynamics can be interpreted as a law of large numbers in time.
 Specifically, when the volatility is uncontrolled, the $1/2$-order convergence is achieved by increasing the sampling frequency for any fixed realization of the system noise.
-Unfortunately, such a strong convergence fails when the volatility is controlled, as shown in the following example.
+However, such a pathwise convergence fails when the volatility is controlled, as shown in the following example.
 Example 4.7.
@@
 Let $a_{t_{i}}=\xi_{i}$ for all $i$, then we have
-$X^{\mathscr{G}}_{t}-\widetilde{X}_{t}=\int_{0}^{t}\left(a_{\delta(s)}-Q^{1/2}\right)\mathrm{d}W_{s}.$ (4.21)
+$X^{\mathscr{G}}_{t}-\widetilde{X}_{t}=\int_{0}^{t}\left(a_{\delta(s)}-\sqrt{Q}\right)\mathrm{d}W_{s}.$ (4.24)
 Note that $W$ is a Brownian motion under the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ with $\mathcal{F}_{t}:=\sigma\{a_{\delta(s)}:s\leq t\}\lor\mathcal{F}^{W}_{t}$.
-It follows from Itô’s isometry that
-$\mathbb{E}\left[|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}|^{2}\right]=\mathbb{E}\left[\int_{0}^{t}(a_{\delta(s)}-Q^{1/2})^{2}\mathrm{d}s\right]=t\mathbb{E}\left[(\xi_{1}-Q^{1/2})^{2}\right],$ (4.22)
+By Itô’s isometry and Fubini’s theorem,
+$\mathbb{E}\left[|X^{\mathscr{G}}_{t}-\widetilde{X}_{t}|^{2}\right]=\mathbb{E}\left[\int_{0}^{t}(a_{\delta(s)}-\sqrt{Q})^{2}\mathrm{d}s\right]=t\mathbb{E}\left[(\xi_{1}-\sqrt{Q})^{2}\right],$ (4.25)
 which does not converge to zero as $|\mathscr{G}|\rightarrow 0$, except that $\xi_{1}$ is a constant (or equivalently $\mu$ is a Dirac distribution).
+4.4. Numerical illustration
+We present numerical results illustrating the theoretical strong and weak convergence rates for two representative one-dimensional cases: uncontrolled volatility $\mathrm{d}X_{t}=a_{t}\mathrm{d}t+\mathrm{d}W_{t}$ and controlled volatility $\mathrm{d}X_{t}=a_{t}\mathrm{d}W_{t}$.
+In both cases, we set the time horizon $[0,1]$ and the initial state $X_{0}=0$, and consider sampling actions from a standard Gaussian policy (i.e., $a_{t}\sim\mathcal{N}(0,1)$) over a uniform grid $\mathscr{G}=\{i\frac{T}{n}\}_{i=0}^{n}$, where $n\in{\mathbb{N}}$ is the number of grid points.
+We consider the quantity $|\mathbb{E}[(X^{\mathscr{G}}_{T})^{4}]-\mathbb{E}[(\widetilde{X}_{T})^{4}]|$ to examine the weak convergence of the sampled dynamics, and the root mean squared error (RMSE) $\sqrt{\mathbb{E}[|X^{\mathscr{G}}_{T}-\widetilde{X}_{T}|^{2}]}$ to assess the strong convergence.
+All expectations are estimated using the Monte Carlo method with 50000 samples.
+Each point in the following plots is an average over 100 independent runs, and the shaded regions represent the 95% confidence intervals, estimated using 1000 bootstrap resamples.
+Figure 1.
+Weak and strong convergence analysis for the uncontrolled volatility case $\mathrm{d}X_{t}=a_{t}\mathrm{d}t+\mathrm{d}W_{t}$ with actions $a_{t}\sim\mathcal{N}(0,1)$.
+The corresponding aggregated dynamics is $\mathrm{d}\tilde{X}_{t}=\mathrm{d}W_{t}$.
+The test function is $f(x)=x^{4}$ and $T=1$.
+Left: Weak error versus the number of grid points $n$.
+Right: RMSE versus the number of grid points $n$.
+Both axes are on log scales.
+Figure 1 presents the convergence results for the case with uncontrolled volatility.
+The left panel shows that the weak error decays linearly with the sampling frequency on a log-log scale, with a slope of approximately 1-1.
+This confirms the first-order convergence rate established in Theorem 4.1.
+The right panel shows that the RMSE also decays linearly on the log-log plot, with a slope of approximately 1/2-1/2, confirming the strong convergence rate of $O(|\mathscr{G}|^{1/2})$ as stated in Theorem 4.4.
+The confidence interval for the strong error is not visible due to the very small variance of the estimator.
+Figure 2.
+Weak and strong convergence analysis for the controlled volatility case $\mathrm{d}X_{t}=a_{t}\mathrm{d}W_{t}$ with actions $a_{t}\sim\mathcal{N}(0,1)$.
+The corresponding aggregated dynamics is $\mathrm{d}\tilde{X}_{t}=\mathrm{d}W_{t}$.
+The test function is $f(x)=x^{4}$ and $T=1$.
+Left: Weak error versus the number of grid points $n$.
+Right: RMSE versus the number of grid points $n$.
+Both axes are on log scales.
+Figure 2 presents the convergence results for the second case, where the volatility is controlled.
+Similar to the first case with uncontrolled volatility, the left panel shows that the weak error exhibits a first-order decay with a slope of 1-1, consistent with Theorem 4.1.
+In contrast, the right panel, which depicts strong convergence, shows that the RMSE remains essentially constant as the number of grid points $n$ increases.
+This provides a clear numerical illustration of the failure of strong convergence under controlled volatility, confirming the theoretical argument presented in Example 4.7.
 5. Applications
@@
 Therefore, the agent can sample the process $X^{\mathscr{G}}_{s}$.
-For simplicity, we assume that the agent can collect and observe the accumulated running reward $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ [^4] If only the instantaneous reward $r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})$ can be observed at the grid point, then the accumulated reward need to be further approximated as $R_{t_{i}}^{\mathscr{G}}\approx r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})(t_{i}-t_{i-1})$ as in the usual Euler scheme.
+For simplicity, we assume that the agent can collect and observe the accumulated running reward $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$.
+[^6] If only the instantaneous reward $r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})$ can be observed at the grid point, then the accumulated reward needs to be further approximated as $R_{t_{i}}^{\mathscr{G}}\approx r(s,X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})(t_{i}-t_{i-1})$ as in the usual Euler scheme.
 Under suitable assumptions on the regularity conditions on the function $r$, the order of our analysis does not change.
@@
 $J^{\mathscr{G}}\coloneqq\mathbb{E}\left[\int_{0}^{T}r(s,X^{\mathscr{G}}_{s},a_{\delta(s)})\mathrm{d}s+h(X^{\mathscr{G}}_{T})\right]=\mathbb{E}\left[\sum_{i=0}^{n-1}R_{t_{i}}^{\mathscr{G}}+h(X^{\mathscr{G}}_{T})\right].$ (5.1)
-In, the expected reward with a continuous-time execution of the policy is given by (2.6):
+In contrast, the expected reward with a continuous-time execution of the policy is given by (2.6):
 $\tilde{J}\coloneqq\mathbb{E}\left[\int_{0}^{T}\tilde{r}(s,\widetilde{X}_{s})\mathrm{d}s+h(\widetilde{X}_{T})\right]=\mathbb{E}\left[\int_{0}^{T}\int_{A}r(s,\widetilde{X}_{s},a)\bm{\pi}(\mathrm{d}a|s,\widetilde{X}_{s})\mathrm{d}s+h(\widetilde{X}_{T})\right].$ (5.2)
@@
   2. (2)
-[…]
+    If (H.4) holds, and $r\in\mathcal{H}_{T}^{(l)},\tilde{r}\in\mathcal{H}_{T}^{(l)},h\in\mathcal{H}^{(l+2)}$ and $\tilde{r}(t,\cdot)\in\mathcal{H}^{(l+2)},\ 0\leq t\leq T$, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that
     $\left|J^{\mathscr{G}}-\tilde{J}\right|\leq C\left(\|h\|^{(l+2)}+\|r\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|^{(l+2)}+\|\tilde{r}\|_{T}^{(l)}\right)|\mathscr{G}|^{\mathcal{X}(l)},$ (5.4)
@@
 Assume (H.1) and (H.2) hold.
-If (H.5) and the conditions in Proposition 5.1(1) hold, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$ such that for every $\varrho\in(0,1)$,
-$\mathbb{P}\left[|J^{\mathscr{G}}(\xi)-\tilde{J}|>C_{p}\left(\|h\|_{C_{p}^{4}}+\|r\|_{C_{p}^{2,0}}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|_{C_{p}^{4}}+\|\tilde{r}\|_{C_{p}^{2}}\right)\sqrt{|\mathscr{G}|}(1+\sqrt{\ln(2/\varrho)})\right]\leq\varrho.$ (5.6)
+If (H.5) and the conditions in Proposition 5.1(1) hold, and $\sup_{a\in A}\|r(\cdot,\cdot,a)\|_{C_{p}^{2}}<+\infty$, then there is a constant $C_{p}$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T,p$ such that for every $\varrho\in(0,1)$,
+$\mathbb{P}\left[|J^{\mathscr{G}}(\xi)-\tilde{J}|>C_{p}\left(\|h\|_{C_{p}^{4}}+\sup_{a\in A}\|r(\cdot,\cdot,a)\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\tilde{r}(t,\cdot)\|_{C_{p}^{4}}+\|\tilde{r}\|_{C_{p}^{2}}\right)\sqrt{|\mathscr{G}|}(1+\sqrt{\ln(2/\varrho)})\right]\leq\varrho.$ (5.6)
 If conditions in Proposition 5.1(2) holds, then there is a constant $C$ depending only on $b,\sigma,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}},T$ such that for every $\varrho\in(0,1)$,
@@
 In particular, suppose $V^{\theta}$ is a family of parameterized functions of $(t,x)$ with $\theta$ a finite-dimensional parameter.
-TD learning concerns the increment in the form “$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” and the instantaneous reward “$r(t,\widetilde{X}_{t})\mathrm{d}t$”.
+TD learning concerns the increment in the form $\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” and the instantaneous reward $r(t,\widetilde{X}_{t})\mathrm{d}t$”.
 In the continuous-time framework, these two increments have different natures.
-“$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” is merely an informal notation that shall be combined with a suitable integrand as in the definition of Itô’s integral, and the latter “$r(t,\widetilde{X}_{t})\mathrm{d}t$” is the usual Lebesgue integral with respect to time.
+$\mathrm{d}V^{\theta}(t,\widetilde{X}_{t})$” is merely an informal notation that shall be combined with a suitable integrand as in the definition of Itô’s integral, and the latter $r(t,\widetilde{X}_{t})\mathrm{d}t$” is the usual Lebesgue integral with respect to time.
 In theory, [21] identify the continuous-time counterpart of TD learning as solving the martingale orthogonality condition.
@@
 We denote the density function by $H^{\psi}(t,x,a)=\frac{\bm{\pi}^{\psi}(\mathrm{d}a|t,x)}{\mu(\mathrm{d}a)}$.
-Then we know the value function under the policy $\bm{\pi}^{\psi}$ (defined in (5.8)) will depend on $\psi$ implicitly, denoted by $\tilde{J}(t,x;\bm{\pi}^{\psi})$.
+Then we know that the value function under the policy $\bm{\pi}^{\psi}$ (defined in (5.8)) will depend on $\psi$ implicitly, denoted by $\tilde{J}(t,x;\bm{\pi}^{\psi})$.
 The policy gradient concerns about the derivative of $\tilde{J}(t,x;\bm{\pi}^{\psi})$ with respect to $\psi$, i.e., $\frac{\partial\tilde{J}(t,x;\bm{\pi}^{\psi})}{\partial\psi}$.
@@
 $S\circ V:=\sum_{i=0}^{n-1}S(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\left(V(t_{i+1},X^{\mathscr{G}}_{t_{i+1}})-V(t_{i},X^{\mathscr{G}}_{t_{i}})+R_{t_{i}}^{\mathscr{G}}\right),$ (5.15)
-where $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ is the observed accumulated reward over $[t_{i},t_{i+1}]$.
+where $R_{t_{i}}^{\mathscr{G}}=\int_{t_{i}}^{t_{i+1}}r(s,X^{\mathscr{G}}_{s},a_{t_{i}})\mathrm{d}s$ is the observed accumulated reward over $t_{i},t_{i+1}]$.
 The form (5.15) also appears in the continuous-time q-learning algorithm proposed in [22], where $S$ is the derivative of the q-function that needs to be learned.
@@
 where $t\mapsto\left<V(\cdot,\widetilde{X}_{\cdot})\right>_{t}$ is the quadratic variation of the process $t\mapsto V(t,\widetilde{X}_{t})$.
-6. Conclusion
-This paper addresses a fundamental problem in the continuous-time RL framework proposed in [36] by answering the relation between the aggregated dynamics (referred to as the “exploratory dynamics” therein) and the actual sampling process.
-In terms of SDE theory, we provide a rigorous treatment of the “law-of-large-numbers” heuristics used in [36] to demonstrate how the original dynamics lead to the aggregated dynamics.
-Our analysis is particularly crucial for understanding how discrete sampling relates to the continuous-time RL problems.
-We establish the convergence and convergence rate of discretely sampled stochastic policies, i.e., the sampled dynamics, to the aggregated dynamics in different senses.
-Furthermore, we illustrate the bias and variance of the estimate for several important quantities in analyzing continuous-time RL algorithms when using sampled dynamics, including values, temporal differences, policy gradients, and quadratic variations.
-These bias and variance analyses are critical for the convergence of stochastic approximation algorithms, and they provide guidance on properly tuning the interaction frequency to foster faster convergence of RL algorithms.
-The results of this paper also stimulate more interesting open questions for future research.
-For example, other sampling methods that could achieve higher accuracy in the RL context, any exact simulation methods for controlled SDEs under stochastic policies if the coefficients are all known/unknown, and a full account for the convergence of the gradient-based continuous-time RL algorithm.
 6. Proofs of Main Results
@@
 6.1.1. Proof of Theorem 4.1
-Proof.
+The following lemma can be obtained by using Itô’s formula and Lemma 3.1. The proof is given in Appendix A.
+Lemma 6.1.
+Suppose (H.1) holds, and $g\in C_{p}^{2,0}$. Then there is a constant $C\geq 0$ such that for all grids $\mathscr{G}$ and $s\in[0,T]$,
+$\left|\mathbb{E}\left[g(s,X^{\mathscr{G}}_{s},a_{\delta(s)})-g(\delta(s),X^{\mathscr{G}}_{\delta(s)},a_{\delta(s)})\right]\right|\leq C\|g\|_{C_{p}^{2,0}}|\mathscr{G}|.$ (6.1)
+Based on Lemma 6.1, we prove Theorem 4.1.
+Proof of Theorem 4.1.
 Without loss of generality, assume $t^{\prime}=t_{j}$, where $1\leq j\leq n$. Consider the following PDE
-$\frac{\partial}{\partial t}v(t,x)+\mathcal{L}v(t,x)=0,\quad t\in[0,t_{j}]\times{\mathbb{R}}^{d};\quad v(t_{j},x)=f(x),\quad x\in{\mathbb{R}}^{d}.$ (7.1)
-Let $v_{f}(t,x;t_{j})$ be the solution of (7.1).
-For simplicity, we dorp the $f$ and $t_{j}$ and use the notation $v(t,x)$.
-By (H.
-3), $v(\cdot,\cdot)$ is smooth and $\|v(\cdot,\cdot)\|_{C_{p}^{4}}<+\infty$.
-Noting that if $v$ is the solution of (7.1), then by the Feynman-Kac formula, $\mathbb{E}[f(\widetilde{X}_{t_{j}})]=v(0,x_{0})$. With this relation, we have
+$\frac{\partial}{\partial t}v(t,x)+\mathcal{L}v(t,x)=0,\quad t\in[0,t_{j}]\times{\mathbb{R}}^{d};\quad v(t_{j},x)=f(x),\quad x\in{\mathbb{R}}^{d},$ (6.2)
+where $\mathcal{L}$ is the generator of the aggregated dynamic $\widetilde{X}$ (2.4) and is defined by (4.3).
+Denote by $v:[0,t_{f}]\to{\mathbb{R}}$ the solution of (6.2).
+By (H.3), $v$ is smooth and $\|v\|_{C_{p}^{4}}<+\infty$.
+By the Feynman-Kac formula, $\mathbb{E}[f(\widetilde{X}_{t_{j}})]=v(0,x_{0})$, and hence
 $\mathbb{E}\left[f(X^{\mathscr{G}}_{t_{j}})-f(\widetilde{X}_{t_{j}})\right]$ $=\mathbb{E}\left[v(t_{j},X^{\mathscr{G}}_{t_{j}})-v(0,x_{0})\right]$ (6.3)
 $=\mathbb{E}\left[\sum_{i=1}^{j}\left(v(t_{i},X^{\mathscr{G}}_{t_{i}})-v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right]=:\sum_{i=1}^{j}\mathbb{E}\left[e_{i}\right].$
-It suffices to estimate every $\mathbb{E}[e_{i}]$.
+It suffices to estimate each $\mathbb{E}[e_{i}]$.
 For simplicity, denote $\partial_{t}:=\partial/\partial t$ and $\partial_{x}:=\partial/\partial x$.
@@
 $\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})\right)\mathrm{d}s\bigg],$
-with $a_{t_{i-1}}\sim\bm{\pi}(\cdot|t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$, where in the second equality, we use the fact that the Itô integral is a martingale, given that $\sigma$ and $\partial_{x}v$ have polynomial growth and the moments of $X^{\mathscr{G}}_{s}$ are finite (by Lemma 3.1).
+with $a_{t_{i-1}}\sim\bm{\pi}(\cdot|t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$, where the second equality used the fact that the Itô integral is a martingale, given that $\sigma$ and $\partial_{x}v$ have polynomial growth and the moments of $X^{\mathscr{G}}_{s}$ are finite (by Lemma 3.1).
 Noting that $\xi_{i-1}$ is independent of $X^{\mathscr{G}}_{t_{i-1}}$, we have
@@
 $=\mathbb{E}\bigg[\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+(\tilde{b}^{\bm{\pi}})^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
-$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}})^{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}=0,$
-where the first equality follows form taking conditional expectation $\mathbb{E}[\cdot|X^{\mathscr{G}}_{t_{i-1}}]$ and the second equality follows from (7.1). Therefore,
+$\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left((\tilde{\sigma}^{\bm{\pi}})^{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg]{\color[rgb]{0,0,0}=0},$
+where the first equality follows from taking conditional expectation $\mathbb{E}[\cdot|X^{\mathscr{G}}_{t_{i-1}}]$ and the second equality follows from (6.2). Therefore,
 $\mathbb{E}\left[e_{i}\right]$ $=\mathbb{E}\bigg[\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})-\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg]$ (6.5)
-$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}v(s,X^{\mathscr{G}}_{s})-\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
-$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right)\mathrm{d}s\bigg{]}$
-$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
-$\quad+\mathbb{E}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
-For $h=\partial_{t}v,\partial_{x}v,\partial_{x}^{2}v,b(\cdot,\cdot,a_{t_{i-1}})\ \text{or}\ \frac{1}{2}\sigma\sigma^{\top}(\cdot,\cdot,a_{t_{i-1}})$, we again use Itô’s lemma to obtain
-$h$ $(s,X^{\mathscr{G}}_{s})-h(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (7.5)
-$=\int_{t_{i-1}}^{s}\left(\partial_{t}h(u,X^{\mathscr{G}}_{u})+b^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}h(u,X^{\mathscr{G}}_{u})\right)\mathrm{d}u$
-$\quad+\int_{t_{i-1}}^{s}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}^{2}h(u,X^{\mathscr{G}}_{u})\right)\mathrm{d}u+\int_{t_{i-1}}^{s}(\partial_{x}h(u,X^{\mathscr{G}}_{u}))^{\top}\sigma(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\mathrm{d}W_{u}.$
-By (H.3), Lemma 3.1 and due to $b,\sigma\sigma^{\top}\in C_{p}^{2,0}$, the integrands on the right hand side of the above equation have finite moments.
-Therefore, if we substitute (7.5) to the first and last two terms of (7.4), then every time integral term is bounded by $C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}$ for constant $C_{p}>0$ depends only on $b,\sigma,\bm{\pi},T,p$, and the Itô integral terms vanish.
-For the second and third terms of (7.4), let $h_{1}=b(\cdot,\cdot,a_{t_{i-1}})\ \text{or}\ \frac{1}{2}\sigma\sigma^{\top}(\cdot,\cdot,a_{t_{i-1}})$ and $h_{2}=\partial_{x}v$.
-We have
-$\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}h_{1}(s,X^{\mathscr{G}}_{s})\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right]$ (7.6)
-$=$ $\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right]$
-$+\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left(h_{1}(s,X^{\mathscr{G}}_{s})-h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\left(h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\right].$
-For the first term of (7.6), we again use the Itô’s Lemma to obtain the upper bound $C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}$.
-Note that by Hölder’s inequality and (7.5), the second term of (7.6) is bounded by
-$\sqrt{\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left|h_{1}(s,X^{\mathscr{G}}_{s})-h_{1}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right|^{2}\mathrm{d}s\right]\mathbb{E}\left[\int_{t_{i-1}}^{t_{i}}\left|h_{2}(s,X^{\mathscr{G}}_{s})-h_{2}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right|^{2}\mathrm{d}s\right]}\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2}.$
-Thus $\mathbb{E}\left[e_{i}\right]\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$ where $C_{p}$ is a constant depending only on $b,\sigma,\bm{\pi},T,p$.
-The desired result follows from $\sum_{i=1}^{j}(t_{i}-t_{i-1})^{2}\leq T|\mathscr{G}|$.
-∎
+$+$ $\mathbb{E}\bigg[\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}v(s,X^{\mathscr{G}}_{s})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg]$
+$+$ $\mathbb{E}\bigg[\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg].$
+Based on (H.3), $\partial_{t}v,b^{\top}\partial_{x}v,\frac{1}{2}\operatorname*{Tr}(\sigma\sigma^{\top}\partial_{x}^{2}v)\in C_{p}^{2,0}$. It follows from Lemma 6.1 that
+$\left|\mathbb{E}\left[e_{i}\right]\right|$ $\leq\int_{t_{i-1}}^{t_{i}}C\left(\|\partial_{t}v\|_{C_{p}^{2}}+\|b^{\top}\partial_{x}v\|_{C_{p}^{2,0}}+\frac{1}{2}\|\operatorname*{Tr}(\sigma\sigma^{\top}\partial_{x}^{2}v)\|_{C_{p}^{2,0}}\right)|\mathscr{G}|\mathrm{d}s$ (6.6)
+$\leq C_{p}\|v\|_{C_{p}^{4}}|\mathscr{G}|(t_{i}-t_{i-1})$
+for a constant $C_{p}$ depending only on $b,\sigma,p$. The desired result follows from (6.3), (6.6), and $\|v\|_{C_{p}^{4}}\leq C\|f\|_{C_{p}^{4}}$ due to (H.3).
 6.1.2. Proof of Theorem 4.2
-The following lemma is similar to [23, Lemma 14.1.6]. One can combine the method in Theorem 4.1 to prove it, so we omit the proof here.
+The following lemma is similar to [25, Lemma 14.1.6]. One can combine the method in Theorem 4.1 to prove it, so we omit the proof here.
 Lemma 6.2.
-Suppose (H.1) holds, $l\in(0,1)\cup(1,2)\cup(2,3)$ and $g\in\mathcal{H}_{T}^{(l)}$. Then there exist a constant $C$, such that for all $s\in[0,T]$,
+Suppose (H.1) holds, $l\in(0,1)\cup(1,2)\cup(2,3)$ and $g\in\mathcal{H}_{T}^{(l)}$. Then there is a constant $C$, such that for all $s\in[0,T]$,
 $\left|\mathbb{E}\left[g(s,X^{\mathscr{G}}_{s},a_{\delta(s)})-g(\delta(s),X^{\mathscr{G}}_{\delta(s)},a_{\delta(s)})\right]\right|\leq C\|g\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)},$ (6.7)
@@
 Based on Lemma 6.2, we prove Theorem 4.2.
-Proof.
+Proof of Theorem 4.2.
 It suffices to prove $|\mathbb{E}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}[f(\widetilde{X}_{t^{\prime}})]|\leq C\|f\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}$ for any $t^{\prime}\in[0,T]$.
 Under (H.
-4), by [24, Chapter 4, Theorem 5.1], for any $f\in\mathcal{H}^{(l+2)}$, the PDE (4.4) has a unique solution $v(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where $C$ is a constant depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}$.
-Note that $\mathcal{H}_{T}^{(l)}\subset\mathcal{H}_{s}^{(l)}$ for every $0\leq t^{\prime}\leq T$.
+4), by [27, Chapter 4, Theorem 5.1], for any $f\in\mathcal{H}^{(l+2)}$, the PDE (4.4) has a unique solution $v(\cdot,\cdot;t^{\prime})\in\mathcal{H}_{t^{\prime}}^{(l+2)}$ and $\|v(\cdot,\cdot;t^{\prime})\|_{t^{\prime}}^{(l+2)}\leq C\|f\|^{(l+2)}$, where $C$ is a constant depending only on $T,\tilde{b}^{\bm{\pi}},\tilde{\sigma}^{\bm{\pi}}$.
+Note that $\mathcal{H}_{T}^{(l)}\subset\mathcal{H}_{{\color[rgb]{0,0,0}t^{\prime}}}^{(l)}$ for every $0\leq t^{\prime}\leq T$.
 Without loss of generality, suppose that $t^{\prime}=T$.
@@
 Recall $\mathcal{F}^{\xi}:=\sigma\{\xi_{i}\mid{i\geq 0}\}$ as defined in Section 4.2.
-To facilitate the proof, we define $\mathcal{F}^{\xi}_{i}:=\sigma\{\xi_{j}\mid j\leq i-1\}$ for all $i\geq 0$, and define the filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with ~$\widetilde{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W}\vee\mathcal{F}^{\xi}$.
-Note that $\{W_{s}\}_{s\geq 0}$ is still a Brownian motion on $(\Omega,\mathcal{F},{\mathbb{P}})$ with filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ and we have $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}\big{[}\cdot\big{|}\widetilde{\mathcal{F}}_{0}\big{]}$.
+To facilitate the proof, we define $\mathcal{F}^{\xi}_{i}:=\sigma\{\xi_{j}\mid j\leq i-1\}$ for all $i\geq 0$, and define the filtration $\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with $\widetilde{\mathcal{F}}_{t}:=\mathcal{F}_{t}^{W}\vee\mathcal{F}^{\xi}$.
+Note that $\{W_{s}\}_{s\geq 0}$ is still a Brownian motion on $(\Omega,\mathcal{F},{\mathbb{P}})$ with filtration $\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ and we have $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}\big[\cdot\big|\widetilde{\mathcal{F}}_{0}\big]$.
 The following lemma shows that $X^{\mathscr{G}}$ has finite moments conditional on $\mathcal{F}^{\xi}$. The proof is given in Appendix A.
@@
 $=\mathbb{E}^{W}\left[\sum_{i=1}^{j}\left(v(t_{i},X^{\mathscr{G}}_{t_{i}})-v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right]=:\sum_{i=1}^{j}\mathbb{E}^{W}\left[e_{i}\right].$
-As in the proof of Theorem 4.1, noting that the Itô integral in (7.5) is mean zero under $\mathbb{P}^{W}$, we can decompose $\mathbb{E}^{W}\left[e_{i}\right]$ into two terms and obtain
+As in the proof of Theorem 4.1, we can decompose $\mathbb{E}^{W}\left[e_{i}\right]$ into two terms and obtain
 $\mathbb{E}^{W}\left[f(X^{\mathscr{G}}_{T})\right]-\mathbb{E}^{W}\left[f(\widetilde{X}_{T})\right]=\sum_{i=1}^{j}A_{i}+\sum_{i=1}^{j}B_{i},$ (6.12)
 where $A_{i}$ is defined by
-$A_{i}:=$ $\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
+$A_{i}$= $\mathbb{E}^{W}\bigg[\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s$
 $\quad+\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg].$
@@
 $B_{i}$ $\coloneqq\mathbb{E}^{W}\bigg[\int_{t_{i-1}}^{t_{i}}\left(\partial_{t}v(s,X^{\mathscr{G}}_{s})-\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg]$ (6.13)
-$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}v(s,X^{\mathscr{G}}_{s})-\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}$
-$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\left(\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\right)\mathrm{d}s\bigg{]}$
-$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\left(b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg{]}$
-$\quad+\mathbb{E}^{W}\bigg{[}\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\right)\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg{]}.$
-We again use the Itô’s lemma to analysis every difference term of (7.13).
-Note that every term in $B_{j}$ can be written as double integrals with respect to time as in the proof of Theorem 4.1, and by conditions (H.5) and (H.3), the integand is bounded by $C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{s}|^{m_{p}})$ for some $m_{p}$ and $C_{p}$ that depend on $p$.
-Applying Lemma 7.2, we obtain
-$|B_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$
-where $C_{p}$ is a constant depending on $b,\sigma,\bm{\pi},T,p$. As a result,
+$+$ $\mathbb{E}^{W}\bigg[\int_{t_{i-1}}^{t_{i}}b^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}v(s,X^{\mathscr{G}}_{s})-b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\mathrm{d}s\bigg]$
+$+$ $\mathbb{E}^{W}\bigg[\int_{t_{i-1}}^{t_{i}}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\partial_{x}^{2}v(s,X^{\mathscr{G}}_{s})-\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\mathrm{d}s\bigg].$
+As in Lemma 6.1, apply Itô’s lemma to every difference term on the right hand side of (6.13) and note that the resulting Itô’s integrals have mean zero under $\mathbb{P}^{W}$.
+Thus by (H.3) and (H.5), the integrands of (6.13) are bounded by $C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{s}|^{m_{p}})(t_{i}-t_{i-1})$ for some $m_{p}$ and $C_{p}$ that depend on $p$.
+This along with Lemma 6.3 implies $|B_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})^{2},$ where $C_{p}$ is a constant depending on $b,\sigma,\bm{\pi},T,p$.
+As a result,
 $\sum_{i=1}^{j}|B_{i}|\leq\sum_{i=1}^{j}C_{p}\|v\|_{C_{p}^{4}}|\mathscr{G}|(t_{i}-t_{i-1})\leq TC_{p}\|v\|_{C_{p}^{4}}|\mathscr{G}|.$ (6.14)
 As for $A_{i}$, note that $\{(A_{i},\mathcal{F}^{\xi}_{i})\}_{i}$ is a martingale difference sequence. In fact, denote
-$S_{i}:=\bigg{(}\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ $+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (7.15)
+$S_{i}=\bigg(\partial_{t}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ $+b^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})$ (6.15)
 $+\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i-1}})\partial_{x}^{2}v(t_{i-1},X^{\mathscr{G}}_{t_{i-1}})\right)\bigg)(t_{i}-t_{i-1}).$
@@
 By conditions (H.3) and (H.5), there is a constant $C_{p}$ and an integer $m_{p}$ such that $|S_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(1+|X^{\mathscr{G}}_{t_{i-1}}|^{m_{p}})$.
-Again using Lemma 7.2, we obtain
-$|A_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1}).$ (7.17)
-By [35, Corollary 2.20], we have
+Again by Lemma 6.3, $|A_{i}|\leq C_{p}\|v\|_{C_{p}^{4}}(t_{i}-t_{i-1})$.
+By the Azuma–Hoeffding inequality [41, Corollary 2.20], for all $t\geq 0$,
 $\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq t\bigg)\leq 2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right).$ (6.17)
-Let $\varrho=2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right)$, we obtain
-$\mathbb{P}\bigg{(}\bigg{|}\sum_{j=1}^{i}A_{i}\bigg{|}\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}\leq\varrho.$ (7.19)
+Hence given $\rho\in(0,1)$, choosing $t\geq 0$ such that $\varrho=2\exp\left(-\frac{t^{2}}{4T(C_{p}\|v\|_{C_{p}^{4}})^{2}|\mathscr{G}|}\right)$ yields
+$\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg)\leq\varrho.$ (6.18)
 It follows from (6.14) that
-$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t^{\prime}})]\right|>(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right)$
-$\quad\leq\mathbb{P}\bigg{(}\bigg{|}\sum_{j=1}^{i}A_{i}\bigg{|}\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\right)$
+$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t_{j}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t_{j}})]\right|>(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\big(1+\sqrt{\ln(2/\varrho)}\big)\right)$
+$\quad\leq\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq 2\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg)+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C_{p}\|v\|_{C_{p}^{4}}\sqrt{|\mathscr{G}|}\right)$
 $\quad\leq\varrho+0=\varrho.$
 The desired result follows from $\|v\|_{C_{p}^{4}}\leq C\|f\|_{C_{p}^{4}}$.
-For the proof of (4.16), we can start from (7.12). Note that $\{A_{i}\}_{i}$ is still a martingale difference sequence. By (H.4), we have
-$|A_{i}|\leq C\|v\|_{t_{j}}^{(l+2)}(t_{i}-t_{i-1})$ (7.20)
-It suffices to analyze (7.13). Applying Lemmas 7.2 and 7.1, one can use a similar method in the proof of Theorem 4.2 to obtain
-$|B_{i}|\leq C\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1}).$ (7.21)
-As a result,
+To prove (4.16), we start from (6.12).
+For (6.13), note that under assumption (H.4), Lemma 6.3 remains valid, and the expectation in Lemma 6.2 can be replaced by $\mathbb{E}^{W}$.
+Based on these observations, one can use a similar method in the proof of Theorem 4.2 to obtain
+$|B_{i}|\leq C\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}(t_{i}-t_{i-1}),$
+which implies that
 $\sum_{i=1}^{j}|B_{i}|\leq CT\|v\|^{(l+2)}_{t_{j}}|\mathscr{G}|^{\mathcal{X}(l)}.$ (6.19)
-By [35, Corollary 2.20], we have
-$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq t\bigg{)}\leq 2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right).$ (7.23)
-Let $\varrho=2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right)$, we obtain
-$\mathbb{P}\bigg{(}\bigg{|}\sum_{i=1}^{j}A_{i}\bigg{|}\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg{)}\leq\varrho.$ (7.24)
+Note that $\{A_{i}\}_{i}$ is still a martingale difference sequence. By (H.4), we have
+$|A_{i}|\leq C\|v\|_{t_{j}}^{(l+2)}(t_{i}-t_{i-1}).$
+Again by the Azuma–Hoeffding inequality [41, Corollary 2.20], for all $t\geq 0$,
+$\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq t\bigg)\leq 2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right).$
+Hence given $\varrho\in(0,1)$, choosing $t\geq 0$ such that $\varrho=2\exp\left({-\frac{t^{2}}{4T(C\|v\|_{t_{j}}^{(l+2)})^{2}|\mathscr{G}|}}\right)$ yields
+$\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg)\leq\varrho.$
 It follows from (6.19) that
-$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t^{\prime}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t^{\prime}})]\right|>(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big{(}1+\sqrt{\ln(2/\varrho)}\big{)}\right)$
+$\mathbb{P}$ $\left(\left|\mathbb{E}^{W}[f(X^{\mathscr{G}}_{t_{j}})]-\mathbb{E}^{W}[f(\widetilde{X}_{t_{j}})]\right|>(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)\land\frac{1}{2}}\big(1+\sqrt{\ln(2/\varrho)}\big)\right)$
 $\quad\leq\mathbb{P}\bigg(\bigg|\sum_{i=1}^{j}A_{i}\bigg|\geq 2\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}\sqrt{|\mathscr{G}|}\sqrt{\ln(2/\varrho)}\bigg)+\mathbb{P}\left(\sum_{i=1}^{j}|B_{i}|\geq(2+\sqrt{T})\sqrt{T}C\|v\|_{t_{j}}^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}\right)$
@@
 We introduce the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$.
-Let $\psi$ be a convex, nondecreasing function with $\psi(0)=0$.
-For a random variable $X(\omega):\Omega^{\xi}\rightarrow\mathbb{R}^{m},m\in\mathbb{N}_{+}$, the Orlicz norm is defined as
-$\|X\|_{\psi}=\inf\left\{C>0:\mathbb{E}\left[\psi\left(\frac{|X|}{C}\right)\right]\leq 1\right\},$ (7.25)
-where ||$\cdot$|| ⋅ | is the Euclid norm on $\mathbb{R}^{m}$, and the corresponding Orlicz space is $L_{\psi}$.
-Note that the $L_{p}$-norm $\|\cdot\|_{p}$ is the corresponding Orlicz norm with the function $\psi(x)=x^{p}$.
-Denote $\psi_{\alpha}=e^{x^{\alpha}}-1,\alpha>0$ (linear near the origin when $10<\alpha<$ 1 in order for $\psi_{\alpha}$ to be convex).
-We say $X$ is sub-exponential if and only if $X\in L_{\psi_{1}}$.
-It follows from $|x|^{p}/p!\leq e^{|x|}-1$ that
-$\|X\|_{p}\leq\sqrt[p]{p!}\|X\|_{\psi_{1}}.$ (7.26)
-Combining the above inequality and [34, Lemma 2.2.2], we derive the following lemma.
+Define $\psi_{1}(x)\coloneqq e^{x}-1$ for all $x>0$, and for a random variable $X:\Omega^{\xi}\rightarrow\mathbb{R}^{m},m\in\mathbb{N}_{+}$, define the Orlicz norm
+$\|X\|_{\psi_{1}}\coloneqq\inf\left\{C>0:\mathbb{E}\left[\psi_{1}\left(\frac{|X|}{C}\right)\right]\leq 1\right\},$ (6.20)
+where $|\cdot|$ is the Euclidean norm on $\mathbb{R}^{m}$. By [40, Proposition 2.7.1], $X$ is sub-exponential if and only if $\|X\|_{\psi_{1}}<\infty$.
+The following lemma follows directly from [39, Lemma 2.2.2] and the fact that $\|X\|_{p}\leq\sqrt[p]{p!}\|X\|_{\psi_{1}}$, due to the inequality that $|x|^{p}/p!\leq e^{|x|}-1$ for all $x\in{\mathbb{R}}^{m}$.
 Lemma 6.4.
-For $p\geq 1$, if $\{X_{i}\}_{i=1}^{n}$ are sub-exponential random variables, then there is a constant $C_{p}$, such that
-$\mathbb{E}\left[\max_{1\leq i\leq n}|X_{i}|^{p}\right]\leq C_{p}\log^{p}(1+n)\max_{i}\|X_{i}\|_{\psi_{1}}^{p}.$ (7.27)
+There is a constant $C\geq 0$ such that for all sub-exponential random variables $\{X_{i}\}_{i=1}^{n}$,
+$\big\|\max_{1\leq i\leq n}|X_{i}|\big\|_{\psi_{1}}\leq C\log(1+n)\max_{1\leq i\leq n}\|X_{i}\|_{\psi_{1}}.$ (6.21)
+Thus, for $p\geq 1$, there is a constant $C_{p}\geq 0$ such that
+$\mathbb{E}\left[\max_{1\leq i\leq n}|X_{i}|^{p}\right]\leq C_{p}\log^{p}(1+n)\max_{1\leq i\leq n}\|X_{i}\|_{\psi_{1}}^{p}.$ (6.22)
 Note that under condition (H.6), $\tilde{b}^{\bm{\pi}}$ and $\tilde{\sigma}^{\bm{\pi}}$ is globally Lipschitz with respect to the state $x$.
 By classical SDE results, e.g.
-[13, Theorem 7.9], we have the following lemma.
+[14, Theorem 7.9], we have the following lemma.
 Lemma 6.5.
@@
 Proof of Theorem 4.4.
-By the Lipschitz condition of $\phi$, $d_{A}(\phi(t_{0},x_{0},\xi),a_{0})$ is sub-exponential if and only if $d_{A}(\phi(0,0,\xi),a_{0})$ is sub-exponential by the triangular inequality and the condition (1) of (H.6).
-we take $(t_{0},x_{0})=(0,0)$ for simplicity.
-Recall the notation $i(t):=\max\{i:t_{i}\leq t\}$.
-Note that
+By (H.6), $d_{A}(\phi(0,0,\xi),a_{0})$ is sub-exponential. Recall the notation $i(t):=\max\{i:t_{i}\leq t\}$ and for all $t\in[0,T]$,
 $X^{\mathscr{G}}_{t}$ $=x_{0}+\int_{0}^{t}b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))\mathrm{d}s+\int_{0}^{t}\sigma(s,X^{\mathscr{G}}_{s})\mathrm{d}W_{s},$ (6.25)
 $\widetilde{X}_{t}$ $=x_{0}+\int_{0}^{t}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s+\int_{0}^{t}\sigma(s,\widetilde{X}_{s})\mathrm{d}W_{s}.$
-Denote $\Delta X_{t}:=X^{\mathscr{G}}_{t}-\widetilde{X}_{t}$. We have
-$\sup_{0\leq u\leq t}|\Delta X_{u}|$ $=\sup_{0\leq u\leq t}\bigg{|}\int_{0}^{u}\left(b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))-b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\right)\mathrm{d}s$ (7.31)
+For all $t\in[0,T]$, $\Delta X_{t}:=X^{\mathscr{G}}_{t}-\widetilde{X}_{t}$ satisfies
+$\sup_{0\leq u\leq t}|\Delta X_{u}|$ (6.26)
+$=\sup_{0\leq u\leq t}\bigg|\int_{0}^{u}\left(b(s,X^{\mathscr{G}}_{s},\phi(\delta(s),X^{\mathscr{G}}_{\delta(s)},\xi_{i(s)}))-b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\right)\mathrm{d}s$
 $\quad+\int_{0}^{u}\left(b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\right)\mathrm{d}s+\int_{0}^{u}\left(\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right)\mathrm{d}W_{s}\bigg|.$
-It follows from the Lipschitz conditions and the Burkholder-Davis-Gundy inequality that for $p\geq 2$,
+By the Lipschitz continuity of $\sigma$ and the Burkholder-Davis-Gundy inequality, for all $p\geq 2$,
 $\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}\left(\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right)\mathrm{d}W_{s}\right|^{p}\right]$ $\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}\left|\sigma(s,X^{\mathscr{G}}_{s})-\sigma(s,\widetilde{X}_{s})\right|^{2}\mathrm{d}s\right)^{\frac{p}{2}}\right]$
-$\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}|\Delta X_{u}|^{2}\mathrm{d}u\right)^{\frac{p}{2}}\right]\leq C_{p}\mathbb{E}\left[\int_{0}^{t}|\Delta X_{u}|^{p}\mathrm{d}u\right].$
-Thus, we have
+$\leq C_{p}\mathbb{E}\left[\left(\int_{0}^{t}|\Delta X_{u}|^{2}\mathrm{d}u\right)^{\frac{p}{2}}\right]\leq C_{p}\mathbb{E}\left[\int_{0}^{t}|\Delta X_{u}|^{p}\mathrm{d}u\right],$
+where $C_{p}>0$ is a generic constant that depends on $p$ and the coefficients, but is independent of the grid $\mathscr{G}$.
+This along with (6.26) and the regularity of $b$ and $\phi$ in (H.6) implies that
 $\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]\leq$ $C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\left|\Delta X_{s}\right|^{p}+\left|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}\right|^{p}\right]\mathrm{d}s$ (6.27)
-$+C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))\mathrm{d}s-\int_{0}^{u}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s\right|^{p}\right].$
-We focus on the third term of the above equation. Denote $U_{s}:=b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})$. Under (H.6), we have
+$+C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}\left(b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\right)\mathrm{d}s\right|^{p}\right].$
+We now bound the last term in (6.27). For all $s\in[0,T]$, write $U_{s}:=b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})$. By (H.6), we have
 $|U_{s}|=\left|b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\right|\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|+d_{A}(\phi(0,0,\xi_{i(s)}),a_{0})\right).$ (6.28)
-Denote $\mathbb{E}_{\xi}\coloneqq\mathbb{E}\left[\cdot|\mathcal{F}^{W}\right]$ and note that $\xi$ is independent of $\mathcal{F}^{W}$.
-$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\leq C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]+C_{p}\mathbb{E}\left[\mathbb{E}_{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right].$ (7.34)
-It follows from (7.33) and Lemma 7.3 that
-$\mathbb{E}\left[\mathbb{E}_{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right]$ $\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\mathbb{E}\left[\max_{1\leq j\leq n}\left(d_{A}(\phi(0,0,\xi_{j}),a_{0})\right)^{p}\right]\right)|\mathscr{G}|^{p}\right]$ (7.35)
+Denote $\mathbb{E}^{\xi}[\cdot]\coloneqq\mathbb{E}\left[\cdot|\mathcal{F}^{W}\right]$ and note that $\xi$ is independent of $\mathcal{F}^{W}$.
+$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\leq C_{p}\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]+C_{p}\mathbb{E}\left[\mathbb{E}^{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right].$ (6.29)
+By (6.28) and Lemma 6.4, the second term on the right hand side of (6.29) satisfies
+$\mathbb{E}\left[\mathbb{E}^{\xi}\left[\sup_{0\leq u\leq t}\left|\int_{\delta(u)}^{u}U_{s}\mathrm{d}s\right|^{p}\right]\right]$ $\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\mathbb{E}\left[\max_{1\leq j\leq n}\left(d_{A}(\phi(0,0,\xi_{j}),a_{0})\right)^{p}\right]\right)|\mathscr{G}|^{p}\right]$ (6.30)
 $\leq C_{p}\mathbb{E}\left[\left(1+\sup_{0\leq t\leq T}\left|\widetilde{X}_{t}\right|^{p}+\log^{p}(1+n)\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}^{p}\right)|\mathscr{G}|^{p}\right]$
 $\leq C_{p}|\mathscr{G}|^{\frac{p}{2}},$
-where in the last inequality, we used the fact that $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}$.
-As for the first term on the right hand side of (7.34), denote $M_{j}:=\int_{t_{j}}^{t_{j+1}}U_{s}\mathrm{d}s$ and $\mathcal{F}_{i}:=\sigma(\xi_{j}:0\leq j\leq i-1)\lor\mathcal{F}^{W}$.
-Note that $\left\{(\sum_{j=0}^{i-1}M_{j},\mathcal{F}_{i})\right\}_{i=1}^{n}$ is a martingale.
-It follows from Doob’s inequality that
-$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]=\mathbb{E}\left[\max_{1\leq i\leq n}\Big{|}\sum_{j=0}^{i-1}M_{j}\Big{|}^{p}\right]\leq C_{p}\mathbb{E}\left[\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}^{p}\right],$ (7.36)
-where $C_{p}$ is a constant depend on $p$.
-By (H.6) and relation (7.33), $M_{j}$ is sub-exponential conditional on $\mathcal{F}^{W}$ and with a little abuse of notations,
-$\|M_{j}\|_{\psi_{1}}\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\|1\|_{\psi_{1}}+\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}\right)(t_{j+1}-t_{j}),$ (7.37)
-where $\|\cdot\|_{\psi_{1}}$ is the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$. It follows from [25, Theorem 6.21] that
-$\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\leq C\left(\mathbb{E}_{\xi}\left[\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}\right]+\|\max_{j}|M_{j}|\|_{\psi_{1}}\right).$ (7.38)
-For the first term on the right hand side of the above equation, using the fact that $\mathbb{E}_{\xi}\left[M_{i}^{\top}M_{j}\right]=0,i\neq j$, it follows from relation (7.33) and $\psi(0,0,\xi)$ is square integrable that
-$\mathbb{E}_{\xi}\bigg{[}\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}\bigg{]}\leq\sqrt{\mathbb{E}_{\xi}\bigg{[}\Big{|}\sum_{j=0}^{n-1}M_{j}\Big{|}^{2}\bigg{]}}=\sqrt{\sum_{j=0}^{n-1}\mathbb{E}_{\xi}\left[|M_{j}|^{2}\right]}\leq C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}.$ (7.39)
-As for the second term, it follows from [34, Lemma 2.2.2] and (7.37) that
-$\|\max_{j}|M_{j}|\|_{\psi_{1}}\leq C\log(1+n)\max_{j}\|M_{j}\|_{\psi_{1}}\leq C\log(1+n)\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|.$ (7.40)
-In conclusion, we have
+where the last inequality used the condition $|\mathscr{G}|^{1/2}\log(1+n)\leq C_{G}$.
+To bound the first term on the right hand side of (6.29), for each $i=0,\ldots n-1$, let $M_{i}:=\int_{t_{i}}^{t_{i+1}}U_{s}\mathrm{d}s$ and $\mathcal{F}_{i}:=\sigma(\xi_{j}:0\leq j\leq i-1)\lor\mathcal{F}^{W}$.
+Since $\left\{(\sum_{j=0}^{i-1}M_{j},\mathcal{F}_{i})\right\}_{i=1}^{n}$ is a martingale, by Doob’s inequality,
+$\mathbb{E}\left[\sup_{0\leq u\leq t}\left|\int_{0}^{\delta(u)}U_{s}\mathrm{d}s\right|^{p}\right]=\mathbb{E}\left[\max_{1\leq i\leq n}\Big|\sum_{j=0}^{i-1}M_{j}\Big|^{p}\right]\leq C_{p}\mathbb{E}\left[\left|\sum_{j=0}^{n-1}M_{j}\right|^{p}\right].$ (6.31)
+In the sequel, we estimate (6.31) by quantifying the tail behavior of $\left|\sum_{j=0}^{n-1}M_{j}\right|$ conditioning on the realization of the Brownian motion $W$.
+Let $\|\cdot\|_{\psi_{1}}$ be the Orlicz norm on $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$, conditioning on $\mathcal{F}^{W}$. Since for all $j=0,\ldots,i-1$,
+$\mathbb{E}^{\xi}[M_{j}]=\int_{t_{j}}^{t_{j+1}}\mathbb{E}^{\xi}[U_{s}]\mathrm{d}s=\int_{t_{j}}^{t_{j+1}}\mathbb{E}^{\xi}[b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{i(s)}))-\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})]\mathrm{d}s=0,$
+and hence by [28, Theorem 6.21],
+$\Big\|\sum_{j=0}^{n-1}M_{j}\Big\|_{\psi_{1}}\leq C\left(\mathbb{E}^{\xi}\left[\Big|\sum_{j=0}^{n-1}M_{j}\Big|\right]+\|\max_{j}|M_{j}|\|_{\psi_{1}}\right).$
+Note that conditioning on $\mathcal{F}^{W}$, $\xi_{i}$ is the only randomness in $M_{i}$, and hence by the independent of $\xi$ and $\xi_{j}$ for $i\not=j$, $\mathbb{E}^{\xi}\left[M_{i}^{\top}M_{j}\right]=\mathbb{E}^{\xi}[M_{i}]^{\top}\mathbb{E}^{\xi}[M_{j}]=0$ for all $i\neq j$.
+Then using (6.28) and the square integrability of $\phi(0,0,\xi)$,
+$\mathbb{E}^{\xi}\bigg[\Big|\sum_{j=0}^{n-1}M_{j}\Big|\bigg]\leq\sqrt{\mathbb{E}^{\xi}\bigg[\Big|\sum_{j=0}^{n-1}M_{j}\Big|^{2}\bigg]}=\sqrt{\sum_{j=0}^{n-1}\mathbb{E}^{\xi}\left[|M_{j}|^{2}\right]}\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|^{\frac{1}{2}}.$
+By (6.28) and (H.6),
+$\|M_{j}\|_{\psi_{1}}\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\|1\|_{\psi_{1}}+\|d_{A}(\phi(0,0,\xi),a_{0})\|_{\psi_{1}}\right)|\mathscr{G}|\leq C\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|,$ (6.32)
+which along with Lemma 6.4 implies that $\left\|\max_{j}|M_{j}|\right\|_{\psi_{1}}\leq C\log(1+n)\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|\right)|\mathscr{G}|$, and
 $\Big\|\sum_{j=0}^{n-1}M_{j}\Big\|_{\psi_{1}}\leq C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}.$ (6.33)
-Conditional on $\mathcal{F}^{W}$, we have
-$\mathbb{P}_{\xi}\left(\bigg{|}\sum_{j=0}^{n-1}M_{j}\bigg{|}^{p}\geq t\right)$ $=\mathbb{P}_{\xi}\left(\psi_{1}\bigg{(}\bigg{|}\sum_{j=0}^{n-1}M_{j}\bigg{|}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\leq\psi_{1}\bigg{(}t^{\frac{1}{p}}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\right)$ (7.42)
-$\leq 1/\psi_{1}\bigg{(}t^{\frac{1}{p}}/\Big{\|}\sum_{j=0}^{n-1}M_{j}\Big{\|}_{\psi_{1}}\bigg{)}\leq 2\exp\left\{-\frac{t^{\frac{1}{p}}}{C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}}\right\}.$
-Integrating the above equation with respect to $t$ from 00 to +$\infty$+ ∞, we have
-$\mathbb{E}\left[\left|\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}b(s,\widetilde{X}_{s},\phi(s,\widetilde{X}_{s},\xi_{j}))\mathrm{d}s-\int_{t_{i}}^{t_{i+1}}\tilde{b}^{\bm{\pi}}(s,\widetilde{X}_{s})\mathrm{d}s\right|^{p}\bigg{|}\mathcal{F}^{W}\right]\leq C_{p}(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p})|\mathscr{G}|^{\frac{p}{2}}.$ (7.43)
-Combining (7.32), (7.34), (7.35) and (7.43), using Lemma 7.4 to bound the time difference of $X^{\mathscr{G}}$ and $\widetilde{X}$ and the moments of $\sup_{t}|\widetilde{X}_{t}|$, we have
-$\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]$ $\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}\left|\Delta X_{u}\right|^{p}\right]+\mathbb{E}\left[|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}|^{p}\right]\mathrm{d}s$ (7.44)
-$\quad\quad+C_{p}|\mathscr{G}|^{\frac{p}{2}}\int_{0}^{t}\mathbb{E}[(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p})]\mathrm{d}s$
-$\leq C_{p}\left|\mathscr{G}\right|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}|\Delta X_{u}|\right]\mathrm{d}s.$
-It follows from the Grönwall’s inequality that
-$\mathbb{E}\left[\sup_{0\leq s\leq T}\left|\Delta X_{s}\right|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}.$ (7.45)
-Now we prove the second part of this theorem. If $|\mathscr{G}|=\frac{T}{n}$, then for $p\geq 2$, $t\geq 0$,
+For all $t>0$, by the conditional Markov inequality and $\mathbb{E}^{\xi}[\psi_{1}(|\sum_{j=1}^{n-1}M_{j}|/\|\sum_{j=1}^{n-1}M_{j}\|_{\psi_{1}})]\leq 1$,
+$\mathbb{P}^{\xi}\left(\bigg|\sum_{j=0}^{n-1}M_{j}\bigg|^{p}\geq t\right)$ $=\mathbb{P}^{\xi}\left(\psi_{1}\left(\frac{\left|\sum_{j=0}^{n-1}M_{j}\right|}{\left\|\sum_{j=0}^{n-1}M_{j}\right\|_{\psi_{1}}}\right)+1\leq\psi_{1}\left(\frac{t^{\frac{1}{p}}}{\left\|\sum_{j=0}^{n-1}M_{j}\right\|_{\psi_{1}}}\right)+1\right)$
+$\leq 2\left(\psi_{1}\left(\frac{t^{\frac{1}{p}}}{\left\|\sum_{j=0}^{n-1}M_{j}\right\|_{\psi_{1}}}\right)+1\right)^{-1}\leq 2\exp\left\{-\frac{t^{\frac{1}{p}}}{C(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|)|\mathscr{G}|^{\frac{1}{2}}}\right\},$
+where the last inequality used that $(\psi_{1}(x)+1)^{-1}=\exp(-x)$ for all $x>0$, and (6.33). Thus
+$\mathbb{E}^{\xi}\left[\bigg|{\sum_{j=0}^{n-1}M_{j}}\bigg|^{p}\right]=\int_{0}^{\infty}\mathbb{P}^{\xi}\left(\bigg|{\sum_{j=0}^{n-1}M_{j}}\bigg|^{p}\geq t\right)\mathrm{d}t\leq C_{p}\left(1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p}\right)|\mathscr{G}|^{\frac{p}{2}}.$ (6.34)
+Combining (6.27), (6.29), (6.30), (6.31) and (6.34), using Lemma 6.5 to bound the time difference of $X^{\mathscr{G}}$ and $\widetilde{X}$ and the moments of $\sup_{t}|\widetilde{X}_{t}|$, we have
+$\mathbb{E}\left[\sup_{0\leq s\leq t}\left|\Delta X_{s}\right|^{p}\right]$ $\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\left(\mathbb{E}\left[\sup_{0\leq u\leq s}\left|\Delta X_{u}\right|^{p}\right]+\mathbb{E}\left[|X^{\mathscr{G}}_{s}-X^{\mathscr{G}}_{\delta(s)}|^{p}\right]\right)\mathrm{d}s$ (6.35)
+$\quad+C_{p}|\mathscr{G}|^{\frac{p}{2}}\int_{0}^{t}\mathbb{E}\left[1+\sup_{0\leq t\leq T}|\widetilde{X}_{t}|^{p}\right]\mathrm{d}s$
+$\leq C_{p}\left|\mathscr{G}\right|^{\frac{p}{2}}+C_{p}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq u\leq s}|\Delta X_{u}|^{p}\right]\mathrm{d}s,$
+which yields the desired estimate $\mathbb{E}\left[\sup_{0\leq s\leq T}\left|\Delta X_{s}\right|^{p}\right]\leq C_{p}|\mathscr{G}|^{\frac{p}{2}}$ due to Grönwall’s inequality.
+It remains to prove the almost sure convergence with $|\mathscr{G}|=\frac{T}{n}$. For $p\geq 2$ and $t\geq 0$,
 $\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>t\right)=\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|^{p}>t^{p}\right)\leq t^{-p}\mathbb{E}\left[\sup_{0\leq t\leq T}|\Delta X_{t}|^{p}\right]\leq C_{p}n^{-\frac{p}{2}}t^{-p}.$ (6.36)
-Let $t=n^{-\frac{1}{2}+\varepsilon}$,
-$\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}n^{-\varepsilon p}.$ (7.47)
+Given $\varepsilon\in(0,1)$, setting $t=n^{-\frac{1}{2}+\varepsilon}$ in (6.36) yields for all $p\geq 2$,
+$\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}n^{-\varepsilon p}.$
 By choosing $p=\frac{1}{\varepsilon}+2$, we have
-$\sum_{n=1}^{\infty}\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}\sum_{n=1}^{\infty}n^{-1-2\varepsilon}<+\infty.$ (7.48)
+$\sum_{n=1}^{\infty}\mathbb{P}\left(\sup_{0\leq t\leq T}|\Delta X_{t}|>n^{-\frac{1}{2}+\varepsilon}\right)\leq C_{p}\sum_{n=1}^{\infty}n^{-1-2\varepsilon}<+\infty.$
 The desired result follows from the Borel-Cantelli lemma. ∎
+6.2. Proof of Corollary 4.1
+Proof of Corollary 4.1.
+Recall that $\mathbb{P}^{W}$ is the conditional probability measure given $\mathcal{F}^{\xi}$.
+Under the measure $\mathbb{P}^{W}$, $\{f(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi))\}_{k=1}^{m}$ are independent random variables taking values in $\|f\|_{\infty},\|f\|_{\infty}]$, where $\|f\|_{\infty}$ is the sup-norm of $f$.
+By Hoeffding’s inequality, for all $t\geq 0$ and $m,n\in{\mathbb{N}}$,
+$\mathbb{P}^{W}\left(\left|\frac{1}{m}\sum_{k=1}^{m}\left(f\left(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)\right)-\mathbb{E}^{W}\left[f\left(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)\right)\right]\right)\right|\geq t\right)\leq 2\exp\left(-\frac{2t^{2}m}{4\|f\|_{\infty}^{2}}\right),$ (6.37)
+For a given $\varrho\in(0,1)$, choosing $t$ such that $\varrho=2\exp(-t^{2}m/(2\|f\|_{\infty}^{2}))$ yields
+$\mathbb{P}^{W}\left(\left|\frac{1}{m}\sum_{i=1}^{m}f\left(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)\right)-\mathbb{E}^{W}\left[f\left(X^{\mathscr{G}_{n}}_{T}(W_{k},\xi)\right)\right]\right|\geq\sqrt{2}\|f\|_{\infty}\frac{1}{\sqrt{m}}\ln(2/\varrho)\right)\leq\varrho.$ (6.38)
+This along with Theorem 4.3 shows that there exists a constant $C\geq 0$ such that for all $n,m\in{\mathbb{N}}$ and $\varrho\in(0,1)$,
+$\mathbb{P}\left(\bigg|\hat{I}_{m,n}-\mathbb{E}[f(\widetilde{X}_{T})]\bigg|\geq C\left(1+\sqrt{\ln(2/\varrho)}\right)(n^{-1/2}+m^{-1/2})\right)\leq\varrho.$ (6.39)
+The desired result follows directly by choosing $n$ and $m$ as in the statement. ∎
 6.3. Proofs of Propositions 5.1, 5.3 and 5.5
@@
 For the first part (1), note that
-$J^{\mathscr{G}}-\tilde{J}=\left(\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right)+\sum_{i=0}^{n-1}\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}(r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s}))\mathrm{d}s\right]$ (7.49)
+$J^{\mathscr{G}}-\tilde{J}=\left(\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right)+\sum_{i=0}^{n-1}\mathbb{E}\left[\int_{t_{i}}^{t_{i+1}}(r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s}))\mathrm{d}s\right].$ (6.40)
 By Theorem 4.1, there is a constant $C_{p}$, such that
 $\left|\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right|\leq C_{p}\|h\|_{C_{p}^{4}}|\mathscr{G}|.$ (6.41)
-As for the second term on the right hand side of (7.49), note that if $t_{i}\leq s<t_{i+1}$, then
+For the second term on the right hand side of (6.40), note that if $t_{i}\leq s<t_{i+1}$, then
 $\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]$ $=\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]+\mathbb{E}\left[r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]$ (6.42)
@@
 Combining these inequalities, we prove the desired result.
-For the second part (2), we start at (7.49). By Theorem 4.2, there is a constant $C$, such that
+For the second part (2), we start at (6.40). By Theorem 4.2, there is a constant $C\geq 0$ such that
 $\left|\mathbb{E}\left[h(X^{\mathscr{G}}_{T})\right]-\mathbb{E}\left[h(\widetilde{X}_{T})\right]\right|\leq C\|h\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (6.46)
-As for the second term on the right hand side of (7.49), we still have (7.51). It follows from Lemma 7.1 that
+For the second term on the right hand side of (6.40), we still have (6.42). By Lemma 6.2,
 $|\mathbb{E}\left[r(s,X^{\mathscr{G}}_{s},a_{t_{i}})-r(t_{i},X^{\mathscr{G}}_{t_{i}},a_{t_{i}})\right]|\leq C\|r\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (6.47)
@@
 $\left|\mathbb{E}\left[\tilde{r}(t_{i},X^{\mathscr{G}}_{t_{i}})-\tilde{r}(t_{i},\widetilde{X}_{t_{i}})\right]\right|\leq C\|\tilde{r}(t_{i},\cdot)\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}\leq C\sup_{0\leq t\leq T}\|\tilde{r}(t,\cdot)\|^{(l+2)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (6.48)
-Note that $\tilde{r}\in\mathcal{H}^{(l)}_{T}$, by the results of [23, Lemma 14.1.6], we have
-$\left|\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]\right|\leq C\|\tilde{r}\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ (7.58)
-Combining these inequalities, we prove the desired result. ∎
+Note that $\tilde{r}\in\mathcal{H}^{(l)}_{T}$, by [25, Lemma 14.1.6], $\left|\mathbb{E}\left[\tilde{r}(t_{i},\widetilde{X}_{t_{i}})-\tilde{r}(s,\widetilde{X}_{s})\right]\right|\leq C\|\tilde{r}\|_{T}^{(l)}|\mathscr{G}|^{\mathcal{X}(l)}.$ Combining these inequalities yield the desired result.
+∎
 Proof of Proposition 5.3.
-For the variance part, note that
+For the variance, note that
 $\mathbb{E}\left[(S\circ V)^{2}\right]$ (6.49)
@@
 $=\mathbb{E}\left[\int_{0}^{T}S(t,X^{\mathscr{G}}_{t})g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right]-\mathbb{E}\left[\int_{0}^{T}\Delta S_{t}g(t,X^{\mathscr{G}}_{t},a_{\delta(t)})\mathrm{d}t\right],$
-where $\Delta S_{t}:=S(t,X^{\mathscr{G}}_{t})-S(\delta(t),X^{\mathscr{G}}_{\delta(t)})$. We have
+[…]
 $\left|\mathbb{E}[S\circ V]-\mathbb{E}^{W}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\left(\mathrm{d}V(t,\widetilde{X}_{t})+\tilde{r}(t,\widetilde{X}_{t})\mathrm{d}t\right)\right]\right|$ (6.52)
@@
 For the variance part, note that
-$\mathrm{Var}\left(S\Box V\right)$
-$\leq\mathbb{E}\left[(S\Box V)^{2}\right]$
+$\mathrm{Var}\left(S\Box V\right)\leq\mathbb{E}\left[(S\Box V)^{2}\right]$
 $=\mathbb{E}\left[\left(\sum_{i=1}^{n}S(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i}})\left(\int_{t_{i-1}}^{t_{i}}g_{1}(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}s+\int_{t_{i-1}}^{t_{i}}(\partial_{x}V(s,X^{\mathscr{G}}_{s}))^{\top}\sigma(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})\mathrm{d}W_{s}\right)^{2}\right)^{2}\right]$
@@
 It converges to $\mathbb{E}\left[\int_{0}^{T}\widetilde{Sg}_{2}(t,\widetilde{X}_{t})\mathrm{d}t\right]$ with error bounds $C_{p}(\|S\|_{C_{p}^{2,0}}\|g_{2}\|_{C_{p}^{0,0}}+\|Sg_{2}\|_{C_{p}^{2,0}}+\|\widetilde{Sg_{2}}\|_{C_{p}^{2}}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|_{C_{p}^{4}})|\mathscr{G}|$ and $C(\|S\|_{T}^{(l)}\|g_{2}\|_{T}^{(0)}+\|Sg_{2}\|_{T}^{(l)}+\|\widetilde{Sg_{2}}\|_{T}^{(l)}+\sup_{t\in[0,T]}\|\widetilde{Sg_{2}}(t,\cdot)\|^{(l)})|\mathscr{G}|^{\mathcal{X}(l)}$ under different types of conditions by the results of Proposition 5.1.
-In particular, if $S$ do not depend on $a$, then
+In particular, if $S$ does not depend on $a$, then
 $\mathbb{E}\left[\int_{0}^{T}\int_{A}S\partial_{x}V^{\top}\sigma\sigma^{\top}\partial_{x}V(t,\widetilde{X}_{t},a)\bm{\pi}(\mathrm{d}a|t,\widetilde{X}_{t})\mathrm{d}t\right]$ $=\mathbb{E}\left[\int_{0}^{T}S(t,\widetilde{X}_{t})\partial_{x}V(t,\widetilde{X}_{t})^{\top}(\tilde{\sigma}^{\bm{\pi}})^{2}(t,\widetilde{X}_{t})\partial_{x}V(t,\widetilde{X}_{t})\mathrm{d}t\right]$
@@
 $\partial_{x}^{4}\lambda^{\frac{1}{2}}$ $=-\frac{15}{16}\lambda^{-\frac{7}{2}}(\partial_{x}\lambda)^{4}+\frac{9}{4}\lambda^{-\frac{5}{2}}(\partial_{x}\lambda)^{2}\partial_{x}^{2}\lambda-\frac{3}{4}\lambda^{-\frac{3}{2}}(\partial_{x}^{2}\lambda)^{2}-\lambda^{-\frac{3}{2}}\partial_{x}\lambda\partial_{x}^{3}\lambda+\frac{1}{2}\lambda^{-\frac{1}{2}}\partial_{x}^{4}\lambda$
-Under the conditions of this proposition, we can exchange the order of integration and differentiation, therefore,
+Under the given conditions, we can exchange the order of integration and differentiation, and obtain
 $\partial_{x}\lambda=\partial_{x}\int_{\mathbb{R}}\sigma^{2}\varphi(u)\mathrm{d}u=\int_{\mathbb{R}}2\sigma(\partial_{x}\sigma+\partial_{a}\sigma(\partial_{x}\mu+\partial_{x}\Sigma u))\varphi(u)\mathrm{d}u.$ (6.58)
@@
 Note that $\sigma\sigma^{\top}$ is uniformly elliptic. There is a $\varepsilon>0$ such that $\lambda^{-1}\leq\varepsilon^{-1}$. The desired result follows from (6.60) and (6.57). ∎
-Appendix A Proofs of Lemmas 3.1 and 7.2
+Acknowledgments
+Yanwei Jia is supported by the Start-up Fund at The Chinese University of Hong Kong, and Hong Kong Research Grants Council (RGC) - Early Career Scheme (ECS) 24211124.
+We are especially indebted to the two anonymous referees for their constructive and detailed comments that have led to an improved version of the paper.
+We thank conference participants at IASM Workshop on “Bridging Theory and Practice in Finance: Stochastic Analysis and Machine Learning”, Workshop on Stochastic Control and Reinforcement Learning at Harbin Institute of Technology, 2nd ETH-Hong Kong-Imperial Mathematical Finance Workshop, and 15th POMS-HK International Conference for their helpful discussions.
+Appendix A Proofs of Lemmas 3.1, 6.1 and 6.3
 Proof of Lemma 3.1.
@@
 Note that $X^{\mathscr{G}}_{0}$ is deterministic and hence is $p$-th power integrable for all $p\geq 2$.
-Suppose that for some $i=0,\ldots,n-2$, (3.3) admits a unique strong solution $X^{\mathscr{G}}$ on $[0,t_{i}]$ satisfying ${\mathbb{E}}[\sup_{t\in[0,t_{i}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
+Suppose that for some $i=0,\ldots,n-2$, (3.3) admits a unique strong solution $X^{\mathscr{G}}$ on $0,t_{i}]$ satisfying ${\mathbb{E}}[\sup_{t\in[0,t_{i}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
 Observe that by (H.1) and the Cauchy-Schwarz inequality, for all $p\geq 2$,
@@
 ${\mathbb{E}}\left[\left(\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{2}{\mathrm{d}}t\right)^{p/2}\right]\leq{\mathbb{E}}\left[\int_{t_{i}}^{t_{i+1}}|\sigma(t,0,a_{t_{i}})|^{p}{\mathrm{d}}t\right]<\infty.$
-Since $x\mapsto b(t,x,a)$ and $x\mapsto\sigma(t,x,a)$ are Lipschitz continuous uniformly in $(t,a)$, by [40, Theorems 3.3.1 and 3.4.3], (3.3) with the initial state $X^{\mathscr{G}}_{t_{i}}$ at $t_{i}$ admits a unique strong solution $X^{\mathscr{G}}$ on $[t_{i},t_{i+1}]$ satisfying ${\mathbb{E}}[\sup_{t\in[t_{i},t_{i+1}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
+Since $x\mapsto b(t,x,a)$ and $x\mapsto\sigma(t,x,a)$ are Lipschitz continuous uniformly in $(t,a)$, by [46, Theorems 3.3.1 and 3.4.3], (3.3) with the initial state $X^{\mathscr{G}}_{t_{i}}$ at $t_{i}$ admits a unique strong solution $X^{\mathscr{G}}$ on $t_{i},t_{i+1}]$ satisfying ${\mathbb{E}}[\sup_{t\in[t_{i},t_{i+1}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
 This along with an inductive argument shows the desired well-posedness result.
@@
 Proof of Lemma 6.1.
-We first prove $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. Observe that by (H.5), for all $p\geq 2$,
+Assume that $s\in[t_{i-1},t_{i}]$, then $\delta(s)=t_{i-1}$. By Itô’s formula,
+$\mathbb{E}[g(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-g(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i_{1}}})]$ (A.5)
+$=\mathbb{E}\bigg[\int_{t_{i-1}}^{s}\left(\partial_{t}g(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})+b^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}g(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\right)\mathrm{d}u$
+$\quad+\int_{t_{i-1}}^{s}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}^{2}g(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\right)\mathrm{d}u\bigg].$
+It follows from the linear growth condition of $b$ and $\sigma$ that
+$\left|\mathbb{E}[g(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-g(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i_{1}}})]\right|$ (A.6)
+$\leq\int_{t_{i-1}}^{s}C\|g\|_{C_{p}^{2,0}}\mathbb{E}\big[(1+|X^{\mathscr{G}}_{u}|^{p}+d_{A}(a_{t_{i-1}},a_{0})^{p})\big(1+(1+|X^{\mathscr{G}}_{u}|+d_{A}(a_{t_{i-1}},a))$
+$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2}(1+|X^{\mathscr{G}}_{u}|+d_{A}(a_{t_{i-1}},a))^{2}\big)\big]\mathrm{d}u$
+$\leq C_{p}\|g\|_{C_{p}^{2,0}}|\mathscr{G}|,$
+for a constant $C_{p}$ depending only on $b,\sigma$ and $p$, where the last inequality used Lemma 3.1. ∎
+Proof of Lemma 6.3.
+We first prove $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. By (H.5), for all $p\geq 2$,
 $\mathbb{E}^{W}\left[\left(\int_{t_{i}}^{t_{i+1}}|b(t,0,a_{t_{i}})|\mathrm{d}t\right)^{p}\right]\leq C<\infty.$ (A.7)
@@
 Note that $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$, and all Itô integrals appear here equal to 0 when taking the conditional expectation $\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$.
-Analogous to the proofs of Lemma 3.1 and [40, Theorem 3.4.3], we obtain that $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
-Next, we aim to prove that the moments of $X^{\mathscr{G}}_{t}$ is independent of $|\mathscr{G}|$ and obtain (7.10).
-Use the notation $\delta(u):=t_{i}$ for $u\in[t_{i},t_{i+1})$, then it satisfies that
+Analogous to the proofs of Lemma 3.1 and [46, Theorem 3.4.3], we obtain that $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
+Next, we prove that the moments of $X^{\mathscr{G}}_{t}$ is independent of $|\mathscr{G}|$ and obtain (6.10). Use the notation $\delta(u):=t_{i}$ for $u\in[t_{i},t_{i+1})$, then it satisfies that
 $X^{\mathscr{G}}_{s}=x_{0}+\int_{0}^{s}b(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}u+\int_{0}^{s}\sigma(u,X^{\mathscr{G}}_{u},a_{\delta(u)})\mathrm{d}W_{u}.$ (A.9)